### v1

# Ant Colony Sampling with GFlowNets for Combinatorial Optimization

Abstract
This paper introduces the Generative Flow Ant Colony Sampler (GFACS), a novel neural-guided meta-heuristic algorithm for combinatorial optimization. GFACS integrates generative flow networks (GFlowNets) with the ant colony optimization (ACO) methodology. GFlowNets, a generative model that learns a constructive policy in combinatorial spaces, enhance ACO by providing an informed prior distribution of decision variables conditioned on input graph instances. Furthermore, we introduce a novel combination of training tricks, including search-guided local exploration, energy normalization, and energy shaping to improve GFACS. Our experimental results demonstrate that GFACS outperforms baseline ACO algorithms in seven CO tasks and is competitive with problem-specific heuristics for vehicle routing problems. The source code is available at https://github.com/ai4co/gfacs.
GFlowNets, Ant Colony Optimization, Neural Combinatorial Optimization
1 Introduction
Combinatorial Optimization Problems (COPs) constitute a category of optimization problems focused on finding the optimal arrangement of discrete structures, such as sets, sequences, and graphs, that maximizes a given objective function. COPs are prevalent in various fields, including compiler scheduling (Zhang et al., 2022b) and hardware design (Kim et al., 2023b), but they are usually NP-hard. Over the decades, heuristics that quickly find near-optimal solutions have been developed to tackle this NP-hardness. However, these problem-specific heuristics require considerable effort and domain knowledge to make them effective. Thus, meta-heuristic algorithms, such as Ant Colony Optimization (ACO) (Dorigo et al., 2006), have been proposed to provide a more generalized approach to designing heuristics.
ACO is a meta-heuristic algorithm inspired by the foraging behaviors of ants, and it has demonstrated effectiveness in solving COPs and general optimization problems. In ACO, a population of artificial ants iteratively builds solutions and updates ‘pheromone trails’ to explore the solution space. At each iteration, solutions are constructed through stochastic moves guided by instance-specific pheromone trails and problem-specific heuristic measures indicating the goodness of sampling certain trails; see Figure 1 for an illustration. While the pheromone trails are updated based on solutions constructed during the iterative procedure of solving an instance, the heuristic measures typically need to be predefined in a matrix form (e.g., heat map). However, developing an effective heuristic matrix for ACO involves expert knowledge and manual tuning, posing challenges in less-explored problem domains where the expertise is limited.
Figure 1: Illustration of ACO. Edge probability is obtained from the multiplication of heuristics and pheromone trails. Ants constructively complete solutions based on the edge probability, then update the pheromone trail with the constructed samples.
Recently, a neural-guided Ant Colony Optimization (ACO) method, DeepACO, was introduced by Ye et al. (2023). This technique utilizes graph neural networks (GNN) (Joshi et al., 2021) to model the heuristic matrix of input instance graphs of COPs. Specifically, this framework is developed as a GNN-based policy that selects an edge (action) based on previously chosen edges (state). Then, the composited policy generates a state trajectory from an empty set (initial state) to a complete solution (terminal state), as in Figure 1. DeepACO incorporates reinforcement learning (RL) to train the GNN-based policy model with a reward signal from the terminal state of the generated trajectory. After RL training, the learned GNN is utilized to construct the heuristic matrix for a given problem instance, thereby providing a good prior to ACO to explore the solution space more efficiently. Note that the trained GNN acts as a parameterized prior, effectively transferring the solution exploration strategy obtained during training to each new instance.
However, we observe certain limitations in DeepACO, primarily in three aspects. The first limitation ( ) is that the usage of a behavior policy is limited since DeepACO uses on-policy RL to train the graph representation. It potentially weakens the performance of heuristic matrices by limiting the choice of behavior policy, which is crucial for the effective exploration of massive combinatorial spaces. Secondly ( ), the reward-maximizing RL of DeepACO fails to satisfactorily capture the diversity of solutions, i.e., the multi-modality of the solution landscape. In CO, a diversity of good candidates is highly advantageous, as it increases the chances of finding a global optimum after a local search; see Figure 2. Lastly ( ), the inherent symmetry in combinatorial solutions, where multiple trajectories can lead to identical combinatorial outcomes, is challenging to consider in the context of RL. For example, in the Traveling Salesman Problem (TSP), distinct paths like (1,2,3,4) and (4,3,2,1) represent the same cyclic solution. As DRL methods train policy on a tree-structured MDP, each trajectory is identified as a different solution.
Figure 2: Importance of capturing multi-modality in combinatorial optimization with local search post-processing ( ). The probability of improved solutions rises after local search, even from initially higher-cost samples, particularly with a sampler adept at capturing multi-modality.
Contribution.
This paper introduces a novel ACO algorithm, the Generative Flow Ant Colony Sampler (GFACS), which addresses the limitations of DeepACO by integrating the emerging amortized inference method of GFlowNets (Bengio et al., 2021, 2023). First ( ), the off-policy training property of GFlowNets allows for the utilization of samples drawn from any behavior policy, including powerful heuristic search algorithms commonly used in COPs. Secondly ( ), unlike maximum reward-seeking algorithms (e.g., RL training in DeepACO), GFlowNets are sampling algorithms and well-suited for modeling the multi-modality nature of certain distributions. Lastly ( ), GFlowNets, which model the solution generative process on a directed acyclic graph (DAG), incorporate the symmetricity of combinatorial solutions by design (Bengio et al., 2023).
To adequately integrate GFlowNet into ACO, we propose an effective combination of training techniques. In our approach, we employ a strategic guided exploration with two distinct behavior policies to collect exploratory and exploitation samples. One aims for a diverse range of samples, and the other aims for locally optimized samples. This approach effectively addresses the over-exploration issue in GFlowNets, which often results in poor performances due to an excessive focus on diversity. To encourage further exploration, we introduce an energy reshaping to compensate samples that have initially high energy but can be significantly improved, as exemplified by the red samples in Figure 2. At last, a shared energy normalization is employed to stabilize the training of conditional GFlowNets, wherein the energy landscape is contingent upon the variability of input problems.
Our GFACS exhibits significant superiority over the current DeepACO and classical ACO algorithms for various representative CO tasks, which vehicle routing problems, scheduling problems, and a grouping problem. Moreover, GFACS outperforms or shows similar performance to competitive deep reinforcement learning-based vehicle routing solvers, which are designed to solve traveling salesman problems (TSP) and capacitated vehicle routing problems (CVRP) in the context of large-scale instances, such as $N=500,1000$.
2 Related Works
2.1 Ant Colony Optimization
Ant Colony Optimization (ACO) is a meta-heuristic inspired by ant behavior in finding optimal paths. Traditional ACO algorithms, such as Ant System (AS) (Dorigo et al., 1996) and MAX-MIN AS (Stützle & Hoos, 2000), rely on specific heuristics for success. To reduce the reliance on heuristic design, Ye et al. (2023) introduce a neural-guided ACO, called DeepACO, utilizing graph representation learning and reinforcement learning to replace manual heuristics with automatically acquired knowledge.
2.2 Neural Combinatorial Optimization
Neural combinatorial optimization (Bello et al., 2016) is a deep learning-based method to train an instance-conditional policy using supervised learning (Vinyals et al., 2015; Fu et al., 2021; Sun & Yang, 2023) or unsupervised learning (Khalil et al., 2017; Kool et al., 2018; Kwon et al., 2020). Usually, supervised learning methods are beneficial when the target problem has an accurate but time-consuming expert solver. The oracle-labeled data is collected from the oracle solvers, and deep neural networks amortize them, allowing for scalability and faster inferences than the expert solver. On the other hand, unsupervised learning proceeds through self-guided trial and error, using the COP objective as a training signal. This makes it applicable even if encountering less-studied problems where powerful expert solvers are unavailable (Bello et al., 2016; Ahn et al., 2020). Our method can be categorized as an unsupervised learning method, which does not need expert labels. See Appendix F for a more comprehensive survey.
2.3 GFlowNets and Combinatorial Optimization
GFlowNets, tailored for sampling combinatorial objects proportionally to a given reward function, play a pivotal role in diversifying the candidate pool for molecule generation in crucial areas like drug discovery (Bengio et al., 2021). Recent efforts to advance GFlowNets have focused on developing new training objectives (Malkin et al., 2022; Madan et al., 2022; Zhang et al., 2023b), enhancing credit assignment (Pan et al., 2023a; Jang et al., 2024), and improving exploration strategies (Pan et al., 2022; Rector-Brooks et al., 2023; Pan et al., 2023a; Kim et al., 2023d, c). Our contribution lies in proposing an efficient exploration technique tailored for combinatorial optimization problems.
GFlowNets have also found successful applications in some COPs (Zhang et al., 2022b, 2023a). Zhang et al. (2023a) effectively addresses the symmetric nature of graph constraint optimization, showcasing remarkable performance in graph CO tasks like maximum independent sets. Zhang et al. (2022b) solves proxy-based scheduling problems. Despite their success, their method has limited applications to either locally decomposable graph CO problems (Ahn et al., 2020) or robust scheduling problems. Our method is a neural-guided meta-heuristic, which aims to be a general-purpose CO solver. Additionally, this paper appears to be the first to address the traveling salesman problem and its variations using GFlowNets. These tasks present a significant challenge due to the global constraints to compare with graph CO tasks.
3 Preliminary
3.1 Travelling Salesman Problem
The traveling salesman problem (TSP) is a fundamental combinatorial optimization problem, which serves as a building block for more complex CO problems like the capacitated vehicle routing problem (CVRP) and orienteering problem (OP). In light of this and for simplicity, this paper presents formulations using the two-dimensional Euclidean TSP as an illustrative example. Note that the formulations can be extended to other COPs straightforwardly.
A TSP instance is defined on a (fully connected) graph $\mathcal{G}=\{\mathcal{V},\mathcal{D}\}$, where $\mathcal{V}$ and $\mathcal{D}$ are a node set and edge set, respectively. Each node $i\in\mathcal{V}$ has two-dimensional coordinates $v_{i}\in[0,1]^{2}$ as a feature, and each edge $(i,j)\in\mathcal{D}$ has a feature $d_{i,j}$ which is the distance between node $i$ and $j$, i.e., $d_{i,j}=\|v_{i}-v_{j}\|_{2}$. A solution to TSP, $x(\mathcal{G})$ corresponds to a Hamiltonian cycle for a given problem instance $\mathcal{G}$, which can be represented as a permutation of nodes $\pi=(\pi_{1},\ldots,\pi_{N})$. The tour length is computed as $f(x(\mathcal{G}))=\sum_{t=1}^{N-1}d_{\pi_{t},\pi_{t+1}}+d_{\pi_{N},\pi_{1}}$. The optimization objective for TSP is to find $x(\mathcal{G})$ that minimizes $f(\cdot,\mathcal{G})$.
3.2 Ant Colony Optimization
Ant colony optimization is a meta-heuristic to refine the solution sampling probability by an iterative procedure of solution construction and pheromone update. Given a problem instance $\mathcal{G}$, a corresponding heuristic matrix $\eta(\mathcal{G})\in\mathbb{R}^{N\times N}$ and a pheromone trails $\rho\in\mathbb{R}^{N\times N}$, a solution is constructed by auto-regressively sampling edges from the probability distribution:
$P(\pi_{t+1}=j|\pi_{t}=i;\rho,\eta(\mathcal{G}))\propto\begin{cases}\rho_{i,j}\eta_{i,j}(\mathcal{G})&\text{if $j\notin\pi_{1:t}$ }\\ 0&\text{otherwise},\end{cases}$ (1)
where $\rho_{i,j}$ and $\eta_{i,j}(\mathcal{G})$ are the pheromone weight and heuristic weight for edge $d_{i,j}$, respectively.
At each iteration, $K$ artificial ants independently sample solutions $\{\pi^{k}\}_{k=1}^{K}$ by following Equation 1. The pheromone trails $\rho$ are then adjusted using the solution set and corresponding costs, thereby updating the solution construction probability. The pheromone update rules vary depending on the ACO algorithm used. For example, in Ant System (AS) (Dorigo et al., 1996), all $K$ solutions contribute to the pheromone update:
$\rho_{i,j}\leftarrow(1-\gamma)\rho_{i,j}+\sum\limits_{k=1}^{K}\Delta^{k}_{i,j},$ (2)
where $\gamma$ is the decaying coefficient (or evaporation rate) and $\Delta^{k}_{i,j}$ is the amount of pheromone laid by the $k$-th ant on the edge $(i,j)$. Here, $\Delta^{k}_{i,j}$ is proportional to the inverse of tour length if $(i,j)$ is included in $\pi^{k}$, and zero otherwise.
Note that the heuristic matrix $\eta(\mathcal{G})\in\mathbb{R}^{N\times N}$ is not updated during the search process, only serving as a prior. Thus, ACO aligns the search process more closely with the most efficacious paths discovered; see Dorigo et al. (2006) for details. In DeepACO (Ye et al., 2023), the learned GNN, parameterized by $\theta$, is utilized to construct the heuristic matrix $\eta(\mathcal{G},\theta)$ for a given problem instance $\mathcal{G}$, thereby providing a good prior for exploring the solution space.
Figure 3: Solution sampling process of GFACS. The GNN is trained with GFlowNets and serves as an expert heuristic in GFACS.
3.3 Generative Flow Networks
Generative Flow Networks, or GFlowNets (Bengio et al., 2021, 2023) provide an effective training framework for learning a policy that samples compositional objects denoted as $x\in\mathcal{X}$. GFlowNets follow a constructive generative process, using discrete actions to iteratively modify a state which represents a partially constructed object. All the ways to construct these objects can be described by a directed acyclic graph (DAG), $\mathcal{H}=(\mathcal{S},\mathcal{A})$, where $\mathcal{S}$ is a finite set of all possible states, and $\mathcal{A}$ is a subset of $\mathcal{S}\times\mathcal{S}$, representing directed edges. Within this framework, we define the children of state $s\in\mathcal{S}$ as the set of states connected by edges whose head is $s$, and the parents of state $s$ as the set of states connected by edges whose tail is $s$.
We define a complete trajectory $\tau=(s_{0}\rightarrow\ldots\rightarrow s_{N})\in\mathcal{T}$ from the initial state $s_{0}$ to terminal state $s_{n}$. We denote $\tau_{{}_{\rightarrow}x}$ when the terminal state $s_{n}$ is finalized to the corresponding object $x\in\mathcal{X}$. We define the forward policy to model the forward transition probability $P_{F}(s^{\prime}|s)$ from $s$ to its child $s^{\prime}$. Similarly, we also consider the backward policy $P_{B}(s|s^{\prime})$ for the backward transition $s^{\prime}\dashrightarrow s$, where $s$ is a parent of $s^{\prime}$. We can use the forward policy to compute in a forward way the probability of a trajectory:
$P_{F}(\tau;\theta)=\prod_{t=1}^{n}P_{F}\left(s_{t}|s_{t-1};\theta\right)$ (3)
Similarly, we can use the backward policy to compute in a backward way the probability of trajectory ending at a given object $x$:
$P_{B}(\tau_{\rightarrow x}|x;\theta)=\prod_{t=1}^{n}P_{B}\left(s_{t-1}|s_{t};\theta\right)$ (4)
Trajectory balance (Malkin et al., 2022).
Trajectory balance is the most widely used learning objective in a GFlowNet framework. When the trajectory balance objective is minimized, we achieve that the probability of sampling $x$ with the forward policy is $P_{F}^{\top}(x)\propto R(x)$, as desired, where $P_{F}^{\top}(x)=\sum_{\tau_{\rightarrow x}\in\mathcal{T}}P_{F}(\tau)$. The trajectory balance loss $\mathcal{L}_{\text{TB}}$ works by training three models, a learnable scalar $Z_{\theta}$ for the partition function, a forward policy $P_{F}(\tau;\theta)$, and a backward policy $P_{B}(\tau_{\rightarrow x}|x;\theta)$ to minimize the following objective:
$\mathcal{L}_{\text{TB}}(\tau;\theta)=\left(\log\frac{Z_{\theta}P_{F}(\tau;\theta)}{R(x)P_{B}(\tau_{\rightarrow x}|x;\theta)}\right)^{2}$ (5)
4 Generative Flow Ant Colony Sampler
4.1 MDP formulation
This subsection formulates a Markov decision process (MDP) by using TSP as an example, similar to the preliminaries. A TSP solution is obtained through the sequential decision-making process of GFlowNets as follows:
State $s$:
The initial state $s_{0}$ is an empty set. The state $s_{t}$ at $t>0$ is a sequence of previous actions: $(a_{0},\ldots,a_{t-1})$.
Action $a$:
The action $a_{t}$ is a selection of a node to visit at time $t$ among un-visited nodes: $a_{t}\in\{1,\ldots,N\}\setminus\{a_{0},\ldots,a_{t-1}\}$. If all nodes are visited, a deterministic terminating action is applied, which maps a final state $s_{N}$ into a solution, $x(\mathcal{G})$.
Reward $R(x(\mathcal{G});\beta)$:
The reward is defined as the unnormalized energy distribution of the solution $x(\mathcal{G})$, i.e.,
$R(x(\mathcal{G});\beta)=e^{-\beta\mathcal{E}(x(\mathcal{G}))}=e^{-\beta f(x(\mathcal{G}))}.$ (6)
Here, $\mathcal{E}(x(\mathcal{G}))$ is energy for solution $x(\mathcal{G})$, which is the same as the tour length of the solution for TSP, i.e., $\mathcal{E}(x(\mathcal{G}))=f(x(\mathcal{G}))$. The $\beta$ is the inverse energy temperature, which is a hyperparameter.
4.2 Policy Configuration
We parameterize the forward policy $P_{F}$ using a graph neural network (GNN) with parameter $\theta$ that maps an input graph instance $\mathcal{G}$ to the output edge feature $\eta(\mathcal{G};\theta)$, following Ye et al. (2023). Then, the forward policy when $t<N$ is expressed as follows:
$\begin{split}&P_{F}(s_{t+1}|s_{t};\eta(\mathcal{G};\theta),\rho)\\ &\propto\begin{cases}\rho_{a_{t-1},a_{t}}\eta_{a_{t-1},a_{t}}(\mathcal{G};\theta)&\text{if $a_{t}\notin s_{t}$ }\\ 0&\text{otherwise}\end{cases}\end{split}$ (7)
Note that $\rho$ and $\eta(\mathcal{G};\theta)$ denote the pheromone trails and learned heuristic matrix in ACO.
Subsequently, we can define a trajectory from the initial state to a terminal solution $\tau=(s_{0},\ldots,s_{N},x(\mathcal{G}))$ and the forward probability of that trajectory:
$P_{F}(\tau;\eta(\mathcal{G};\theta),\rho)=\prod_{t=0}^{N-1}P_{F}(s_{t+1}|s_{t};\eta(\mathcal{G};\theta),\rho)$ (8)
Note that the forward probability for the transition from $s_{N}$ to $x(\mathcal{G})$ is omitted as it follows a deterministic rule.
For the backward policy, we set $P_{B}=\frac{1}{2N}$, emulating the maximum entropy GFlowNets (Zhang et al., 2022a) to make uniform exploration over symmetric trajectories which terminate into identical solution. Here, $P_{B}$ represents a uniform distribution across all $2N$ feasible trajectories leading to a solution $x(\mathcal{G})$. This is particularly pertinent in the context of the TSP involving $N$ cities, where $2N$ permutations culminate in the same Hamiltonian cycle $x(\mathcal{G})$. The count of symmetric trajectories depends on the task’s characteristics. For an exhaustive discussion on solution symmetry across various COPs, please refer to Appendix E.
4.3 Solution Sampling with GFACS
In the sampling process, we assume the existence of the learned GNN parameterized with $\theta^{*}$; see Section 5 for details of the training procedure. The learned GNN provides the heuristic matrix $\eta(\mathcal{G};\theta^{*})$ for ACO algorithm. Thus, the remaining process is to update pheromone parameter $\rho$ with solutions sampled by $P_{F}(\tau;\eta(\mathcal{G};\theta^{*}),\rho)$ during the solution searching process. As the $\eta(\mathcal{G};\theta^{*})$ is trained with GFlowNets to sample diverse high-rewarded (i.e., low energy) samples, $P_{F}$ generates good candidates for updating $\rho$. The overall process is illustrated in Figure 3.
5 Training GFACS
This section introduces how to train GFACS, especially the GNN component $\theta$, in the GFlowNet framework. GFACS is trained via iterations of the following two steps. First (Step A; Section 5.1), we collect experiences, each being a triplet of input instance graph $\mathcal{G}$, a complete trajectory $\tau_{\rightarrow x(\mathcal{G})}$, and the energy of the solution $\mathcal{E}(x(\mathcal{G}))$. We employ $\beta$ annealing and guided exploration for better exploration and energy reshaping for proper credit assignment. Next (Step B; Section 5.1), the parameterized forward policy $P_{F}$ is trained using the collected experiences. Here, we utilize a shared energy normalization to stabilize training.
5.1 Step A: Experience Collection
For every iteration, we first sample $M$ problem instances from a random sampler $P(\mathcal{G})$, i.e., $\mathcal{G}_{1},\ldots,\mathcal{G}_{M}\sim P(\mathcal{G})$. Details of instance sampling are described in Appendix B.
Collecting exploration samples with $\beta$ annealing. For each problem instance $\mathcal{G}_{m}$, we collect $K$ exploration solutions by following Eq. 8, resulting in an exploration batch, $\mathcal{B}_{\text{explore}}=\{(\mathcal{G}_{m},\tau^{m,k}_{\rightarrow x_{m,k}},\mathcal{E}(x_{m,k})\}_{m=1,\ldots,M,k=1,\ldots,K}$. To encourage exploration for solution sampling, we introduce an annealing scheme for the inverse energy temperature $\beta$. Rather than set the $\beta$ to a fixed value, we gradually increase the beta toward the target value along the training procedure. This makes the policy in the early stage of training to be exploratory, preventing it from getting stuck into a local optimum.
Figure 4: Training of GFACS using GFlowNets and our novel training tricks of guided exploration and energy reshaping.
Collecting exploitation samples with guided exploration. To cope with the over-exploration problem of GFlowNet, we also collect $K$ exploitation samples for each $\mathcal{G}_{m}$ using guided exploration. We use a heuristic local search operator ($LS$), such as 2-opt for TSP, as external guidance. The local search operator directly refines a given solution into a locally optimized one, i.e., $x^{\prime}\leftarrow LS(x)$. We then use $P_{B}$ to collect a trajectory $\tau_{\rightarrow x^{\prime}}$ for each $x^{\prime}$, obtaining a batch of exploitation experiences, $\mathcal{B}_{\text{exploit}}=\{(\mathcal{G}_{m},\tau^{m,k}_{\rightarrow x^{\prime}_{m,k}},\mathcal{E}(x^{\prime}_{m,k})\}_{m=1,\ldots,M,k=1,\ldots,K}$.
However, a powerful local search operator may not exist for some problems, especially for less-explored ones. In such cases, we use the Top-$L$ guidance technique, described as follows. For a given problem instance $\mathcal{G}_{m}$, we first update the pheromone trails using the exploration samples collected before. After that, $L\times K$ new sample solutions are generated with the updated pheromone, where $L$ is a hyperparameter. Among the new samples, we select top $L$ solutions with the lowest energy and also select $(K-L)$ random solutions, resulting in $K$ exploitation samples for each $\mathcal{G}_{m}$, and thus obtaining $B_{\text{exploit}}$ like before.
Energy reshaping for exploratory samples. In the case where we have a local search operator, we reshape the energy for the exploration $x$ in $\mathcal{B}_{\text{explore}}$ using the energy of the refined solution $x^{\prime}=LS(x)$ as follows:
$\mathcal{E}_{\text{reshaped}}(x)=\alpha\mathcal{E}(x^{\prime})+(1-\alpha)\mathcal{E}(x)$ (9)
This is to evaluate an underrated sample that initially had high energy (i.e., low reward) but can be easily changed to have low energy (i.e., high reward) after the local search (see Figure 2). This reshaping changes the reward values in $\mathcal{B}_{\text{explore}}$ accordingly, while $\mathcal{B}_{\text{exploit}}$ remains the same. Note that this reshaping cannot be applied for TopK guidance because $x^{\prime}$ cannot be determined solely with $x$ in that case.
5.2 Step B: Training with Collected Experiences
Given the collected experience batches $\mathcal{B}_{\text{explore}}$ and $\mathcal{B}_{\text{exploit}}$, we aim to minimize the following loss function:
$\mathcal{L}(\theta)=\mathcal{L}(\theta;\mathcal{B}_{\text{explore}})+\mathcal{L}(\theta;\mathcal{B}_{\text{exploit}}),$ (10)
where ($\begin{split}&\mbox{where }\mathcal{L}(\theta;\mathcal{B})\\ &=\sum_{m=1}^{M}\sum_{k=1}^{K}\left(\log\frac{2NP_{F}(\tau^{m,k}_{\rightarrow x_{m,k}};\eta(\mathcal{G}_{m};\theta),\rho)Z_{\theta}(\mathcal{G}_{m})}{R(x_{m,k};\beta)}\right)^{2}.\end{split}$ (11)
The $2N$ comes from the fact that TSP has $2N$ permutations, as discussed in Section 4.2. In addition, our target sampling distribution is proportional to reward $p(x;\mathcal{G})\propto R(x(\mathcal{G}))$. However, in most COPs, including TSP, the scale of $R(x(\mathcal{G}))$ is significantly different, possibly leading to unstable training due to the high variance of rewards. Similarly to previous works that stabilize REINFORCE training (Kwon et al., 2020; Kim et al., 2022b), we normalize the energy $\mathcal{E}=-\log R$ using the sample mean over a $K$ samples for the same $\mathcal{G}_{m}$, i.e.,
$-\log\tilde{R}(x_{m,k})=\tilde{\mathcal{E}}(x_{m,k})=\mathcal{E}(x_{m,k})-\frac{1}{K}\sum_{k=1}^{K}\mathcal{E}(x_{m,k}).$ (12)
We call this technique a shared energy normalization. Note that the shared energy normalization is applied independently to the two batches, $\mathcal{B}_{\text{explore}}$ and $\mathcal{B}_{\text{exploit}}$, as the energy distribution between the two batches is often significantly different to each other, especially when a heuristic local search is applied.
Figure 5: Results of ACO algorithms on various CO tasks. Our GFACS outperforms every ACO baseline. The results are averaged over 3 independent models evaluated on the held-out test sets, and the shade indicates the min-max range of the 3 models. GFACS, without guided exploration, gives poorer performances. This verifies the effectiveness of our algorithm.
6 Experiments
This section presents the experimental findings of the GFACS framework. First, we test GFACS against established ACO baselines in various CO tasks. Subsequently, we explore the trade-offs between diversity and optimality in relation to the inverse energy temperature $\beta$. Next, an ablation study is conducted to evaluate the effectiveness of our proposed training techniques. Lastly, we compare GFACS with the RL-based combinatorial optimization methods in two representative routing problems, the TSP and the capacitated vehicle routing problem (CVRP).
6.1 Setup
Implementation and hyperparameters. GFACS is built upon the implementation established by DeepACO (Ye et al., 2023). As our methodology mainly focuses on improving training algorithms rather than network architecture, we simply adopt the GNN architecture proposed in DeepACO, except the additional parameters for partition function $Z_{\theta}$ in Eq. 11. For detailed network architecture and hyperparameters, please refer to Appendix A. One important hyperparameter that GFACS newly introduces is the inverse energy temperature $\beta$, described in Eq. 6. We offer an analysis of the impact and significance of $\beta$ in Section 6.4.
Test settings. Each algorithm is tested on the held-out datasets generated or obtained in advance. Unless otherwise stated, the results presented throughout this section are the averaged values from three independent models, each trained with a distinct random seed.
6.2 Comparison with ACO algorithms
Here, we compare GFACS against the classical ACO and the recent DeepACO algorithm across various CO tasks.
Problems. Following the approach of Ye et al. (2023), we evaluate our algorithm in a range of COPs. This includes four routing problems – TSP, Capacitated Vehicle Routing Problem (CVRP), Prize-Collecting Vehicle Routing Problem (PCTSP), and Orienteering Problem (OP) – along with two scheduling problems – Sequence Ordering Problem (SOP) and Single Machine Total Weighted Tardiness Problem (SMTWTP) – and a grouping problem, the Bin Packing Problem (BPP). For detailed problem configurations, we direct readers to Appendix B or Ye et al. (2023) as we mostly follow their problem settings. For TSP and CVRP, we use local search algorithms with neural-guided perturbation following Ye et al. (2023). For other problems, we use TopK guidance as described in Section 5.1.
Baselines. We use ACO and DeepACO as baseline algorithms. Both baselines and GFACS use Ant System (Dorigo et al., 1996) for pheromone update (Eq. 2). We also test GFACS without guided exploration to analyze its effect on the performance. We use 100 ants and $T=100$ ACO iterations for all the algorithms.
Results. As shown in Figure 5, GFACS consistently surpasses both the classic ACO algorithm and DeepACO across all benchmark problems. Especially for routing problems, GFACS outperforms DeepACO by a considerable margin. This is particularly remarkable given that both algorithms utilize identical neural networks and local search operators, highlighting the crucial role of our training strategy. Without guided exploration, GFACS generally performs badly, likely due to the over-exploration problem of GFlowNets. This issue is effectively mitigated with the proposed guided exploration, showing the efficacy of our method. Interestingly, in CVRP, OP, and PCTSP, GFACS without guided exploration outperforms DeepACO after a certain number of ACO iterations. We believe that this is attributed to the increased solution diversity of GFACS, which comes from the generative nature of the policy and implied DAG structure that incorporates the solution symmetricity.
Figure 6: The trade-off between diversity and optimality as the inverse energy temperature $\beta$ changes. By adjusting $\beta$, we can control the trade-off of having a Pareto frontier to compare with DeepACO. The reported values are obtained with $T=1$, i.e., the solutions are not affected by the pheromone.
6.3 Analysis of $\beta$ and Diversity-Optimality Trade-off
In this subsection, we explored the diversity-optimality trade-off by varying the inverse energy temperature $\beta$ (Eq. 6) from low to high values. For analysis, we generate $K$ solution samples from policies trained with GFACS with different values of $\beta$ and DeepACO (with $\rho=1$ so that the pheromone does not affect the result). We use average pairwise Jaccard distance as a diversity metric (formally defined in Appendix D). Results for TSP200 (Figure 6) show that lower $\beta$ values lead to more diverse but less optimal solutions, while higher $\beta$ values result in less diversity but higher optimality. GFACS confirms the anticipated trade-off between diversity and optimality, establishing a Pareto frontier over DeepACO by offering more diverse and high-quality solutions. This advantage is particularly beneficial when an effective local search mechanism is available, increasing the likelihood of achieving superior local optima, as illustrated in Figure 2. However, excessively high $\beta$ (e.g., $2000$ in our result) degrades the optimality, supposedly because too high $\beta$ value prevents exploration during training, which in turn results in sub-optimality. This empirical evidence supports our design choice of $\beta$ annealing, which encourages exploration in the early stage of training.
Table 1: An ablation study examining various design elements of our method for TSP200 and CVRP200. Here, the optimality gap is used as a metric, where lower values indicate better performance.
Method TSP200 CVRP200
Ours 0.26% 1.09%
-- $\beta$ annealing 0.27% 1.11%
-- guided exloration 0.88% 1.42%
-- energy reshaping 0.59% 1.09%
-- shared energy normalization 3.11% 2.88%
6.4 Ablation Study
We also conducted an ablation study for four key training techniques introduced in Section 5: $\beta$ annealing, guided exploration, energy reshaping, and shared energy normalization. The ablation for $\beta$ annealing is conducted by fixing $\beta$ to the maximum value of the annealed $\beta$ schedule, while for other elements, we simply remove them from our algorithm. The empirical results for TSP200 and CVRP200 reported in Table 1 validate the effectiveness of our design choices.
Table 2: Experimental result on TSP. Results for methods with * are drawn from Ye et al. (2023), Cheng et al. (2023a), and Jin et al. (2023). The reported time is the average computation time for solving a single instance.
TSP200 TSP500 TSP1000
Method Obj. Gap(%) Time(s) Obj. Gap(%) Times(s) Obj. Gap(%) Time(s)
LKH-3 10.72 0.00 9.80 16.55 0.00 31.70 23.14 0.00 102
AM* 12.00 11.91 0.2 21.46 29.07 0.6 33.55 45.10 1.5
POMO* 10.89 1.62 0.10 20.57 24.40 0.6 32.90 42.30 4.1
POMO + EAS 10.75 0.30 120 18.25 10.27 300 30.77 32.96 600
POMO + SGBS 10.77 0.43 133 18.86 13.96 318 28.58 23.51 605
DIMES+MCTS* - - - 17.01 2.78 11 24.45 5.75 32
SO* 10.78 0.63 10 16.94 2.40 15 23.77 2.80 26
Pointerformer* 10.79 0.68 5.5 17.14 3.56 14 24.80 7.90 40
H-TSP* - - - - - - 24.65 6.62 0.3
ACO $(T=10)$ 10.91 1.76 2.1 17.38 5.02 14 24.76 6.99 57
DeepACO $(T=10)$ 10.77 0.45 2.5 16.84 1.75 15 23.78 2.77 66
GFACS $(T=10)$ 10.75 0.26 2.5 16.80 1.54 15 23.72 2.53 66
Table 3: Experimental result on CVRP. The reported time is the average computation time for solving a single instance.
CVRP200 CVRP500 CVRP1000
Method Obj. Gap(%) Time Obj. Gap(%) Time Obj. Gap(%) Time
LKH-3 28.02 0.00 2.4m 63.07 0.00 18.6m 119.52 0.00 2.1h
POMO 29.12 3.94 1s 68.69 8.90 2s 145.74 21.94 3s
POMO + EAS 27.99 -0.11 2.0m 64.76 2.67 5.0m 126.25 5.63 10.0m
POMO + SGBS 28.07 0.19 2.0m 65.44 3.75 5.2m 127.90 7.01 10.1m
SymNCO 29.15 4.03 1s 68.81 9.10 2s 141.82 18.66 3s
SymNCO + EAS 28.01 0.27 2.0m 64.63 2.47 5.0m 125.58 5.07 10.0m
SymNCO + SGBS 28.10 0.30 2.0m 65.40 3.69 5.2m 127.53 6.70 10.1m
ACO $(T=10)$ 28.45 1.55 13s 64.62 2.45 36s 122.82 2.76 1.3m
DeepACO $(T=10)$ 28.44 1.49 14s 64.49 2.25 36s 122.15 2.20 1.3m
GFACS $(T=10)$ 28.32 1.09 14s 64.19 1.76 36s 121.80 1.91 1.3m
6.5 Comparison with Other Reinforcement Learning Methods for Combinatorial Optimization
In the domain of machine learning for combinatorial optimization, significant strides have been made using various approaches, including sampling-based methods (Sun et al., 2023), heatmap-guided searches (Sun & Yang, 2023; Qiu et al., 2022), and constructive methods that employ (Drakulic et al., 2023; Kool et al., 2018; Kwon et al., 2020). Our study compares our methods with RL-based approaches in TSP and CVRP for sizes $N=200,500,1000$, which do not require expert-labeled data. This comparison is particularly relevant as our unsupervised approach aligns with sequence decision-making processes akin to RL.
Baselines. For TSP baselines, we select the AM (Kool et al., 2018), POMO (Kwon et al., 2020), Sym-NCO (Kim et al., 2022b) and Pointerformer (Jin et al., 2023) as RL-based constructive methods, DIMES (Qiu et al., 2022) as RL-based heatmap-based methods, the Select and Optimize (SO) (Cheng et al., 2023b) as RL-based improvement methods, and H-TSP (Pan et al., 2023b) as a RL-based hierarchical methods. We also applied two search algorithms, the Efficient Active Search (EAS) (Hottung et al., 2021) and Simulation Guided Beam Search (SGBS) (Choo et al., 2022), to POMO for a more thorough comparison. For CVRP [^1] We also adopt TAM as a CVRP baseline; however, due to the inaccessibility of source code, we compare it using the reported result in the paper (Hou et al., 2022). See Section C.2.. We choose POMO and Sym-NCO as baselines with EAS and SGBS. Note that we restrict the duration of these searches to no more than 2, 5, and 10 minutes for problems with 200, 500, and 1000 nodes, respectively. We use LKH-3 (Helsgaun, 2017) as an oracle heuristic solver for TSP and CVRP to obtain near-optimal solutions that serve as baselines with $0.00\%$ gap. Please refer to Section A.4 for the parameter settings of LKH-3.
Results. Among the baseline methods evaluated, our GFACS algorithm consistently delivers either superior or highly competitive performances. When compared to faster methods like AM, POMO or H-TSP, GFACS demonstrates substantially better results. Even after an additional search with EAS or SGBS, they fail to match GFACS’s achievements, except in CVRP200. Compared to algorithms like DIMES and Pointformer, which operate at similar speeds, GFACS achieves improved performance. GFACS performs slightly worse than the oracle heuristic solver LKH-3, but offers an advantage of faster inference, especially in CVRP. Therefore, we can conclude that GFACS, as a general-purpose algorithm, holds significant appeal in the highly competitive benchmarks of both TSP and CVRP.
7 Conclusion
This paper proposes a novel neural-guided meta-heuristic algorithm called generative flow ant colony sampler (GFACS). Our algorithm was designed by integrating generative flow networks (GFlowNets) and ant colony optimization (ACO) with a novel combination of training techniques. GFACS consistently outperforms other ACO methods in several representative combinatorial optimization tasks. GFACS also shows competitive results both in traveling salesman problems and capacitated vehicle routing problems compared to a range of RL-based methods.
Appendix A Implementation
A.1 Learning architecture and optimizer
We utilize the anisotropic graph neural network with edge gating mechanisms, which is proposed by (Joshi et al., 2021) and adopted by (Ye et al., 2023), as the foundational GNN architecture for the construction of a heuristic matrix. This network offers node and edge embeddings that are more beneficial for edge prediction than other popular GNN variants, such as Graph Attention Networks (Veličković et al., 2018) or Graph Convolutional Networks (Kipf & Welling, 2016).
We represent the node feature of node $v_{i}$ at the $\ell$-th layer as $h_{i}^{\ell}$, and the feature of the edge between nodes $i$ and $j$ at the same layer as $e_{ij}^{\ell}$. These features undergo propagation to the next layer $\ell+1$ through an anisotropic message passing scheme, detailed as follows:
$h_{i}^{\ell+1}=h_{i}^{\ell}+\alpha(\text{BN}(U^{\ell}h_{i}^{\ell}+\mathcal{A}_{j\in\mathcal{N}_{i}}(\sigma(e^{\ell}_{ij})\odot V^{\ell}h_{j}^{\ell})))$ (13)
$m^{\ell}=P^{\ell}e_{ij}^{\ell}+Q^{\ell}h_{i}^{\ell}+R^{\ell}h_{j}^{\ell}$ (14)
$e_{ij}^{\ell+1}=e_{ij}^{\ell}+\alpha(\text{BN}(m^{\ell}))$ (15)
where $U^{\ell},V^{\ell},P^{\ell},Q^{\ell},R^{\ell}\in R^{d\times d}$ are the learnable parameters of layer $\ell$, $\alpha$ denotes the activation function, BN(Ioffe & Szegedy, 2015) denotes batch normalization, $\mathcal{A}$ denote the aggregation function, $\sigma$ is the sigmoid function, $\odot$ is Hadamard product, $\mathcal{N}_{i}$ denotes the neighborhoods of node $v_{i}$. In this paper we use $\alpha$ is SiLU(Elfwing et al., 2018) and $\mathcal{A}$ is mean pooling. We transform the extracted edge features into real-valued heuristic metrics by employing a 3-layer Multilayer Perceptron (MLP) that includes skipping connections to the embedded node feature $h$. The SiLU is the activation function for all layers except the output layer, in which the sigmoid function is applied to generate normalized outputs. The training process is optimized using the AdamW optimizer(Loshchilov & Hutter, 2017) in conjunction with a Cosine Annealing Scheduler(Loshchilov & Hutter, 2016)
A.2 Hyperparameters
For a task of COP (e.g., TSP200), we train GFACS for $N_{\text{epoch}}$ epochs. In each epoch, we use $N_{\text{inst}}$ number of instances using mini-batch with size $M$; thus $N_{\text{inst}}/M$ gradient steps are taken per epoch. For each problem instance, we sample $K$ solutions to train with, as discussed in Section 5.1. Our main baseline algorithm, DeepACO, is also trained using the same hyperparameters to ensure a fair comparison. We summarize used value for $N_{\text{epoch}}$, $N_{\text{inst}}$, $M$ and $K$ in Table 4. Note that for TSP1000 and CVRP1000, we use model checkpoints pre-trained in TSP500 and CVRP500, respectively, to reduce the processing time for training.
Table 4: Training hyperparameter settings for each task. These hyperparameter settings are used both for GFACS and DeepACO training. The values with * indicate that the values are used for fine-tuning a pre-trained model.
TSP200 TSP500 TSP1000 CVRP200 CVRP500 CVRP1000 OP100 PCTSP100 SOP100 SMTWTP100 BPP120
$N_{\text{epoch}}$ 50 50 20* 50 50 20* 20 20 20 20 20
$N_{\text{inst}}$ 400 400 200* 200 200 100* 100 100 100 100 100
$M$ 20 20 10 10 10 5 5 5 5 5 5
$K$ 30 30 20 20 20 15 30 30 30 30 30
There are other hyperparameters to be considered specifically for GFACS. $\beta$ is one of the most important hyperparameters for training GFACS, which is analyzed in Section 6.3. As we’ve discussed in Section 5.1, we use annealing for $\beta$ using the following log-shaped scheduling:
$\beta=\beta_{\text{min}}+(\beta_{\text{max}}-\beta_{\text{min}})\min\left(\frac{\log(i_{\text{epoch}})}{\log(N_{\text{epoch}}-N_{\text{flat}})},1\right)$
where $i_{\text{epoch}}$ is the index of epoch and $N_{\text{flat}}$ is the number of last epochs to have $\beta=\beta_{\text{max}}$.
We report the $\beta$ value used for each task in Table 5. Note that $\beta$ remains consistent across different scales of the same problem class.
Table 5: Settings of $\beta$ for each task.
TSP CVRP OP PCTSP SOP SMTWTP BPP
$\beta_{\text{min}}$ 200 500 5 20 10 10 500
$\beta_{\text{max}}$ 1000 2000 10 50 20 20 5000
For TSP and CVRP, where we have local search operators, we also need $\alpha$ for energy reshaping (Eq. 9). We empirically found that linearly increasing $\alpha$ from $0.5$ to $1.0$ along the epochs works well, and we use that setting for all TSP and CVRP experiments.
A.3 Computing resource
Throughout the experiments, we utilize a single CPU, specifically an AMD EPYC 7542 32-Core Processor, and a single GPU, the NVIDIA RTX A6000, for training neural networks. After thet, the inference was conducted on a separate, standalone PC with a single AMD Ryzen 9 5900X 12-Core Processor and NVIDIA RTX 4070 Ti to assess the wall time of each algorithm fairly.
A.4 Hyperparameters for LKH-3.
LKH-3 (Helsgaun, 2017), an extension of the Lin-Kernighan heuristic (Lin & Kernighan, 1973), is a powerful open-source heuristic solver designed for solving routing problems. It not only addresses TSP but also handles a wide array of constrained routing problems, such as CVRP and PDP, by minimizing a penalty term that measures constraint violations.
In the following, we describe the parameter settings for LKH-3 used in this work. For both TSP and CVRP, we set MAX_TRIALS to be $100\times N$ where $N$ is the number of nodes and SEED to $1234$. We set RUNS to 10 for TSP but 1 for CVRP to limit the search time in a reasonable range. For other parameters not stated, we followed the default settings.
Appendix B Problem Settings
In this section, we provide an overview of the combinatorial optimization problems we address. These encompass vehicle routing problems, bin packing problems, and scheduling problems. For a comprehensive understanding of each problem’s specific settings, we direct readers to the work by Ye et al. (2023). Furthermore, for access to the exact source code pertaining to each problem, please refer to the repository available at  [^2] https://github.com/henry-yeh/DeepACO from Ye et al. (2023).
Traveling Salesman Problem
In the Traveling Salesman Problem (TSP), we aim to find the shortest route that visits a set of nodes once and returns to the origin. We consider Euclidean TSP, where the distance between two nodes is determined by Euclidean distance. A TSP instance can be determined as a set of nodes with 2D coordinates. We generate a random TSP instance by sampling the coordinate from a unit square, $[0,1]^{2}$.
Capacitated Vehicle Routing Problem
Similar to TSP, Capacitated Vehicle Routing Problem (CVRP) also seeks the shortest path that visits all the nodes once. In CVRP, however, a vehicle has a limited capacity, and each node has a demand to be satisfied. The sum of the demand of each subroute should not exceed the capacity, which limits the number of nodes that can be visited by one vehicle. In our setting, CVRP is also defined in Euclidean space.
A CVRP instance can be represented by a set of customer nodes, a single depot node, and a capacity $C$. Each customer node has a position (2D coordinates) and a demand, while the depot only has a position. A random CVRP instance is generated by sampling the coordinates of both customer and depot from a unit square $[0,1]^{2}$ and sampling demands for each customer from a predefined uniform distribution, $U[a,b]$. We use $a=1$ and $b=9$, and a fixed $C=50$ for all problem scales, 200, 500 and 1000.
Orienteering Problem
In the Orienteering Problem (OP), the objective is to maximize the total prize of visited nodes within a given tour length limitation. This problem is similar to CVRP but focuses on maximizing gains rather than minimizing costs.
An OP instance comprises a set of nodes to visit, each with its position and prize, a depot, and a maximum length constraint. We sample a random OP instance by first sampling the position of nodes, including the starting node, from a unit square $[0,1]^{2}$. And we set the prize for each node, following (Fischetti et al., 1998; Ye et al., 2023), to be $p_{i}=(1+\lfloor 99\cdot\frac{d_{0,i}}{\max_{j=1}^{n}{d_{0,j}}}\rfloor)/100$, where $d_{0,i}$ is the Euclidean distance from the depot to node $i$. We set the maximum length to 4 in the OP100 task, following Kool et al. (2018).
Prize-Collecting Traveling Salesman Problem
The Prize-Collecting Traveling Salesman Problem (PCTSP) extends the TSP by associating a prize for visiting a node and a penalty for not visiting it. The goal is to find a tour that minimizes the total travel cost plus penalties for the unvisited nodes and also satisfies a predefined prize threshold. In PCTSP, balancing the need to collect prizes against the cost of visiting more nodes matters.
A PCTSP instance comprises a set of nodes to visit, a depot, and a minimum total prize threshold. The coordinates of nodes and depot are sampled in the unit square. The prize for each node is sampled from $U[0,1]$, while the penalty for each node is sampled from $U[0,C_{N}]$, where $C_{N}$ is set to $0.12$ for $N=100$. Lastly, the minimum total prize threshold is $N/4$ for problems with ${N}$ nodes, i.e., $25$ for PCTSP100.
Sequential Ordering Problem
The Sequential Ordering Problem (SOP) seeks to find a linear order of jobs that minimizes total processing time with a set of precedence constraints among the jobs. An SOP instance is a set of jobs and precedence constraints. Each job has its own processing time, and each pair has a waiting (transition) time.
We sample an SOP instance by sampling the processing time and waiting time from $U[0,1]$. The precedence constraints for each job pair are laid with probability 0.2, while ensuring feasibility, i.e., when $i<j$ means that the constraint that job $i$ precedes the job $j$, the feasibility of a problem requires $i<j$ and $j<k$ then $i<k$.
Single Machine Total Weighted Tardiness Problem
This problem involves scheduling jobs on a single machine where each job has a due date, processing time, and weight. The objective is to minimize the total weighted tardiness, which is the sum of the weights of the jobs multiplied by their tardiness (the amount of time a job is completed after its due date).
To generate an SMTWTP instance, we sample a due date from $U[0,N]$, the weight from $U[0,1]$, and the processing time from $U[0,2]$, where $N$ is the number of jobs to process.
Bin Packing Problem
The Bin Packing Problem (BPP) involves packing objects of different volumes into a finite number of bins or containers of fixed capacity to minimize the number of bins used. In this work, we use the latter setting.
To generate a BPP instance, we sample the $N$ items with a random size from $U[20,100]$. And we use 150 for the bin capacity.
Appendix C Additional Experiments
C.1 Experiments on TSPLIB
This section presents a comparison of performance between the baseline models and our model on real-world Traveling Salesman Problem (TSP) instances. We evaluated the models on TSP200, TSP500, and TSP1000 under their respective conditions. Our model outperforms the other Ant Colony Optimization (ACO) baselines, demonstrating superior generalization capabilities.
Table 6: Evaluation of ACO methods on the real-world TSP problem. ‘#’ denotes the number of instances in each set.
Size # ACO DeepACO Ours
$300100\leq N<$ 300 30 1.71% 1.25% 1.21%
$700300\leq N<$ 700 10 4.26% 2.70% 2.60%
$1500700\leq N<$ 1500 12 7.01% 4.08% 3.88%
ALL 52 3.43% 2.24% 2.16%
C.2 Comparison against TAM
We compare GFACS and DeepACO against CVRP instances used in TAM (Hou et al., 2022). These instances have a different capacity constraint from those we used for the main experiment. Please refer to (Hou et al., 2022) for more details regarding the instance generation. Moreover, to match the processing time to be comparative to DeepACO, we set $K$ (equal to the number of ants for ACO) to 16.
Table 7: Results for CVRP instances used in TAM. The result for AM and TAM is drawn from Hou et al. (2022), and the result for DeepACO is drawn from Ye et al. (2023).
CVRP100 CVRP400
Method Obj. Time (s) Obj. Time (s)
AM* 16.42 0.06 29.33 0.20
TAM* 16.08 0.086 25.93 1.35
DeepACO ($T=4$) 16.08 2.97 25.31 3.65
DeepACO ($T=10$) 15.77 3.87 25.27 5.89
GFACS ($T=4$) 15.77 1.07 24.95 3.23
GFACS ($T=10$) 15.72 1.40 24.87 5.21
Appendix D Diversity Calculation
The diversity of $K$ solutions generated from a solver is defined as the average pairwise Jaccard distance, $\frac{1}{K(K-1)}\sum_{i}\sum_{j}d_{J}(x_{i},x_{j})$, where $d_{J}$ is defined as follows:
$d_{J}(x_{i},x_{j})=1-J(x_{i},x_{j})=1-\cfrac{|E(x_{i})\cap E(x_{j})|}{|E(x_{i})\cup E(x_{j})|}$ (16)
where $E_{x}$ is the set of edges composing the solution $x$. In TSP or CVRP, the (undirected) edge is a set of adjoining nodes in the solution path.
Appendix E Symmetric Solutions of Combinatorial Optimization
This section delves into the symmetric characteristics of combinatorial solutions, wherein multiple trajectories, denoted as $\tau$, can map to an identical combinatorial solution. We begin with a formal analysis focusing on two quintessential tasks: the Traveling Salesman Problem (TSP) and the Capacitated Vehicle Routing Problem (CVRP). This analysis lays the foundation for understanding the connections to other related combinatorial optimization problems.
Proposition E.1 (Symmetry Solution in TSP (Hamiltonian Cycle)).
Let $\mathcal{G}=(\mathcal{V},\mathcal{D})$ be a finite, undirected graph where $\mathcal{V}$ is the set of vertices and $\mathcal{D}$ is the set of edges. Assume $\mathcal{G}$ contains a Hamiltonian cycle. Denote $N=|\mathcal{V}|$ as the number of vertices. If $N>1$, there exist exactly $2N$ symmetric trajectories (including the cycle and its reverse) starting from each vertex in the cycle. For $N=1$, there is exactly one identical trajectory.
Proof.
A Hamiltonian cycle in $\mathcal{G}$ is a cycle that visits every vertex exactly once. For $N>1$, each cycle can be traversed in two distinct directions (clockwise and counterclockwise) from each vertex, yielding $2$ trajectories per vertex. Since there are $N$ vertices, this results in $2N$ symmetric trajectories. For $N=1$, the cycle is trivial and only has one trajectory. ∎
Proposition E.2 (Symmetric Solutions in CVRP).
Consider a Capacitated Vehicle Routing Problem (CVRP) defined on a graph $\mathcal{G}=(\mathcal{V},\mathcal{D})$ with a designated depot vertex. Let there be $K$ distinct Hamiltonian cycles (or routes) originating and terminating at the depot, and denote $S$ as the number of single vertex cycles (excluding the depot) among these $K$ cycles. The total number of symmetric solutions, considering the directionality of each cycle, is given by $K!\times 2^{K-S}$.
Proof.
Each of the $K$ cycles can be ordered in $K!$ ways. Excluding the $S$ single vertex cycles, each of the remaining $K-S$ cycles can be traversed in two directions (forward and reverse). Therefore, for these $K-S$ cycles, there are $2^{K-S}$ ways to choose a direction. The product $K!\times 2^{K-S}$ gives the total number of symmetric solutions for the CVRP under these conditions. ∎
The propositions discussed previously can be extended to various variants of vehicle routing problems. Specifically, in the context of the orienteering problem (OP) and prize collecting Traveling Salesman Problem (PCTSP), which represent special cases where $K=1$, the number of symmetric trajectories is determined by the number of vertices, $N$, in the solution. When $N>1$, there are exactly 2 symmetric trajectories for each route. In contrast, for a single-vertex solution ($N=1$), the trajectory is identical, reflecting the singularity of the route.
Appendix F Extended Related Works
F.1 Deep Constructive Policy for Neural Combinatorial Optimization
Constructive policy methods systematically build solutions by incrementally adding elements, beginning with an empty set and culminating in a complete solution. This approach ensures compliance with constraints by progressively narrowing the scope of actions, making it versatile for application in domains like vehicle routing (Nazari et al., 2018), graph combinatorial optimization (Khalil et al., 2017), and scheduling (Zhang et al., 2020).
In deep learning, various techniques adopt this constructive policy framework. Notably, the attention model (AM) (Kool et al., 2018) has been pivotal in vehicle routing problems (VRPs), establishing a new standard for stabilizing REINFORCE training. Subsequent advancements in AM have included enhanced decoder architecture (Son et al., 2023a; Xin et al., 2021; Luo et al., 2023), improved REINFORCE (Williams, 1992) baseline assessments (Kwon et al., 2020; Kim et al., 2022b), leveraging symmetries for sample efficiency (Kim et al., 2023a), the application of population-based reinforcement learning (RL) (Grinsztajn et al., 2022), and refined solution-building methods (Drakulic et al., 2023). Recent advancements in hierarchical algorithms exploit AM-based policies for solution optimization: one policy broadly generates solutions while another refines them locally (Kim et al., 2021b; Hou et al., 2022; Ye et al., 2024). Additionally, AM-driven constructive policies have been effectively applied in hardware routing (Kim et al., 2020, 2021a) and placement tasks (Park et al., 2022; Kim et al., 2023b).
Graph combinatorial optimization, particularly in problems like maximum independent sets, has diversified from contrive VRPs policies due to their local decomposability and the relative ease of training value functions for partial solutions where VRPs are not locally decomposable due to their global contains. Starting with the influential works by Khalil et al. (2017), which employed deep Q networks (Mnih et al., 2013) for graph policy training, creating solutions constructively, Ahn et al. (2020) introduced an innovative Markov Decision Process (MDP) to generate multiple solution components per iteration by exploiting the local decomposability of graph combinatorial problems. Zhang et al. (2023a) further developed this concept, training their MDP with forward-looking GFlowNets (Pan et al., 2023a).
Our research aligns with the deep constructive policy framework, where GFlowNets constructively build actions from an initial empty solution ($s_{0}$) to a final complete solution ($s_{n}$). Additionally, our methods include a heatmap-based approach, detailed in Section F.3.
F.2 Deep Improvement Policy for Neural Combinatorial Optimization
The improvement policy focuses on refining complete solutions, similar to local search methods like 2opt (Croes, 1958). Its advantage lies in achieving comparable performances to constructive policies, especially when allotted sufficient time for extensive improvement iterations. Researchers typically adapt existing heuristic local search techniques, creating learning policies that emulate these methods through maximizing expected returns via reinforcement learning (d O Costa et al., 2020; Hottung & Tierney, 2019; Ma et al., 2021, 2022, 2023; Kim et al., 2022a). Also, recent works have explored active search to refine solutions via policy adjustments (Hottung et al., 2021; Son et al., 2023b). Moreover, alternative strategies using discrete Langevin sampling instead of neural networks have shown promising outcomes (Sun et al., 2023). This approach serves as a complementary extension to deep constructive policies, as it often enhances solutions initially formulated by constructive methods. Additionally, our method, which integrates local search into the training loop, could see further enhancements if advancements in this area are incorporated into our training procedures.
F.3 Heatmap-based Neural Combinatorial Optimization
The heatmap-based method (Joshi et al., 2021) entails the acquisition of a graph representation, referred to as the “Heatmap,” from input graph instances. This Heatmap serves as a prior distribution, influencing the generation of a solution. However, this approach is not self-contained and requires an additional solution construction step, which may involve techniques like beam search (Joshi et al., 2021), dynamic programming (Kool et al., 2022), or Monte Carlo tree search (Fu et al., 2021; Qiu et al., 2022; Sun & Yang, 2023). The key advantage of the heatmap-based method lies in its exceptional generalizability when compared to constructive policies. It operates at an abstract level, free from the constraints inherent in generating actual combinatorial solutions.
Our approach falls under heatmap-based methodologies, where we learn a heuristic metric (i.e., heatmap) and employ ant colony optimization for solution construction. A key advantage of our method over many other supervised learning-based heatmap approaches is its independence from optimally labeled data.

### v2

@@
 Abstract
-This paper introduces the Generative Flow Ant Colony Sampler (GFACS), a novel neural-guided meta-heuristic algorithm for combinatorial optimization.
-GFACS integrates generative flow networks (GFlowNets) with the ant colony optimization (ACO) methodology.
-GFlowNets, a generative model that learns a constructive policy in combinatorial spaces, enhance ACO by providing an informed prior distribution of decision variables conditioned on input graph instances.
-Furthermore, we introduce a novel combination of training tricks, including search-guided local exploration, energy normalization, and energy shaping to improve GFACS.
+This paper introduces the Generative Flow Ant Colony Sampler (GFACS), a neural-guided probabilistic search algorithm for solving combinatorial optimization (CO).
+GFACS integrates generative flow networks (GFlowNets), an emerging amortized inference method, with ant colony optimization (ACO), a promising probabilistic search algorithm.
+Specifically, we use GFlowNets to learn a constructive policy in combinatorial spaces for enhancing ACO by providing an informed prior distribution over decision variables conditioned on input graph instances.
+Furthermore, we introduce a novel off-policy training algorithm for scaling conditional GFlowNets into large-scale combinatorial spaces by leveraging local search and shared energy normalization.
 Our experimental results demonstrate that GFACS outperforms baseline ACO algorithms in seven CO tasks and is competitive with problem-specific heuristics for vehicle routing problems.
-The source code is available at https://github.com/ai4co/gfacs.
-GFlowNets, Ant Colony Optimization, Neural Combinatorial Optimization
 1 Introduction
-Combinatorial Optimization Problems (COPs) constitute a category of optimization problems focused on finding the optimal arrangement of discrete structures, such as sets, sequences, and graphs, that maximizes a given objective function.
-COPs are prevalent in various fields, including compiler scheduling (Zhang et al., 2022b) and hardware design (Kim et al., 2023b), but they are usually NP-hard.
-Over the decades, heuristics that quickly find near-optimal solutions have been developed to tackle this NP-hardness.
-However, these problem-specific heuristics require considerable effort and domain knowledge to make them effective.
-Thus, meta-heuristic algorithms, such as Ant Colony Optimization (ACO) (Dorigo et al., 2006), have been proposed to provide a more generalized approach to designing heuristics.
-ACO is a meta-heuristic algorithm inspired by the foraging behaviors of ants, and it has demonstrated effectiveness in solving COPs and general optimization problems.
-In ACO, a population of artificial ants iteratively builds solutions and updates ‘pheromone trails’ to explore the solution space.
-At each iteration, solutions are constructed through stochastic moves guided by instance-specific pheromone trails and problem-specific heuristic measures indicating the goodness of sampling certain trails; see Figure 1 for an illustration.
-While the pheromone trails are updated based on solutions constructed during the iterative procedure of solving an instance, the heuristic measures typically need to be predefined in a matrix form (e.g., heat map).
-However, developing an effective heuristic matrix for ACO involves expert knowledge and manual tuning, posing challenges in less-explored problem domains where the expertise is limited.
-Figure 1: Illustration of ACO.
-Edge probability is obtained from the multiplication of heuristics and pheromone trails.
-Ants constructively complete solutions based on the edge probability, then update the pheromone trail with the constructed samples.
-Recently, a neural-guided Ant Colony Optimization (ACO) method, DeepACO, was introduced by Ye et al.
-(2023).
-This technique utilizes graph neural networks (GNN) (Joshi et al., 2021) to model the heuristic matrix of input instance graphs of COPs.
-Specifically, this framework is developed as a GNN-based policy that selects an edge (action) based on previously chosen edges (state).
-Then, the composited policy generates a state trajectory from an empty set (initial state) to a complete solution (terminal state), as in Figure 1.
-DeepACO incorporates reinforcement learning (RL) to train the GNN-based policy model with a reward signal from the terminal state of the generated trajectory.
-After RL training, the learned GNN is utilized to construct the heuristic matrix for a given problem instance, thereby providing a good prior to ACO to explore the solution space more efficiently.
-Note that the trained GNN acts as a parameterized prior, effectively transferring the solution exploration strategy obtained during training to each new instance.
-However, we observe certain limitations in DeepACO, primarily in three aspects.
-The first limitation ( ) is that the usage of a behavior policy is limited since DeepACO uses on-policy RL to train the graph representation.
-It potentially weakens the performance of heuristic matrices by limiting the choice of behavior policy, which is crucial for the effective exploration of massive combinatorial spaces.
-Secondly ( ), the reward-maximizing RL of DeepACO fails to satisfactorily capture the diversity of solutions, i.e., the multi-modality of the solution landscape.
-In CO, a diversity of good candidates is highly advantageous, as it increases the chances of finding a global optimum after a local search; see Figure 2.
-Lastly ( ), the inherent symmetry in combinatorial solutions, where multiple trajectories can lead to identical combinatorial outcomes, is challenging to consider in the context of RL.
-For example, in the Traveling Salesman Problem (TSP), distinct paths like (1,2,3,4) and (4,3,2,1) represent the same cyclic solution.
-As DRL methods train policy on a tree-structured MDP, each trajectory is identified as a different solution.
-Figure 2: Importance of capturing multi-modality in combinatorial optimization with local search post-processing ( ).
-The probability of improved solutions rises after local search, even from initially higher-cost samples, particularly with a sampler adept at capturing multi-modality.
+Combinatorial optimization (CO) is a crucial domain involving high-impact applications, including vehicle routing, hardware design optimization, and biological sequence design.
+Ant colony optimization (ACO) [14, 68] is a general-purpose meta-heuristic algorithm designed to address these tasks.
+ACO employs a constructive solution generation scheme that sequentially adds elements to a partial solution, known for its flexibility across CO tasks.
+This approach involves collecting experiences from multiple agents (ants) to iteratively update a distribution (pheromone update) based on a prior distribution known as the heuristic matrix.
+The major benefits of ACO include: (1) its robustness against local optima, enabled by the algorithm’s stochastic nature, which dynamically balances exploration and exploitation, preventing premature convergence [14]; and (2) its flexible solution generation scheme, which allows it to effectively solve a wide range of combinatorial optimization tasks [79].
+The main challenge with ACO lies in designing effective prior distributions (i.e., the heuristic matrix) for their probability models.
+Inadequate priors can lead to time-consuming pheromone update iterations.
+Recent advancements, known as DeepACO [79], address this by using policy gradient reinforcement learning (RL) [76, 71] to pretrain the priors conditioned on given problem instances.
+RL trains these networks to guide highly rewarded regions, allowing generalization to new problems and enhanced performance.
+Motivated by DeepACO, we propose a deep learning-based prior distribution design.
+However, this paper tackles challenges related to the design of prior distribution through deep amortized inference rather than the policy gradient RL method.
+This is motivated by the following three primary limitations in DeepACO’s use of policy gradient RL.
+First ( ), the on-policy nature of DeepACO restricts the flexible use of a behavior policy, potentially undermining performance by limiting effective exploration of vast combinatorial spaces.
+Second ( ), the reward maximizing form of RL in DeepACO struggles to capture the diverse solutions necessary in CO, even though having a variety of good candidates could significantly enhance the likelihood of finding a global optimum after local searches.
+Last ( ), the inherent symmetry in combinatorial solutions—where different paths can lead to identical outcomes—presents a significant challenge not adequately addressed in DeepACO.
 Contribution.
-This paper introduces a novel ACO algorithm, the Generative Flow Ant Colony Sampler (GFACS), which addresses the limitations of DeepACO by integrating the emerging amortized inference method of GFlowNets (Bengio et al., 2021, 2023).
-First ( ), the off-policy training property of GFlowNets allows for the utilization of samples drawn from any behavior policy, including powerful heuristic search algorithms commonly used in COPs.
-Secondly ( ), unlike maximum reward-seeking algorithms (e.g., RL training in DeepACO), GFlowNets are sampling algorithms and well-suited for modeling the multi-modality nature of certain distributions.
-Lastly ( ), GFlowNets, which model the solution generative process on a directed acyclic graph (DAG), incorporate the symmetricity of combinatorial solutions by design (Bengio et al., 2023).
-To adequately integrate GFlowNet into ACO, we propose an effective combination of training techniques.
-In our approach, we employ a strategic guided exploration with two distinct behavior policies to collect exploratory and exploitation samples.
-One aims for a diverse range of samples, and the other aims for locally optimized samples.
-This approach effectively addresses the over-exploration issue in GFlowNets, which often results in poor performances due to an excessive focus on diversity.
-To encourage further exploration, we introduce an energy reshaping to compensate samples that have initially high energy but can be significantly improved, as exemplified by the red samples in Figure 2.
-At last, a shared energy normalization is employed to stabilize the training of conditional GFlowNets, wherein the energy landscape is contingent upon the variability of input problems.
-Our GFACS exhibits significant superiority over the current DeepACO and classical ACO algorithms for various representative CO tasks, which vehicle routing problems, scheduling problems, and a grouping problem.
-Moreover, GFACS outperforms or shows similar performance to competitive deep reinforcement learning-based vehicle routing solvers, which are designed to solve traveling salesman problems (TSP) and capacitated vehicle routing problems (CVRP) in the context of large-scale instances, such as $N=500,1000$.
+This paper introduces a novel ACO algorithm, the Generative Flow Ant Colony Sampler (GFACS), which addresses the limitations of DeepACO by integrating the emerging amortized inference method of GFlowNets [4, 5].
+Our twofold contribution can be summarized as follows.
+(1) We introduce a new viewpoint to ACO using amortized inference instead of policy gradient RL.
+(2) Additionally, we present a novel GFlowNet training algorithm for large scale combinatorial tasks, applicable to a range of further uses over ACO.
+  • ACO as off-policy amortized inference with GFlowNets.
+There are three reasons that GFlowNets can improve over the existing DeepACO algorithm.
+First ( ), the off-policy training property of GFlowNets allows for the utilization of samples drawn from any behavior policy, including powerful heuristic search algorithms commonly used in CO problems.
+Secondly ( ), unlike maximum reward-seeking algorithms (e.g., RL training in DeepACO), GFlowNets are sampling algorithms that are well-suited for modeling the multimodal nature of certain distributions.
+Lastly ( ), GFlowNets model solutions as a generative process on a directed acyclic graph (DAG), which can represent multiple trajectories that share identical states.
+This approach allows for accurate probabilistic modeling by considering the symmetry of combinatorial solutions [5].
+  • Scaling conditional GFlowNets into large-scale combinatorial space.
+GFlowNets often struggle with limited scalability due to challenging exploration problems [58, 39], credit assignment issues [59, 26], and unstable training when inferring conditional distributions [82].
+In combinatorial optimization, both limitations must be addressed because we deal with (1) large-scale combinatorial spaces and (2) conditional inference on problem instances.
+This study introduces a new off-policy training method for GFlowNets in large combinatorial spaces by effectively leveraging local search operators.
+It also proposes a novel training method to address instability in training conditional GFlowNets by properly normalizing conditional reward values with sample mean.
+In comparative experiments, our GFACS exhibits significant superiority over the current DeepACO and classical ACO algorithms for various representative CO tasks – vehicle routing problems, scheduling problems, and a grouping problem.
+Moreover, GFACS outperforms or shows similar performance to competitive unsupervised learning-based vehicle routing solvers, which are designed to solve traveling salesman problems (TSP) and capacitated vehicle routing problems (CVRP) in the context of large-scale instances, such as $N=500,1000$.
+Lastly, we compared our approach with previous GFlowNet methods for solving combinatorial optimization problems [82, 84] in vehicle routing tasks, which are known to pose challenges for credit assignment due to their locally undecomposable properties, as highlighted by Ahn et al.
+[1].
 2 Related works
@@
 Ant Colony Optimization (ACO) is a meta-heuristic inspired by ant behavior in finding optimal paths.
-Traditional ACO algorithms, such as Ant System (AS) (Dorigo et al., 1996) and MAX-MIN AS (Stützle & Hoos, 2000), rely on specific heuristics for success.
+Traditional ACO algorithms, such as Ant System (AS) [14] and MAX-MIN AS [68], rely on specific heuristics for success.
 To reduce the reliance on heuristic design, Ye et al.
-(2023) introduce a neural-guided ACO, called DeepACO, utilizing graph representation learning and reinforcement learning to replace manual heuristics with automatically acquired knowledge.
+[79] introduce a neural-guided ACO, called DeepACO, utilizing graph representation learning and reinforcement learning to replace manual heuristics with automatically acquired knowledge.
 2.2 Neural combinatorial optimization
-Neural combinatorial optimization (Bello et al., 2016) is a deep learning-based method to train an instance-conditional policy using supervised learning (Vinyals et al., 2015; Fu et al., 2021; Sun & Yang, 2023) or unsupervised learning (Khalil et al., 2017; Kool et al., 2018; Kwon et al., 2020).
+Neural combinatorial optimization [3] is a deep learning-based method to train an instance-conditional policy using supervised learning [75, 20, 70] or unsupervised learning [30, 41, 43, 54].
 Usually, supervised learning methods are beneficial when the target problem has an accurate but time-consuming expert solver.
-The oracle-labeled data is collected from the oracle solvers, and deep neural networks amortize them, allowing for scalability and faster inferences than the expert solver.
+The oracle-labeled data is collected from the oracle solvers, and deep neural networks amortize the search process, converting an expensive run-time optimization into an also expensive but one-time cost of training the neural net, allowing for scalability and faster run-time inferences than the expert solver.
 On the other hand, unsupervised learning proceeds through self-guided trial and error, using the COP objective as a training signal.
-This makes it applicable even if encountering less-studied problems where powerful expert solvers are unavailable (Bello et al., 2016; Ahn et al., 2020).
+This makes it applicable even if encountering less-studied problems where powerful expert solvers are unavailable [3, 1].
 Our method can be categorized as an unsupervised learning method, which does not need expert labels.
@@
 2.3 GFlowNets and combinatorial optimization
-GFlowNets, tailored for sampling combinatorial objects proportionally to a given reward function, play a pivotal role in diversifying the candidate pool for molecule generation in crucial areas like drug discovery (Bengio et al., 2021).
-Recent efforts to advance GFlowNets have focused on developing new training objectives (Malkin et al., 2022; Madan et al., 2022; Zhang et al., 2023b), enhancing credit assignment (Pan et al., 2023a; Jang et al., 2024), and improving exploration strategies (Pan et al., 2022; Rector-Brooks et al., 2023; Pan et al., 2023a; Kim et al., 2023d, c).
-Our contribution lies in proposing an efficient exploration technique tailored for combinatorial optimization problems.
-GFlowNets have also found successful applications in some COPs (Zhang et al., 2022b, 2023a).
+GFlowNets are designed for sampling combinatorial objects proportionally to a given reward function [4].
+Theoretical studies have connected GFlowNets with variational inference [53] and maximum entropy reinforcement learning [56, 13].
+Recent algorithmic advancements in GFlowNets have focused on developing new training objectives [52, 51, 85], enhancing credit assignment [59, 26], and improving exploration strategies [58, 63, 59, 39, 38].
+GFlowNets have also found successful applications in some combinatorial optimization problems (COPs) [82, 84].
 Zhang et al.
-(2023a) effectively addresses the symmetric nature of graph constraint optimization, showcasing remarkable performance in graph CO tasks like maximum independent sets.
-Zhang et al.
-(2022b) solves proxy-based scheduling problems.
-Despite their success, their method has limited applications to either locally decomposable graph CO problems (Ahn et al., 2020) or robust scheduling problems.
-Our method is a neural-guided meta-heuristic, which aims to be a general-purpose CO solver.
-Additionally, this paper appears to be the first to address the traveling salesman problem and its variations using GFlowNets.
-These tasks present a significant challenge due to the global constraints to compare with graph CO tasks.
+[84] effectively address the symmetric nature of graph constraint optimization, demonstrating remarkable performance in tasks such as finding maximum independent sets.
+Despite their success, there are limitations in expanding to other combinatorial tasks due to their reliance on local credits for the learning objective proposed by Pan et al.
+[59].
+Conversely, Zhang et al.
+[82] address proxy-based scheduling problems by developing a problem-specific neural architecture and introducing novel training methods for general-purpose conditional GFlowNets.
+One of their significant contributions is the proposal of a loss function based on the sample variance of the log partition function, known as VarGrad, which enhances the training process and performance.
 3 Preliminary
-3.1 Travelling Salesman Problem
+3.1 Traveling salesman problem
 The traveling salesman problem (TSP) is a fundamental combinatorial optimization problem, which serves as a building block for more complex CO problems like the capacitated vehicle routing problem (CVRP) and orienteering problem (OP).
 In light of this and for simplicity, this paper presents formulations using the two-dimensional Euclidean TSP as an illustrative example.
-Note that the formulations can be extended to other COPs straightforwardly.
 A TSP instance is defined on a (fully connected) graph $\mathcal{G}=\{\mathcal{V},\mathcal{D}\}$, where $\mathcal{V}$ and $\mathcal{D}$ are a node set and edge set, respectively.
@@
 A solution to TSP, $x(\mathcal{G})$ corresponds to a Hamiltonian cycle for a given problem instance $\mathcal{G}$, which can be represented as a permutation of nodes $\pi=(\pi_{1},\ldots,\pi_{N})$.
-The tour length is computed as $f(x(\mathcal{G}))=\sum_{t=1}^{N-1}d_{\pi_{t},\pi_{t+1}}+d_{\pi_{N},\pi_{1}}$.
-The optimization objective for TSP is to find $x(\mathcal{G})$ that minimizes $f(\cdot,\mathcal{G})$.
+The tour length is computed as $\mathcal{E}(x(\mathcal{G}))=d_{\pi_{N},\pi_{1}}+\sum_{t=1}^{N-1}d_{\pi_{t},\pi_{t+1}}$.
+The optimization objective for TSP is to find $x(\mathcal{G})$ that minimizes $\mathcal{E}(x(\mathcal{G}))$.
 3.2 Ant colony optimization
@@
 The pheromone update rules vary depending on the ACO algorithm used.
-For example, in Ant System (AS) (Dorigo et al., 1996), all $K$ solutions contribute to the pheromone update:
+For example, in Ant System (AS) [14], all $K$ solutions contribute to the pheromone update:
 $\rho_{i,j}\leftarrow(1-\gamma)\rho_{i,j}+\sum\limits_{k=1}^{K}\Delta^{k}_{i,j},$ (2)
@@
 Here, $\Delta^{k}_{i,j}$ is proportional to the inverse of tour length if $(i,j)$ is included in $\pi^{k}$, and zero otherwise.
-Note that the heuristic matrix $\eta(\mathcal{G})\in\mathbb{R}^{N\times N}$ is not updated during the search process, only serving as a prior.
-Thus, ACO aligns the search process more closely with the most efficacious paths discovered; see Dorigo et al.
-(2006) for details.
-In DeepACO (Ye et al., 2023), the learned GNN, parameterized by $\theta$, is utilized to construct the heuristic matrix $\eta(\mathcal{G},\theta)$ for a given problem instance $\mathcal{G}$, thereby providing a good prior for exploring the solution space.
-Figure 3: Solution sampling process of GFACS. The GNN is trained with GFlowNets and serves as an expert heuristic in GFACS.
+$\Delta^{k}_{i,j}=\begin{cases}C/f(\pi^{k})&\text{if $(i,j)\in\pi^{k}$}\\ 0&\text{otherwise}\end{cases}$ (3)
+where $C$ is a constant. The smaller the cost $f(\pi^{k})$, the more pheromone is laid on the edges of $\pi^{k}$.
+Note that the heuristic matrix $\eta(\mathcal{G})\in\mathbb{R}^{N\times N}$ is not updated during the search process; it only serves as a prior.
+ACO aligns the search process more closely with the most efficacious paths discovered; see Dorigo et al.
+[15] for details.
+In DeepACO [79], the learned graph neural network (GNN), parameterized by $\theta$, is utilized to construct the heuristic matrix $\eta(\mathcal{G},\theta)$ for a given problem instance $\mathcal{G}$, thereby providing a good prior for guiding probabilistic solution search process of ACO.
+Figure 1: Solution sampling process of GFACS. The GNN, pretraind with a GFlowNet loss, serves as an expert heuristic to pick the next action to construct a solution, such as a tour in the TSP.
 3.3 Generative flow networks
-Generative Flow Networks, or GFlowNets (Bengio et al., 2021, 2023) provide an effective training framework for learning a policy that samples compositional objects denoted as $x\in\mathcal{X}$.
+Generative Flow Networks, or GFlowNets [4, 5] provide an effective training framework for learning a policy that samples compositional objects denoted as $x\in\mathcal{X}$ towards the objective of generating $x$ with probability proportional to a given non-negative reward function $R(x)$.
 GFlowNets follow a constructive generative process, using discrete actions to iteratively modify a state which represents a partially constructed object.
@@
 Similarly, we also consider the backward policy $P_{B}(s|s^{\prime})$ for the backward transition $s^{\prime}\dashrightarrow s$, where $s$ is a parent of $s^{\prime}$.
-We can use the forward policy to compute in a forward way the probability of a trajectory:
-$P_{F}(\tau;\theta)=\prod_{t=1}^{n}P_{F}\left(s_{t}|s_{t-1};\theta\right)$ (3)
-Similarly, we can use the backward policy to compute in a backward way the probability of trajectory ending at a given object $x$:
-$P_{B}(\tau_{\rightarrow x}|x;\theta)=\prod_{t=1}^{n}P_{B}\left(s_{t-1}|s_{t};\theta\right)$ (4)
-Trajectory balance (Malkin et al., 2022).
-Trajectory balance is the most widely used learning objective in a GFlowNet framework.
+We can use the forward policy to compute in a forward way the probability of a trajectory:$P_{F}(\tau;\theta)=\prod_{t=1}^{n}P_{F}\left(s_{t}|s_{t-1};\theta\right)$.
+Similarly, we can use the backward policy to compute in a backward way the probability of trajectory ending at a given object $x=s_{n}$: $P_{B}(\tau_{\rightarrow x}|x;\theta)=\prod_{t=1}^{n}P_{B}\left(s_{t-1}|s_{t};\theta\right)$.
+Trajectory balance [52].
+Trajectory balance is the most widely used training objective for GFlowNets.
 When the trajectory balance objective is minimized, we achieve that the probability of sampling $x$ with the forward policy is $P_{F}^{\top}(x)\propto R(x)$, as desired, where $P_{F}^{\top}(x)=\sum_{\tau_{\rightarrow x}\in\mathcal{T}}P_{F}(\tau)$.
 The trajectory balance loss $\mathcal{L}_{\text{TB}}$ works by training three models, a learnable scalar $Z_{\theta}$ for the partition function, a forward policy $P_{F}(\tau;\theta)$, and a backward policy $P_{B}(\tau_{\rightarrow x}|x;\theta)$ to minimize the following objective:
-$\mathcal{L}_{\text{TB}}(\tau;\theta)=\left(\log\frac{Z_{\theta}P_{F}(\tau;\theta)}{R(x)P_{B}(\tau_{\rightarrow x}|x;\theta)}\right)^{2}$ (5)
-4 Generative Flow Ant Colony Sampler
+$\mathcal{L}_{\text{TB}}(\tau;\theta)=\left(\log\frac{Z_{\theta}P_{F}(\tau;\theta)}{R(x)P_{B}(\tau_{\rightarrow x}|x;\theta)}\right)^{2}.$ (4)
+4 Generative flow ant colony sampler (GFACS)
+In this section, we provide the inference procedure for the Generative Flow Ant Colony Sampler (GFACS).
+GFACS is a sampling algorithm that enhances Ant Colony Optimization (ACO) by integrating it with GFlowNets.
+The central concept is to use GFlowNets to generate an appropriate instance-conditioned prior (edge) distribution (heuristic matrix), which more effectively guides the probabilistic search procedure of ACO.
 4.1 MDP formulation
-This subsection formulates a Markov decision process (MDP) by using TSP as an example, similar to the preliminaries.
-A TSP solution is obtained through the sequential decision-making process of GFlowNets as follows:
-State $s$:
-The initial state $s_{0}$ is an empty set. The state $s_{t}$ at $t>0$ is a sequence of previous actions: $(a_{0},\ldots,a_{t-1})$.
-Action $a$:
-The action $a_{t}$ is a selection of a node to visit at time $t$ among un-visited nodes: $a_{t}\in\{1,\ldots,N\}\setminus\{a_{0},\ldots,a_{t-1}\}$.
+The solution construction procedure by GFlowNets can be modeled as a Markov decision process (MDP).
+We illustrate this procedure using the Traveling Salesman Problem (TSP) as an example.
+This formulation can be easily adapted to general combinatorial optimization problem with minor modifications.
+State $s$: The initial state $s_{0}$ is an empty set.
+The state $s_{t}$ at $0<t<n$ is a partial combinatorial solution of prensented with previous actions: $\{a_{0},\ldots,a_{t-1}\}$ and $s_{n}$ is completed solution.
+Action $a$: The action $a_{t}$ is a selection of a node to visit at time $t$ among un-visited nodes: $a_{t}\in\{1,\ldots,N\}\setminus\{a_{0},\ldots,a_{t-1}\}$.
 If all nodes are visited, a deterministic terminating action is applied, which maps a final state $s_{N}$ into a solution, $x(\mathcal{G})$.
-Reward $R(x(\mathcal{G});\beta)$:
-The reward is defined as the unnormalized energy distribution of the solution $x(\mathcal{G})$, i.e.,
-$R(x(\mathcal{G});\beta)=e^{-\beta\mathcal{E}(x(\mathcal{G}))}=e^{-\beta f(x(\mathcal{G}))}.$ (6)
-Here, $\mathcal{E}(x(\mathcal{G}))$ is energy for solution $x(\mathcal{G})$, which is the same as the tour length of the solution for TSP, i.e., $\mathcal{E}(x(\mathcal{G}))=f(x(\mathcal{G}))$.
-The $\beta$ is the inverse energy temperature, which is a hyperparameter.
+Reward $R(x(\mathcal{G});\beta)$: The reward is defined as the unnormalized Boltzmann distribution of the solution $x(\mathcal{G})$ with energy $\mathcal{E}(x(\mathcal{G}))$.
+Specifically, the reward is given by $R(x(\mathcal{G});\beta)=e^{-\beta\mathcal{E}(x(\mathcal{G}))}$.
+Here, $\mathcal{E}(x(\mathcal{G}))$ represents the energy of the solution $x(\mathcal{G})$, which corresponds to metrics such as the tour length in the TSP.
+The coefficient $\beta$, known as the inverse temperature, is a hyperparameter.
 4.2 Policy configuration
-We parameterize the forward policy $P_{F}$ using a graph neural network (GNN) with parameter $\theta$ that maps an input graph instance $\mathcal{G}$ to the output edge feature $\eta(\mathcal{G};\theta)$, following Ye et al.
-(2023).
+We parameterize the forward policy $P_{F}$ using a graph neural network (GNN) with parameter $\theta$ that maps an input graph instance $\mathcal{G}$ to the output edge matrix $\eta(\mathcal{G};\theta)$, following Ye et al.
+[79].
 Then, the forward policy when $t<N$ is expressed as follows:
-$\begin{split}&P_{F}(s_{t+1}|s_{t};\eta(\mathcal{G};\theta),\rho)\\ &\propto\begin{cases}\rho_{a_{t-1},a_{t}}\eta_{a_{t-1},a_{t}}(\mathcal{G};\theta)&\text{if $a_{t}\notin s_{t}$ }\\ 0&\text{otherwise}\end{cases}\end{split}$ (7)
+$P_{F}(s_{t+1}|s_{t};\eta(\mathcal{G};\theta),\rho)\propto\begin{cases}\rho_{a_{t-1},a_{t}}\eta_{a_{t-1},a_{t}}(\mathcal{G};\theta)&\text{if $a_{t}\notin s_{t}$ }\\ 0&\text{otherwise}\end{cases}$ (5)
 Note that $\rho$ and $\eta(\mathcal{G};\theta)$ denote the pheromone trails and learned heuristic matrix in ACO.
-Subsequently, we can define a trajectory from the initial state to a terminal solution $\tau=(s_{0},\ldots,s_{N},x(\mathcal{G}))$ and the forward probability of that trajectory:
-$P_{F}(\tau;\eta(\mathcal{G};\theta),\rho)=\prod_{t=0}^{N-1}P_{F}(s_{t+1}|s_{t};\eta(\mathcal{G};\theta),\rho)$ (8)
+Subsequently, we can define a trajectory from the initial state to a terminal solution, with $\tau=(s_{0},\ldots,s_{N},x(\mathcal{G}))$, and the forward probability of that trajectory,
+$P_{F}(\tau;\eta(\mathcal{G};\theta),\rho)=\prod_{t=0}^{N-1}P_{F}(s_{t+1}|s_{t};\eta(\mathcal{G};\theta),\rho).$ (6)
 Note that the forward probability for the transition from $s_{N}$ to $x(\mathcal{G})$ is omitted as it follows a deterministic rule.
-For the backward policy, we set $P_{B}=\frac{1}{2N}$, emulating the maximum entropy GFlowNets (Zhang et al., 2022a) to make uniform exploration over symmetric trajectories which terminate into identical solution.
+For the backward policy, we set $P_{B}=\frac{1}{2N}$, emulating the maximum entropy GFlowNets [83] to make uniform exploration over symmetric trajectories which terminate into identical solution.
 Here, $P_{B}$ represents a uniform distribution across all $2N$ feasible trajectories leading to a solution $x(\mathcal{G})$.
@@
 The overall process is illustrated in Figure 1.
-5 Training GFACS
-This section introduces how to train GFACS, especially the GNN component $\theta$, in the GFlowNet framework.
-GFACS is trained via iterations of the following two steps.
-First (Step A; Section 5.1), we collect experiences, each being a triplet of input instance graph $\mathcal{G}$, a complete trajectory $\tau_{\rightarrow x(\mathcal{G})}$, and the energy of the solution $\mathcal{E}(x(\mathcal{G}))$.
-We employ $\beta$ annealing and guided exploration for better exploration and energy reshaping for proper credit assignment.
-Next (Step B; Section 5.1), the parameterized forward policy $P_{F}$ is trained using the collected experiences.
-Here, we utilize a shared energy normalization to stabilize training.
-5.1 Step A: Experience Collection
+5 Off-policy training of GFACS
+This section introduces how to train GFACS, especially the GNN component parameterized with $\theta$, in the GFlowNet framework.
+GFACS is trained via iterations of the off-policy experience collection (Step A) and training (Step B).
+5.1 Step A: Off-policy experience collection with local search and energy reshaping
+Figure 2: Hindsight observation with local search and energy reshaping to control the trade-off between exploration and exploitation.
+In this step, we collect experiences on given graph instances.
+The experiences in step A will be used for training GFACS in step B.
 For every iteration, we first sample $M$ problem instances from a random sampler $P(\mathcal{G})$, i.e., $\mathcal{G}_{1},\ldots,\mathcal{G}_{M}\sim P(\mathcal{G})$.
 Details of instance sampling are described in Appendix B.
-Collecting exploration samples with $\beta$ annealing.
-For each problem instance $\mathcal{G}_{m}$, we collect $K$ exploration solutions by following Eq.
-8, resulting in an exploration batch, $\mathcal{B}_{\text{explore}}=\{(\mathcal{G}_{m},\tau^{m,k}_{\rightarrow x_{m,k}},\mathcal{E}(x_{m,k})\}_{m=1,\ldots,M,k=1,\ldots,K}$.
-To encourage exploration for solution sampling, we introduce an annealing scheme for the inverse energy temperature $\beta$.
-Rather than set the $\beta$ to a fixed value, we gradually increase the beta toward the target value along the training procedure.
-This makes the policy in the early stage of training to be exploratory, preventing it from getting stuck into a local optimum.
-Figure 4: Training of GFACS using GFlowNets and our novel training tricks of guided exploration and energy reshaping.
-Collecting exploitation samples with guided exploration.
-To cope with the over-exploration problem of GFlowNet, we also collect $K$ exploitation samples for each $\mathcal{G}_{m}$ using guided exploration.
-We use a heuristic local search operator ($LS$), such as 2-opt for TSP, as external guidance.
+For each problem instance $\mathcal{G}_{m}$, we collect $K$ solutions by following Eq.
+6 as $\mathcal{B}_{\text{explore}}=\{(\mathcal{G}_{m},\tau^{m,k}_{\rightarrow x_{m,k}},\mathcal{E}(x_{m,k})\}_{m=1,\ldots,M,k=1,\ldots,K}$.
+To exploit high-reward regions in the large-scale combinatorial space, we also collect $K$ exploitation samples for each $\mathcal{G}_{m}$ using a local search operator ($LS$), such as 2-opt for TSP, as a powerful behavioral policy.
 The local search operator directly refines a given solution into a locally optimized one, i.e., $x^{\prime}\leftarrow LS(x)$.
-We then use $P_{B}$ to collect a trajectory $\tau_{\rightarrow x^{\prime}}$ for each $x^{\prime}$, obtaining a batch of exploitation experiences, $\mathcal{B}_{\text{exploit}}=\{(\mathcal{G}_{m},\tau^{m,k}_{\rightarrow x^{\prime}_{m,k}},\mathcal{E}(x^{\prime}_{m,k})\}_{m=1,\ldots,M,k=1,\ldots,K}$.
-However, a powerful local search operator may not exist for some problems, especially for less-explored ones.
-In such cases, we use the Top-$L$ guidance technique, described as follows.
-For a given problem instance $\mathcal{G}_{m}$, we first update the pheromone trails using the exploration samples collected before.
-After that, $L\times K$ new sample solutions are generated with the updated pheromone, where $L$ is a hyperparameter.
-Among the new samples, we select top $L$ solutions with the lowest energy and also select $(K-L)$ random solutions, resulting in $K$ exploitation samples for each $\mathcal{G}_{m}$, and thus obtaining $B_{\text{exploit}}$ like before.
-Energy reshaping for exploratory samples.
-In the case where we have a local search operator, we reshape the energy for the exploration $x$ in $\mathcal{B}_{\text{explore}}$ using the energy of the refined solution $x^{\prime}=LS(x)$ as follows:
-$\mathcal{E}_{\text{reshaped}}(x)=\alpha\mathcal{E}(x^{\prime})+(1-\alpha)\mathcal{E}(x)$ (9)
-This is to evaluate an underrated sample that initially had high energy (i.e., low reward) but can be easily changed to have low energy (i.e., high reward) after the local search (see Figure 2).
-This reshaping changes the reward values in $\mathcal{B}_{\text{explore}}$ accordingly, while $\mathcal{B}_{\text{exploit}}$ remains the same.
-Note that this reshaping cannot be applied for TopK guidance because $x^{\prime}$ cannot be determined solely with $x$ in that case.
-5.2 Step B: Training with Collected Experiences
+We then use the backward policy $P_{B}$ to sample a trajectory $\tau_{\rightarrow x^{\prime}}$ for each $x^{\prime}$, obtaining a batch of “exploitation experiences,” $\mathcal{B}_{\text{exploit}}=\{(\mathcal{G}_{m},\tau^{m,k}_{\rightarrow x^{\prime}_{m,k}},\mathcal{E}(x^{\prime}_{m,k})\}_{m=1,\ldots,M,k=1,\ldots,K}$.
+Please note that we also suggest a general-purpose local search operator that partially destroys a solution and then repairs for problems that lack domain-specific local search methods.
+Please refer to Appendix C for details about the destroy-and-repair local saerch.
+While the concept of using local search for exploitation has been explored in recent GFlowNet works [39], our distinct approach involves utilizing the local search (LS) operation to facilitate hindsight observation—a forward-looking insight derived from extending the search beyond the current state.
+We focus on evaluating samples that were initially overlooked due to their high energy (i.e., low reward) but can be transformed to exhibit low energy (i.e., high reward) following the local search.
+To achieve this, we modify the energy (in a way we call energy reshaping) for the exploration around $x$ in $\mathcal{B}_{\text{explore}}$ by using the energy of the refined solution $x^{\prime}=LS(x)$ as follows:
+$\mathcal{E}_{\text{reshaped}}(x)=\alpha\mathcal{E}(x^{\prime})+(1-\alpha)\mathcal{E}(x).$ (7)
+5.2 Step B: Training with collected experiences with shared energy normalization
 Given the collected experience batches $\mathcal{B}_{\text{explore}}$ and $\mathcal{B}_{\text{exploit}}$, we aim to minimize the following loss function:
 $\mathcal{L}(\theta)=\mathcal{L}(\theta;\mathcal{B}_{\text{explore}})+\mathcal{L}(\theta;\mathcal{B}_{\text{exploit}}),$ (8)
-where ($\begin{split}&\mbox{where }\mathcal{L}(\theta;\mathcal{B})\\ &=\sum_{m=1}^{M}\sum_{k=1}^{K}\left(\log\frac{2NP_{F}(\tau^{m,k}_{\rightarrow x_{m,k}};\eta(\mathcal{G}_{m};\theta),\rho)Z_{\theta}(\mathcal{G}_{m})}{R(x_{m,k};\beta)}\right)^{2}.\end{split}$ (11)
+where ($\mbox{where }\mathcal{L}(\theta;\mathcal{B})=\sum_{m=1}^{M}\sum_{k=1}^{K}\left(\log\frac{2NP_{F}(\tau^{m,k}_{\rightarrow x_{m,k}};\eta(\mathcal{G}_{m};\theta),\rho)Z_{\theta}(\mathcal{G}_{m})}{R(x_{m,k};\beta)}\right)^{2}.$ (9)
 The $2N$ comes from the fact that TSP has $2N$ permutations, as discussed in Section 4.2.
 In addition, our target sampling distribution is proportional to reward $p(x;\mathcal{G})\propto R(x(\mathcal{G}))$.
-However, in most COPs, including TSP, the scale of $R(x(\mathcal{G}))$ is significantly different, possibly leading to unstable training due to the high variance of rewards.
-Similarly to previous works that stabilize REINFORCE training (Kwon et al., 2020; Kim et al., 2022b), we normalize the energy $\mathcal{E}=-\log R$ using the sample mean over a $K$ samples for the same $\mathcal{G}_{m}$, i.e.,
+However, in most COPs, including TSP, the scale of $R(x(\mathcal{G}))$ can vary significantly across instances, possibly leading to unstable training due to the high variance of logarithms of rewards.
+Similarly to previous works that stabilize REINFORCE training [43, 37], we normalize the energy $\mathcal{E}=-\log R$ using the sample mean over $K$ samples for the same $\mathcal{G}_{m}$, i.e.,
 $-\log\tilde{R}(x_{m,k})=\tilde{\mathcal{E}}(x_{m,k})=\mathcal{E}(x_{m,k})-\frac{1}{K}\sum_{k=1}^{K}\mathcal{E}(x_{m,k}).$ (10)
 We call this technique a shared energy normalization.
-Note that the shared energy normalization is applied independently to the two batches, $\mathcal{B}_{\text{explore}}$ and $\mathcal{B}_{\text{exploit}}$, as the energy distribution between the two batches is often significantly different to each other, especially when a heuristic local search is applied.
+Note that the shared energy normalization is applied independently to the two batches, $\mathcal{B}_{\text{explore}}$ and $\mathcal{B}_{\text{exploit}}$, as the energy distribution between the two batches is often significantly different from each other, especially when a heuristic local search is applied.
+To this end, the gradient loss is computed as follows:
+$\nabla_{\theta}\mathcal{L(\theta)}\approx\nabla_{\theta}\sum_{m=1}^{2M}\sum_{i=1}^{I}\left(\log\frac{2NP_{F}(\tau^{i,m}_{\rightarrow x_{i,m}};\eta(\mathcal{G}_{m};\theta),\rho)Z_{\theta}(\mathcal{G}_{m};\theta)}{\tilde{R}(x_{i,m})}\right)^{2}.$ (11)
+6 Experiments
+This section presents experimental results to validate our algorithm [^1] Source code is available: https://github.com/ai4co/gfacs.
+.
+We compare GFACS with other ACO algorithms, and with competitive reinforcement learning algorithms for vehicle routing problems.
+Moreover, we conduct an ablation study, and perform an empirical comparison with previous GFlowNets training methods, which serve a similar purpose to our method.
+Below, we provide a brief overview of the experimental setup.
+Implementation and hyperparameters.
+GFACS is built upon the implementation established by DeepACO [79].
+As our methodology mainly focuses on improving training algorithms rather than network architecture, we simply adopt the GNN architecture proposed in DeepACO, except the additional parameters for partition function $Z_{\theta}$ in Equation 9.
+For detailed network architecture, hyperparameters and hardware settings, please refer to Appendix A.
+It is worth noting that one important hyperparameter that GFACS newly introduces is the inverse energy temperature $\beta$, which enables us to control the diversity-optimality trade-offs (see Section D.2 for detailed analysis).
+In addition, we employ Ant System [14] for the pheromone updates throughout the experiments, while we also provide experimental results with other ACO variants in Section D.3.
+Test settings.
+Each algorithm is tested on the held-out datasets, each with 100 instances generated in advance except TSP, where we obtain the TSP test dataset with 128 instances from Ye et al.
+[79].
+Unless otherwise stated, the results presented throughout this section are the averaged values from three independent models, each trained with a distinct random seed.
+6.1 Comparison with ACO algorithms
+Here, we compare GFACS against the vanilla ACO with and the DeepACO [79] algorithm across various CO tasks.
+Problems.
+Following the approach of Ye et al.
+[79], we evaluate our algorithm in a range of COPs.
+This includes four routing problems – TSP, Capacitated Vehicle Routing Problem (CVRP), CVRP with Time Window (CVRPTW), Prize-Collecting Vehicle Routing Problem (PCTSP), and Orienteering Problem (OP) – along with a scheduling problem – Single Machine Total Weighted Tardiness Problem (SMTWTP) – and a grouping problem, the Bin Packing Problem (BPP).
+For detailed problem configurations, we direct readers to Appendix B.
+For TSP, CVRP and CVRPTW, we use local search algorithms with perturbation similar to Ye et al.
+[79].
+For others, we use the destroy-and-repair method (see Appendix C for more details).
 Figure 3: Results of ACO algorithms on various CO tasks.
@@
 The results are averaged over 3 independent models evaluated on the held-out test sets, and the shade indicates the min-max range of the 3 models.
-GFACS, without guided exploration, gives poorer performances.
-This verifies the effectiveness of our algorithm.
-6 Experiments
-This section presents the experimental findings of the GFACS framework.
-First, we test GFACS against established ACO baselines in various CO tasks.
-Subsequently, we explore the trade-offs between diversity and optimality in relation to the inverse energy temperature $\beta$.
-Next, an ablation study is conducted to evaluate the effectiveness of our proposed training techniques.
-Lastly, we compare GFACS with the RL-based combinatorial optimization methods in two representative routing problems, the TSP and the capacitated vehicle routing problem (CVRP).
-6.1 Setup
-Implementation and hyperparameters.
-GFACS is built upon the implementation established by DeepACO (Ye et al., 2023).
-As our methodology mainly focuses on improving training algorithms rather than network architecture, we simply adopt the GNN architecture proposed in DeepACO, except the additional parameters for partition function $Z_{\theta}$ in Eq.
-11.
-For detailed network architecture and hyperparameters, please refer to Appendix A.
-One important hyperparameter that GFACS newly introduces is the inverse energy temperature $\beta$, described in Eq.
-6.
-We offer an analysis of the impact and significance of $\beta$ in Section 6.4.
-Test settings.
-Each algorithm is tested on the held-out datasets generated or obtained in advance.
-Unless otherwise stated, the results presented throughout this section are the averaged values from three independent models, each trained with a distinct random seed.
-6.2 Comparison with ACO algorithms
-Here, we compare GFACS against the classical ACO and the recent DeepACO algorithm across various CO tasks.
-Problems.
-Following the approach of Ye et al.
-(2023), we evaluate our algorithm in a range of COPs.
-This includes four routing problems – TSP, Capacitated Vehicle Routing Problem (CVRP), Prize-Collecting Vehicle Routing Problem (PCTSP), and Orienteering Problem (OP) – along with two scheduling problems – Sequence Ordering Problem (SOP) and Single Machine Total Weighted Tardiness Problem (SMTWTP) – and a grouping problem, the Bin Packing Problem (BPP).
-For detailed problem configurations, we direct readers to Appendix B or Ye et al.
-(2023) as we mostly follow their problem settings.
-For TSP and CVRP, we use local search algorithms with neural-guided perturbation following Ye et al.
-(2023).
-For other problems, we use TopK guidance as described in Section 5.1.
 Baselines.
-We use ACO and DeepACO as baseline algorithms.
-Both baselines and GFACS use Ant System (Dorigo et al., 1996) for pheromone update (Eq.
-2).
-We also test GFACS without guided exploration to analyze its effect on the performance.
-We use 100 ants and $T=100$ ACO iterations for all the algorithms.
+We use vanilla ACO with a human-designed heuristic matrix, DeepACO as baseline algorithms.
+We also test GFACS without the proposed off-policy training scheme (called On-policy TB).
+For inference, we use $K=100$ ants and $T=100$ ACO iterations for all the algorithms.
 Results.
-As shown in Figure 5, GFACS consistently surpasses both the classic ACO algorithm and DeepACO across all benchmark problems.
-Especially for routing problems, GFACS outperforms DeepACO by a considerable margin.
-This is particularly remarkable given that both algorithms utilize identical neural networks and local search operators, highlighting the crucial role of our training strategy.
-Without guided exploration, GFACS generally performs badly, likely due to the over-exploration problem of GFlowNets.
-This issue is effectively mitigated with the proposed guided exploration, showing the efficacy of our method.
-Interestingly, in CVRP, OP, and PCTSP, GFACS without guided exploration outperforms DeepACO after a certain number of ACO iterations.
-We believe that this is attributed to the increased solution diversity of GFACS, which comes from the generative nature of the policy and implied DAG structure that incorporates the solution symmetricity.
-Figure 6: The trade-off between diversity and optimality as the inverse energy temperature $\beta$ changes.
-By adjusting $\beta$, we can control the trade-off of having a Pareto frontier to compare with DeepACO.
-The reported values are obtained with $T=1$, i.e., the solutions are not affected by the pheromone.
-6.3 Analysis of $\beta$ and Diversity-Optimality Trade-off
-In this subsection, we explored the diversity-optimality trade-off by varying the inverse energy temperature $\beta$ (Eq.
-6) from low to high values.
-For analysis, we generate $K$ solution samples from policies trained with GFACS with different values of $\beta$ and DeepACO (with $\rho=1$ so that the pheromone does not affect the result).
-We use average pairwise Jaccard distance as a diversity metric (formally defined in Appendix D).
-Results for TSP200 (Figure 6) show that lower $\beta$ values lead to more diverse but less optimal solutions, while higher $\beta$ values result in less diversity but higher optimality.
-GFACS confirms the anticipated trade-off between diversity and optimality, establishing a Pareto frontier over DeepACO by offering more diverse and high-quality solutions.
-This advantage is particularly beneficial when an effective local search mechanism is available, increasing the likelihood of achieving superior local optima, as illustrated in Figure 2.
-However, excessively high $\beta$ (e.g., $2000$ in our result) degrades the optimality, supposedly because too high $\beta$ value prevents exploration during training, which in turn results in sub-optimality.
-This empirical evidence supports our design choice of $\beta$ annealing, which encourages exploration in the early stage of training.
-Table 1: An ablation study examining various design elements of our method for TSP200 and CVRP200. Here, the optimality gap is used as a metric, where lower values indicate better performance.
-Method TSP200 CVRP200
-Ours 0.26% 1.09%
--- $\beta$ annealing 0.27% 1.11%
--- guided exloration 0.88% 1.42%
--- energy reshaping 0.59% 1.09%
--- shared energy normalization 3.11% 2.88%
-6.4 Ablation Study
-We also conducted an ablation study for four key training techniques introduced in Section 5: $\beta$ annealing, guided exploration, energy reshaping, and shared energy normalization.
-The ablation for $\beta$ annealing is conducted by fixing $\beta$ to the maximum value of the annealed $\beta$ schedule, while for other elements, we simply remove them from our algorithm.
-The empirical results for TSP200 and CVRP200 reported in Table 1 validate the effectiveness of our design choices.
-Table 2: Experimental result on TSP.
-Results for methods with * are drawn from Ye et al.
-(2023), Cheng et al.
-(2023a), and Jin et al.
-(2023).
-The reported time is the average computation time for solving a single instance.
-TSP200 TSP500 TSP1000
-Method Obj. Gap(%) Time(s) Obj. Gap(%) Times(s) Obj. Gap(%) Time(s)
-LKH-3 10.72 0.00 9.80 16.55 0.00 31.70 23.14 0.00 102
-AM* 12.00 11.91 0.2 21.46 29.07 0.6 33.55 45.10 1.5
-POMO* 10.89 1.62 0.10 20.57 24.40 0.6 32.90 42.30 4.1
-POMO + EAS 10.75 0.30 120 18.25 10.27 300 30.77 32.96 600
-POMO + SGBS 10.77 0.43 133 18.86 13.96 318 28.58 23.51 605
-DIMES+MCTS* - - - 17.01 2.78 11 24.45 5.75 32
-SO* 10.78 0.63 10 16.94 2.40 15 23.77 2.80 26
-Pointerformer* 10.79 0.68 5.5 17.14 3.56 14 24.80 7.90 40
-H-TSP* - - - - - - 24.65 6.62 0.3
-ACO $(T=10)$ 10.91 1.76 2.1 17.38 5.02 14 24.76 6.99 57
-DeepACO $(T=10)$ 10.77 0.45 2.5 16.84 1.75 15 23.78 2.77 66
-GFACS $(T=10)$ 10.75 0.26 2.5 16.80 1.54 15 23.72 2.53 66
-Table 3: Experimental result on CVRP. The reported time is the average computation time for solving a single instance.
-CVRP200 CVRP500 CVRP1000
-Method Obj. Gap(%) Time Obj. Gap(%) Time Obj. Gap(%) Time
-LKH-3 28.02 0.00 2.4m 63.07 0.00 18.6m 119.52 0.00 2.1h
-POMO 29.12 3.94 1s 68.69 8.90 2s 145.74 21.94 3s
-POMO + EAS 27.99 -0.11 2.0m 64.76 2.67 5.0m 126.25 5.63 10.0m
-POMO + SGBS 28.07 0.19 2.0m 65.44 3.75 5.2m 127.90 7.01 10.1m
-SymNCO 29.15 4.03 1s 68.81 9.10 2s 141.82 18.66 3s
-SymNCO + EAS 28.01 0.27 2.0m 64.63 2.47 5.0m 125.58 5.07 10.0m
-SymNCO + SGBS 28.10 0.30 2.0m 65.40 3.69 5.2m 127.53 6.70 10.1m
-ACO $(T=10)$ 28.45 1.55 13s 64.62 2.45 36s 122.82 2.76 1.3m
-DeepACO $(T=10)$ 28.44 1.49 14s 64.49 2.25 36s 122.15 2.20 1.3m
-GFACS $(T=10)$ 28.32 1.09 14s 64.19 1.76 36s 121.80 1.91 1.3m
+Figure 3 show GFACS consistently outperforms both the classic ACO algorithm and DeepACO across all benchmark problems.
+Notably, for routing problems, GFACS surpasses DeepACO by a significant margin.
+This is particularly remarkable given that both algorithms use identical neural networks and local search operators, underscoring the importance of our training strategy.
+Additionally, the proposed off-policy training method enhances GFACS performance, demonstrating its effectiveness for GFlowNets.
 6.2 Comparison with other reinforcement learning methods for combinatorial optimization
-In the domain of machine learning for combinatorial optimization, significant strides have been made using various approaches, including sampling-based methods (Sun et al., 2023), heatmap-guided searches (Sun & Yang, 2023; Qiu et al., 2022), and constructive methods that employ (Drakulic et al., 2023; Kool et al., 2018; Kwon et al., 2020).
-Our study compares our methods with RL-based approaches in TSP and CVRP for sizes $N=200,500,1000$, which do not require expert-labeled data.
-This comparison is particularly relevant as our unsupervised approach aligns with sequence decision-making processes akin to RL.
+In this section, we present experimental results comparing our method with reinforcement learning (RL)-based approaches, as both aim to develop combinatorial solvers without the need for labeled optimal data points, unlike supervised learning methods (e.g., [16, 70]).
 Baselines.
-For TSP baselines, we select the AM (Kool et al., 2018), POMO (Kwon et al., 2020), Sym-NCO (Kim et al., 2022b) and Pointerformer (Jin et al., 2023) as RL-based constructive methods, DIMES (Qiu et al., 2022) as RL-based heatmap-based methods, the Select and Optimize (SO) (Cheng et al., 2023b) as RL-based improvement methods, and H-TSP (Pan et al., 2023b) as a RL-based hierarchical methods.
-We also applied two search algorithms, the Efficient Active Search (EAS) (Hottung et al., 2021) and Simulation Guided Beam Search (SGBS) (Choo et al., 2022), to POMO for a more thorough comparison.
-For CVRP [^1] We also adopt TAM as a CVRP baseline; however, due to the inaccessibility of source code, we compare it using the reported result in the paper (Hou et al., 2022).
-See Section C.2..
-We choose POMO and Sym-NCO as baselines with EAS and SGBS.
-Note that we restrict the duration of these searches to no more than 2, 5, and 10 minutes for problems with 200, 500, and 1000 nodes, respectively.
-We use LKH-3 (Helsgaun, 2017) as an oracle heuristic solver for TSP and CVRP to obtain near-optimal solutions that serve as baselines with $0.00\%$ gap.
-Please refer to Section A.4 for the parameter settings of LKH-3.
+For TSP baselines, we select the AM [41], POMO [43], Sym-NCO [37] and Pointerformer [28] as RL-based constructive methods, DIMES [62] as RL-based heatmap-based methods, the Select and Optimize (SO) [9] as RL-based improvement methods.
+We also applied two search algorithms, the Efficient Active Search (EAS) [23] and Simulation Guided Beam Search (SGBS) [10], to POMO for a more thorough comparison.
+For CVRP, [^2] We also adopt TAM as a CVRP baseline; however, due to the inaccessibility of source code, we compare it using the reported result in the paper [24].
+See Section D.5.
+we employ POMO and Sym-NCO as baselines with EAS and SGBS.
+Note that we restrict the duration of these searches to no more than 5 and 10 minutes for problems with 500 and 1000 nodes, respectively.
+We use Concorde [2] for TSP and PyVRP [77] for CVRP as oracle heuristic solvers to obtain near-optimal solutions that serve as baselines with $0.00\%$ gap.
+For ACO algorithms, we use 100 ants with $T=10$ with local search algorithms similar to Ye et al.
+[79].
 Results.
 Among the baseline methods evaluated, our GFACS algorithm consistently delivers either superior or highly competitive performances.
-When compared to faster methods like AM, POMO or H-TSP, GFACS demonstrates substantially better results.
-Even after an additional search with EAS or SGBS, they fail to match GFACS’s achievements, except in CVRP200.
+When compared to faster methods like AM or POMO, GFACS demonstrates substantially better results.
+Even after an additional search with EAS or SGBS, they fail to match GFACS’s performance.
 Compared to algorithms like DIMES and Pointformer, which operate at similar speeds, GFACS achieves improved performance.
@@
 Therefore, we can conclude that GFACS, as a general-purpose algorithm, holds significant appeal in the highly competitive benchmarks of both TSP and CVRP.
-7 Conclusion
-This paper proposes a novel neural-guided meta-heuristic algorithm called generative flow ant colony sampler (GFACS).
-Our algorithm was designed by integrating generative flow networks (GFlowNets) and ant colony optimization (ACO) with a novel combination of training techniques.
-GFACS consistently outperforms other ACO methods in several representative combinatorial optimization tasks.
-GFACS also shows competitive results both in traveling salesman problems and capacitated vehicle routing problems compared to a range of RL-based methods.
+Table 1: Experimental results on TSP and CVRP.
+Results for methods with * are drawn from Ye et al.
+[79], Cheng et al.
+[8], and Jin et al.
+[28].
+The reported objective value is averaged over 128 instances of TSP and 100 instances of CVRP.
+We report the average computation time per instance.
+TSP500 TSP1000 CVRP500 CVRP1000
+Method Obj. Gap(%) Time Obj. Gap(%) Time Method Obj. Gap(%) Time Obj. Gap(%) Time
+Concorde 16.55 - 10.7s 23.12 - 108s PyVRP 62.96 - 21.0m 119.29 - 1.4h
+LKH-3 16.55 0.02 31.7s 23.14 0.10 102s LKH-3 63.07 0.18 18.6m 119.52 0.20 2.1h
+AM* 21.46 29.07 0.6s 33.55 45.10 1.5s
+POMO* 20.57 24.40 0.6s 32.90 42.30 4.1s POMO 68.69 9.08 2s 145.74 22.14 3s
+    +EAS 18.25 10.29 300s 30.77 33.06 600s     +EAS 64.76 2.85 5.0m 126.25 5.83 10.0m
+    +SGBS 18.86 13.98 318s 28.58 23.61 605s     +SGBS 65.44 3.93 5.2m 127.90 7.21 10.1m
+DIMES* 17.01 2.78 11s 24.45 5.75 32s SymNCO 68.81 9.28 2s 141.82 18.86 3s
+SO* 16.94 2.40 15s 23.77 2.80 26s     +EAS 64.63 2.65 5.0m 125.58 5.27 10.0m
+Pointerformer* 17.14 3.56 14s 24.80 7.90 40s     +SGBS 65.40 3.87 5.2m 127.53 6.90 10.1m
+ACO 17.38 5.04 14s 24.76 7.09 57s ACO 64.62 2.63 36s 122.82 2.96 1.3m
+DeepACO 16.84 1.77 15s 23.78 2.87 66s DeepACO 64.49 2.43 36s 122.15 2.40 1.3m
+GFACS (ours) 16.80 1.56 15s 23.72 2.63 66s GFACS (ours) 64.19 1.95 36s 121.80 2.11 1.3m
+Table 2: Experimental results on real-world datasets, TSPLib and CVRPLib-X. ‘#’ denotes the number of instances in each set. We report the average optimality gap using the best-known costs.
+TSPLib # ACO DeepACO Ours CVRPLib-X # ACO DeepACO Ours
+100-299 30 1.71% 1.25% 1.21% 100-299 43 2.50% 3.27% 2.69%
+300-699 10 4.26% 2.70% 2.60% 300-699 40 3.75% 4.22% 3.71%
+700-1499 12 7.01% 4.08% 3.88% 700-1001 17 4.64% 5.07% 4.32%
+Evaluation on real-world datasets.
+We compare the performance of our model against baseline models on real-world TSP and CVRP instances from the TSPLib [64] and CVRPLib [72] X datasets.
+The models were trained on random uniform instances of TSP/CVRP with sizes 200 for dataset sizes 100-299, 500 for sizes 300-699, and 1000 for sizes 700-1499, with all other settings remaining the same.
+Table 2 presents the results.
+GFACS outperforms the baselines in TSPLib and large CVRPLib instances but slightly underperforms compared to vanilla ACO on small CVRPLib instances, where ACO with Swap* local search suffices for small-scale CVRP.
+As the problem scale increases, ACO’s performance declines, underscoring the importance of effective priors using GFlowNets for efficient pheromone posterior updates in large combinatorial spaces.
+6.3 Ablation study
+Table 3: Results of the ablation study for the proposed training techniques.
+Method TSP500 CVRP500 BPP120
+GFACS 16.80 64.19 51.84
+   – Off-policy Training 17.09 64.42 52.32
+   – Energy Reshaping 16.87 64.41 52.07
+   – Shared $\varepsilon$-norm 17.52 65.51 90.73
+    + VarGrad [82] 16.86 64.28 51.95
+This section reports the ablation study results, validating our proposed training techniques for stabilizing and enhancing GFlowNets in large-scale combinatorial optimization tasks.
+As detailed in Table 3, our off-policy training method, incorporating local search, energy reshaping to value potential solutions, and shared $\epsilon$-norm to stabilize conditional GFlowNets training, significantly improves performance across three tasks: TSP, CVRP, and BPP.
+For comprehensive results, please refer to Table 10.
+6.4 Comparative analysis with previous GFlowNets training methods
+While this paper proposes a novel training method for GFlowNets designed for large-scale combinatorial optimization problems that are not locally decomposable due to their cyclic nature (e.g., TSP, CVRP) [1], we also compare existing GFlowNets training methods within this domain.
+Figure 4: Validation costs during training with different GFlowNet losses.
+Forward-looking detailed balance (FL-DB) [59, 84].
+FL-DB leverages local credits when the application benefits from useful local rewards for partial solutions.
+While graph combinatorial optimization tasks are generally locally decomposable [1], yielding useful local rewards, tasks like TSP do not offer substantial local rewards due to their cyclic nature, which imposes a global constraint, thus complicating local credit assignment.
+As shown in Figure 4, various GFlowNet methods were used to train the ACO’s heuristic matrix $\eta(\mathcal{G};\theta)$.
+While FL-DB did not surpass on-policy TB for TSP, our off-policy method demonstrated significant improvement.
+This trend holds across different deep learning models and inference tasks beyond ACO (see Section D.1).
+Gradient of log-parition variance (VarGrad) [65, 82].
+The VarGrad loss, initially proposed by Richter et al.
+[65] for variational inference, was applied to conditional GFlowNets for solving scheduling problems by Zhang et al.
+[82].
+This approach is crucial for combinatorial optimization tasks that require conditional inference on problem instances ($\mathcal{G}$).
+The key idea involves using a generative distribution and energy to compute the log-partition function through the consistency property of log-partitions over trajectories conditioned on identical instances ($\mathcal{G}$), thereby bypassing the explicit training of the (conditional) log-partition ($\log Z(\mathcal{G})$).
+Explicit training of the log-partition is challenging in conditional GFlowNets due to the varying energy landscapes across different instance conditions ($\mathcal{G}$).
+On the other hand, our method, which utilizes explicit training of the log-partition (see Equation 9), outperforms VarGrad in training conditional GFlowNets (see Table 3).
+This improvement is achieved because our Share $\epsilon$-norms enable stable and explicit training of the log-partition by normalizing energy values into sample means, resulting in an energy distribution across different instances that maintains a mean of zero.
+Note that accurately estimated the explicit log-partition value is known to stabilize the amortized inference process in GFlowNets [53].
+7 Discussion
+This paper proposes a novel probabilistic search method leveraging the off-policy training capabilities of GFlowNets.
+This strategy can be applied not only to ant-colony optimization (ACO) but also to other probabilistic methods requiring an effective learned prior (see Section D.4).
+Our off-policy framework enables the integration of learning algorithms with offline datasets containing optimal combinatorial solutions for specific problem instances, combined with online exploration to fine-tune the policy through off-to-on training.
+This approach holds promise for future work.
+Limitation.
+A limitation of our method is that its performance depends on the effectiveness of the local search.
+A promising future direction could involve leveraging advanced neural-guided improvement heuristics, including neural local search approaches [48, 50], to create a hybrid algorithm that combines GFlowNets with neural improvement-based local search.
 Appendix A Implementation
 A.1 Learning architecture and optimizer
-We utilize the anisotropic graph neural network with edge gating mechanisms, which is proposed by (Joshi et al., 2021) and adopted by (Ye et al., 2023), as the foundational GNN architecture for the construction of a heuristic matrix.
-This network offers node and edge embeddings that are more beneficial for edge prediction than other popular GNN variants, such as Graph Attention Networks (Veličković et al., 2018) or Graph Convolutional Networks (Kipf & Welling, 2016).
+We utilize the anisotropic graph neural network with edge gating mechanisms, which is proposed by [29] and adopted by [79], as the foundational GNN architecture for the construction of a heuristic matrix.
+This network offers node and edge embeddings that are more beneficial for edge prediction than other popular GNN variants, such as Graph Attention Networks [73] or Graph Convolutional Networks [40].
 We represent the node feature of node $v_{i}$ at the $\ell$-th layer as $h_{i}^{\ell}$, and the feature of the edge between nodes $i$ and $j$ at the same layer as $e_{ij}^{\ell}$.
 These features undergo propagation to the next layer $\ell+1$ through an anisotropic message passing scheme, detailed as follows:
-$h_{i}^{\ell+1}=h_{i}^{\ell}+\alpha(\text{BN}(U^{\ell}h_{i}^{\ell}+\mathcal{A}_{j\in\mathcal{N}_{i}}(\sigma(e^{\ell}_{ij})\odot V^{\ell}h_{j}^{\ell})))$ (13)
+$h_{i}^{\ell+1}=h_{i}^{\ell}+\text{ACT}(\text{BN}(U^{\ell}h_{i}^{\ell}+\mathcal{A}_{j\in\mathcal{N}_{i}}(\sigma(e^{\ell}_{ij})\odot V^{\ell}h_{j}^{\ell})))$ (12)
 $m^{\ell}=P^{\ell}e_{ij}^{\ell}+Q^{\ell}h_{i}^{\ell}+R^{\ell}h_{j}^{\ell}$ (13)
-$e_{ij}^{\ell+1}=e_{ij}^{\ell}+\alpha(\text{BN}(m^{\ell}))$ (15)
-where $U^{\ell},V^{\ell},P^{\ell},Q^{\ell},R^{\ell}\in R^{d\times d}$ are the learnable parameters of layer $\ell$, $\alpha$ denotes the activation function, BN(Ioffe & Szegedy, 2015) denotes batch normalization, $\mathcal{A}$ denote the aggregation function, $\sigma$ is the sigmoid function, $\odot$ is Hadamard product, $\mathcal{N}_{i}$ denotes the neighborhoods of node $v_{i}$.
-In this paper we use $\alpha$ is SiLU(Elfwing et al., 2018) and $\mathcal{A}$ is mean pooling.
+$e_{ij}^{\ell+1}=e_{ij}^{\ell}+\text{ACT}(\text{BN}(m^{\ell}))$ (14)
+where $U^{\ell},V^{\ell},P^{\ell},Q^{\ell},R^{\ell}\in R^{d\times d}$ are the learnable parameters of layer $\ell$, ACT denotes the activation function, BN[25] denotes batch normalization, $\mathcal{A}$ denote the aggregation function, $\sigma$ is the sigmoid function, $\odot$ is Hadamard product, $\mathcal{N}_{i}$ denotes the neighborhoods of node $v_{i}$.
+In this paper we use $\alpha$ is SiLU[17] and $\mathcal{A}$ is mean pooling.
 We transform the extracted edge features into real-valued heuristic metrics by employing a 3-layer Multilayer Perceptron (MLP) that includes skipping connections to the embedded node feature $h$.
 The SiLU is the activation function for all layers except the output layer, in which the sigmoid function is applied to generate normalized outputs.
-The training process is optimized using the AdamW optimizer(Loshchilov & Hutter, 2017) in conjunction with a Cosine Annealing Scheduler(Loshchilov & Hutter, 2016)
+The training process is optimized using the AdamW optimizer[46] in conjunction with a Cosine Annealing Scheduler[45]
 A.2 Hyperparameters
+Table 4: Training hyperparameter settings for each task.
+These hyperparameter settings are used both for GFACS and DeepACO training.
+The values with * indicate that the values are used for fine-tuning a pre-trained model.
+TSP200 TSP500 TSP1000 CVRP200 CVRP500 CVRP1000 OP100 PCTSP100 SMTWTP100 BPP120
+$N_{\text{epoch}}$ 50 50 20* 50 50 20* 20 20 20 20
+$N_{\text{inst}}$ 400 400 200* 200 200 100* 100 100 100 100
+$M$ 20 20 10 10 10 5 5 5 5 5
+$K$ 30 30 20 20 20 15 30 30 30 30
 For a task of COP (e.g., TSP200), we train GFACS for $N_{\text{epoch}}$ epochs.
@@
 Note that for TSP1000 and CVRP1000, we use model checkpoints pre-trained in TSP500 and CVRP500, respectively, to reduce the processing time for training.
-Table 4: Training hyperparameter settings for each task.
-These hyperparameter settings are used both for GFACS and DeepACO training.
-The values with * indicate that the values are used for fine-tuning a pre-trained model.
-TSP200 TSP500 TSP1000 CVRP200 CVRP500 CVRP1000 OP100 PCTSP100 SOP100 SMTWTP100 BPP120
-$N_{\text{epoch}}$ 50 50 20* 50 50 20* 20 20 20 20 20
-$N_{\text{inst}}$ 400 400 200* 200 200 100* 100 100 100 100 100
-$M$ 20 20 10 10 10 5 5 5 5 5 5
-$K$ 30 30 20 20 20 15 30 30 30 30 30
 There are other hyperparameters to be considered specifically for GFACS.
-$\beta$ is one of the most important hyperparameters for training GFACS, which is analyzed in Section 6.3.
-As we’ve discussed in Section 5.1, we use annealing for $\beta$ using the following log-shaped scheduling:
+$\beta$ is one of the most important hyperparameters for training GFACS, which is analyzed in Section D.2.
+We use annealing for $\beta$ using the following log-shaped scheduling:
 $\beta=\beta_{\text{min}}+(\beta_{\text{max}}-\beta_{\text{min}})\min\left(\frac{\log(i_{\text{epoch}})}{\log(N_{\text{epoch}}-N_{\text{flat}})},1\right)$
 where $i_{\text{epoch}}$ is the index of epoch and $N_{\text{flat}}$ is the number of last epochs to have $\beta=\beta_{\text{max}}$.
+Table 5: Settings of $\beta$ for each task.
+TSP CVRP CVRPTW OP PCTSP SMTWTP BPP
+$\beta_{\text{min}}$ 200 500 500 5 20 10 1000
+$\beta_{\text{max}}$ 1000 2000 2000 10 50 20 2000
 We report the $\beta$ value used for each task in Table 5. Note that $\beta$ remains consistent across different scales of the same problem class.
-Table 5: Settings of $\beta$ for each task.
-TSP CVRP OP PCTSP SOP SMTWTP BPP
-$\beta_{\text{min}}$ 200 500 5 20 10 10 500
-$\beta_{\text{max}}$ 1000 2000 10 50 20 20 5000
 For TSP and CVRP, where we have local search operators, we also need $\alpha$ for energy reshaping (Eq.
-9).
+7).
 We empirically found that linearly increasing $\alpha$ from $0.5$ to $1.0$ along the epochs works well, and we use that setting for all TSP and CVRP experiments.
@@
 After thet, the inference was conducted on a separate, standalone PC with a single AMD Ryzen 9 5900X 12-Core Processor and NVIDIA RTX 4070 Ti to assess the wall time of each algorithm fairly.
-A.4 Hyperparameters for LKH-3.
-LKH-3 (Helsgaun, 2017), an extension of the Lin-Kernighan heuristic (Lin & Kernighan, 1973), is a powerful open-source heuristic solver designed for solving routing problems.
-It not only addresses TSP but also handles a wide array of constrained routing problems, such as CVRP and PDP, by minimizing a penalty term that measures constraint violations.
-In the following, we describe the parameter settings for LKH-3 used in this work.
-For both TSP and CVRP, we set MAX_TRIALS to be $100\times N$ where $N$ is the number of nodes and SEED to $1234$.
-We set RUNS to 10 for TSP but 1 for CVRP to limit the search time in a reasonable range.
-For other parameters not stated, we followed the default settings.
 Appendix B Problem settings
@@
 For a comprehensive understanding of each problem’s specific settings, we direct readers to the work by Ye et al.
-(2023).
+[79].
 Furthermore, for access to the exact source code pertaining to each problem, please refer to the repository available at  [^3] https://github.com/henry-yeh/DeepACO from Ye et al.
-(2023).
+[79].
 Traveling salesman problem
@@
 We use $a=1$ and $b=9$, and a fixed $C=50$ for all problem scales, 200, 500 and 1000.
+Capacitated vehicle routing problem with time window
+A uniform-capacity fleet of delivery vehicles is tasked with servicing customers who have specific demand and operating hours (earliest-latest arrival) for a single commodity.
+The vehicles should deliver to customer only the time available for service.
+The goals are to minimize the total travel distance.
+A CVRPTW instance consists of a set of customer nodes, a single depot node, the demands of each customer node, the vehicle capacity (as in CVRP), and additionally, the earliest and latest arrival times for each node, as well as the service time at each node.
+We generate the CVRPTW instance following the method provided by Berto et al.
+[6], which can be accessed here [^4] https://github.com/ai4co/rl4co/blob/main/rl4co/envs/routing/cvrptw/generator.py.
 Orienteering problem
@@
 We sample a random OP instance by first sampling the position of nodes, including the starting node, from a unit square $[0,1]^{2}$.
-And we set the prize for each node, following (Fischetti et al., 1998; Ye et al., 2023), to be $p_{i}=(1+\lfloor 99\cdot\frac{d_{0,i}}{\max_{j=1}^{n}{d_{0,j}}}\rfloor)/100$, where $d_{0,i}$ is the Euclidean distance from the depot to node $i$.
+And we set the prize for each node, following [19, 79], to be $p_{i}=(1+\lfloor 99\cdot\frac{d_{0,i}}{\max_{j=1}^{n}{d_{0,j}}}\rfloor)/100$, where $d_{0,i}$ is the Euclidean distance from the depot to node $i$.
 We set the maximum length to 4 in the OP100 task, following Kool et al.
-(2018).
+[41].
 Prize-collecting traveling salesman problem
@@
 Lastly, the minimum total prize threshold is $N/4$ for problems with ${N}$ nodes, i.e., $25$ for PCTSP100.
-Sequential Ordering Problem
-The Sequential Ordering Problem (SOP) seeks to find a linear order of jobs that minimizes total processing time with a set of precedence constraints among the jobs.
-An SOP instance is a set of jobs and precedence constraints.
-Each job has its own processing time, and each pair has a waiting (transition) time.
-We sample an SOP instance by sampling the processing time and waiting time from $U[0,1]$.
-The precedence constraints for each job pair are laid with probability 0.2, while ensuring feasibility, i.e., when $i<j$ means that the constraint that job $i$ precedes the job $j$, the feasibility of a problem requires $i<j$ and $j<k$ then $i<k$.
 Single machine total weighted tardiness problem
@@
 Bin packing problem
-The Bin Packing Problem (BPP) involves packing objects of different volumes into a finite number of bins or containers of fixed capacity to minimize the number of bins used.
-In this work, we use the latter setting.
+The bin packing problem (BPP) aims to minimize the number of bins when objects with different volumes need to be packed into a finite number of bins.
+The cost functions are often modified to maximize the utility of used bins [18, 79].
+However, the direct use of the objective function employed in DeepACO can lead to an excessively large $\beta$ since the objective values tend to be saturated near 1.0; note that the objective values can not exceed 1.0 in that setting.
+Therefore, we slightly modify them to minimize a weighted sum of the inverse of the utilization of the used bins throughout all our BPP experiments.
 To generate a BPP instance, we sample the $N$ items with a random size from $U[20,100]$. And we use 150 for the bin capacity.
-Appendix C Additional Experiments
-C.1 Experiments on TSPLIB
-This section presents a comparison of performance between the baseline models and our model on real-world Traveling Salesman Problem (TSP) instances.
-We evaluated the models on TSP200, TSP500, and TSP1000 under their respective conditions.
-Our model outperforms the other Ant Colony Optimization (ACO) baselines, demonstrating superior generalization capabilities.
-Table 6: Evaluation of ACO methods on the real-world TSP problem. ‘#’ denotes the number of instances in each set.
-Size # ACO DeepACO Ours
-$300100\leq N<$ 300 30 1.71% 1.25% 1.21%
-$700300\leq N<$ 700 10 4.26% 2.70% 2.60%
-$1500700\leq N<$ 1500 12 7.01% 4.08% 3.88%
-ALL 52 3.43% 2.24% 2.16%
-C.2 Comparison against TAM
-We compare GFACS and DeepACO against CVRP instances used in TAM (Hou et al., 2022).
-These instances have a different capacity constraint from those we used for the main experiment.
-Please refer to (Hou et al., 2022) for more details regarding the instance generation.
-Moreover, to match the processing time to be comparative to DeepACO, we set $K$ (equal to the number of ants for ACO) to 16.
-Table 7: Results for CVRP instances used in TAM. The result for AM and TAM is drawn from Hou et al. (2022), and the result for DeepACO is drawn from Ye et al. (2023).
+Appendix C Implementation of destroy-and-repair local search
+As demonstrated in Section 5, our algorithm utilizes the local search for off-policy training.
+The refined solutions from local search serve as exploitation batches, and they are also used for energy-reshaping (Eq.
+7).
+For some problems like traveling salesman problem (TSP), capacitated vehicle routing problem (CVRP), and CVRP with time window (CVRPTW), we used the well-established local search operators such as 2-opt, Swap* [74] and granular neighborhood search from PyVRP [77].
+While domain-specific local search techniques are well developed for many combinatorial optimization problems, there are still tasks that lack effective local search solvers like Swap*.
+A potential candidate for a general-purpose local search solver in combinatorial optimization is the destroy-and-reconstruct local search operator.
+This method is inspired by the large neighborhood search approach described by Pisinger and Ropke [61].
+Our implementation utilizes an evolutionary-style algorithm that repeatedly performs destruction, reconstruction, and top-K selection over a specified number of rounds.
+Symmetricity-aware destruction.
+The most naive way to destroy a solution $x$ is simply dropping $N_{destroy}$ of the last nodes from the sampled trajectory $\tau_{\to x}$.
+However, it only performs the destruction with one trajectory, while there can be multiple trajectories that correspond to the same solution $x$.
+Considering this symmetric nature, our destruction phase consists of 1) selecting $\tau^{\prime}_{\to x}$ randomly from the set of symmetric trajectories and then 2) destroying $\tau^{\prime}_{x}$ by dropping the last $N_{destroy}$ nodes, resulting in a partial solution.
+Note that the number of symmetric trajectories for a given solution differs from problem types, as discussed in Appendix E.
+Reconstruction with temperature annealing.
+Given a partial solution, we reconstruct a new solution $x^{\prime}$ using the policy similar to the decoding steps explained in Section 4.2.
+Considering the purpose of the local search (refining a given solution), we further promote the exploitation by introducing an inverse temperature for the policy.
+Formally, we modify Equation 5 with the inverse temperature $\delta_{r}$ as follows:
+$P_{F}(s_{t+1}|s_{t};\eta(\mathcal{G};\theta),\rho,\delta_{r})\propto\begin{cases}\left(\rho_{a_{t-1},a_{t}}\eta_{a_{t-1},a_{t}}(\mathcal{G};\theta)\right)^{\delta_{r}}&\text{if $a_{t}\notin s_{t}$ }\\ 0&\text{otherwise}\end{cases}$ (15)
+We increase $\delta_{r}$ linearly from 1.0 to a specified maximum value, $\delta_{\text{max}}$, as the local search round progresses.
+Top-K selection.
+For a given solution $x$, we perform the aforementioned destruction and reconstruction multiple times (in batch), obtaining a set of refined solutions.
+Then we evaluate them and leave only top-K solutions among them in terms of solution quality.
+The remaining solutions are used as an initial solution for the next round of the search.
+Appendix D Additional experiments
+D.1 Comparison between trajectory balance and forward-looking detailed balance
+We trained Attention Model (AM) [41] and heatmap-based policy from Joshi et al.
+[29] in an unsupervised way using FL-DB and TB loss.
+Table 6 shows the test results for 128 random TSP50 instances.
+The result is aligned with the previous observation from GFACS, showing the limitation of FL-DB loss for solving complex CO tasks.
+Table 6: Comparison between forward-looking detailed balance (FL-DB) and trajectory balance (TB) loss in TSP50, using Attention Model (AM) and heatmap-based method from Joshi et al.
+[29] as backbone policies.
+TSP50
+Method Obj. Gap (%)
+Concorde 5.69 0.00
+[…]
+[…]
+[…]
+[…]
+D.2 Analysis of $\beta$ and diversity-optimality trade-off
+In this subsection, we explored the diversity-optimality trade-off by varying the inverse energy temperature $\beta$ from low to high values.
+For analysis, we generate $K$ solution samples from policies trained with GFACS with different values of $\beta$ and DeepACO (with $\rho=1$ so that the pheromone does not affect the result).
+We use average pairwise Jaccard distance as a diversity metric.
+Formally, the diversity of a set of $K$ solutions is calculated as $\frac{1}{K(K-1)}\sum_{i}\sum_{j}d_{J}(x_{i},x_{j})$, where $d_{J}$ is a Jaccard distance between two solutions, defined as follows:
+$d_{J}(x_{i},x_{j})=1-J(x_{i},x_{j})=1-\cfrac{|E(x_{i})\cap E(x_{j})|}{|E(x_{i})\cup E(x_{j})|}$ (16)
+$E(x)$ is the set of edges composing the solution $x$.
+In TSP or CVRP, the (undirected) edge is a set of adjoining nodes in the solution path.
+For some CO problems, such as SMTWTP, the edge should be directed due to its asymmetric nature.
+Figure 5: The trade-off between diversity and optimality as the inverse energy temperature $\beta$ changes.
+Figure 5 show that lower $\beta$ values lead to more diverse solutions with higher costs, while higher $\beta$ values result in less diverse solutions with lower costs.
+GFACS confirms the anticipated trade-off between diversity and optimality, establishing a Pareto frontier over DeepACO by offering more diverse and high-quality solutions.
+This advantage is particularly beneficial when an effective local search mechanism is available, increasing the likelihood of achieving superior local optima.
+However, excessively high $\beta$ (e.g., $2000$ in our result) degrades both diversity and solution quality, supposedly because too much high $\beta$ hinders exploration during training.
+This empirical evidence supports our design choice of $\beta$ annealing Section A.2, which encourages exploration in the early stage of training.
+D.3 Evaluation with other ACO variants
+Table 7: Evaluation results with different pheromone update rules, Elitist AS and MAX-MIN AS. We run $T=10$ ACO rounds with 100 ants in all settings.
+TSP200 CVRP200
+Method Obj. Gap (%) Obj. Gap (%)
+Concorde / PyVRP 10.72 0.00 27.93 0.00
+AS 10.91 1.77 28.45 1.86
+DeepACO + AS 10.77 0.45 28.44 1.80
+GFACS + AS 10.75 0.26 28.32 1.40
+Elitist AS 10.97 2.31 28.46 1.90
+DeepACO + Elitist AS 10.77 0.49 28.45 1.84
+GFACS + Elitist AS 10.75 0.28 28.34 1.47
+MAX-MIN AS 10.97 2.34 28.46 1.88
+DeepACO + MAX-MIN AS 10.78 0.57 28.44 1.82
+GFACS + MAX-MIN AS 10.75 0.32 28.33 1.43
+Our method is focused on enhancing the design of the heuristic matrix within the ACO framework.
+Thus, the advancement of the ACO algorithm itself (usually in pheromone update and action selection rule) is orthogonal to our work.
+To validate this claim, we evaluate our method with different ACO variants other than AS, namely Elitist Ant System [14], MAX-MIN Ant System [68].
+The results in Table 7 demonstrate that GFACS consistently outperforms both ACO with human-designed heuristic and DeepACO under all pheromone update rules tested.
+This underscores the superiority of GFACS in generating a high-quality heuristic matrix.
+D.4 Applying our method to other search algorithms
+Table 8: Evaluation results combine with active search as posterior inference method.
+TSP200 CVRP200
+Method Obj. Gap (%) Obj. Gap (%)
+Concorde / PyVRP 10.72 0.00 27.93 0.00
+Pretrained GFACS
+   – Off-policy TB Active Search 10.75 0.29 28.34 1.47
+   – On-policy RL Active Search 10.76 0.40 28.38 1.62
+Pretrained DeepACO
+   – On-policy RL Active Search 10.79 0.65 28.43 1.79
+In this subsection, we study active search [3] rather than the pheromone update of ACO.
+In ACO, pheromone update is a posterior inference given a prior distribution (heuristic) for the probability model that generates the solutions.
+Similarly, we can think of active search as a posterior update that implicitly updates a given probability model by adjusting the parameters that generate it.
+We sample 20 solutions and take 5 gradient steps to update the parameter at each iteration.
+After 50 iterations, we report the average performance of the active search using our method with 3 seeds of 128 instances for each task.
+Table 8 shows that GAFCS integrates better with active search than DeepACO.
+This indicates that GAFCS is effective not only with heuristic-based posterior inference but also with learning-based posterior inference.
+D.5 Comparison against TAM
+Table 9: Results for CVRP instances used in TAM. The results for AM and TAM is drawn from Hou et al. [24], and the result for DeepACO is drawn from Ye et al. [79].
 CVRP100 CVRP400
@@
 GFACS ($T=10$) 15.72 1.40 24.87 5.21
-Appendix D Diversity Calculation
-The diversity of $K$ solutions generated from a solver is defined as the average pairwise Jaccard distance, $\frac{1}{K(K-1)}\sum_{i}\sum_{j}d_{J}(x_{i},x_{j})$, where $d_{J}$ is defined as follows:
-$d_{J}(x_{i},x_{j})=1-J(x_{i},x_{j})=1-\cfrac{|E(x_{i})\cap E(x_{j})|}{|E(x_{i})\cup E(x_{j})|}$ (16)
-where $E_{x}$ is the set of edges composing the solution $x$. In TSP or CVRP, the (undirected) edge is a set of adjoining nodes in the solution path.
+We compare GFACS and DeepACO against CVRP instances used in TAM [24].
+These instances have a different capacity constraint from those we used for the main experiment (Table 1).
+Please refer to [24] for more details regarding the instance generation.
+Moreover, to match the processing time to be comparative to DeepACO, we set $K$ (equal to the number of ants for ACO) to 16.
+D.6 Ablation study
+Table 10: Results of the ablation study.
+The mean and standard deviation are reported based on three independent runs.
+The reported gap is calculated against Concorde [2] for TSP and PyVRP [77] for CVRP.
+TSP500 CVRP500 BPP120
+Method Obj. Gap (%) Obj. Gap (%) Obj.
+GFACS 16.0842 $\pm$ 0.0043 1.56 64.1858 $\pm$ 0.0071 1.95 51.8428 $\pm$ 0.0706
+   – Off-policy Training 17.0861 $\pm$ 0.0028 3.26 64.4248 $\pm$ 0.0155 2.33 52.3158 $\pm$ 0.1171
+   – Energy Reshaping 16.8718 $\pm$ 0.0049 1.97 64.4121 $\pm$ 0.0069 2.31 52.0654 $\pm$ 0.0998
+   – Shared $\varepsilon$-norm 17.5209 $\pm$ 0.0062 5.87 65.5069 $\pm$ 0.0171 4.05 90.7332 $\pm$ 9.1649
+    + VarGrad [82] 16.8570 $\pm$ 0.0086 1.88 64.2771 $\pm$ 0.0293 2.09 51.9504 $\pm$ 0.1222
+In this section, we present the comprehensive results of our ablation study, incorporating standard deviations from multiple independent runs (three runs) to ensure statistical significance.
+The findings demonstrate that each component of our technical proposals contributes to improvement, supported by empirical and statistically rigorous evidence.
 Appendix E Symmetric solutions of combinatorial optimization
@@
 Constructive policy methods systematically build solutions by incrementally adding elements, beginning with an empty set and culminating in a complete solution.
-This approach ensures compliance with constraints by progressively narrowing the scope of actions, making it versatile for application in domains like vehicle routing (Nazari et al., 2018), graph combinatorial optimization (Khalil et al., 2017), and scheduling (Zhang et al., 2020).
+This approach ensures compliance with constraints by progressively narrowing the scope of actions, making it versatile for application in domains like vehicle routing [57, 41], graph combinatorial optimization [30, 1, 84], and scheduling [81].
 In deep learning, various techniques adopt this constructive policy framework.
-Notably, the attention model (AM) (Kool et al., 2018) has been pivotal in vehicle routing problems (VRPs), establishing a new standard for stabilizing REINFORCE training.
-Subsequent advancements in AM have included enhanced decoder architecture (Son et al., 2023a; Xin et al., 2021; Luo et al., 2023), improved REINFORCE (Williams, 1992) baseline assessments (Kwon et al., 2020; Kim et al., 2022b), leveraging symmetries for sample efficiency (Kim et al., 2023a), the application of population-based reinforcement learning (RL) (Grinsztajn et al., 2022), and refined solution-building methods (Drakulic et al., 2023).
-Recent advancements in hierarchical algorithms exploit AM-based policies for solution optimization: one policy broadly generates solutions while another refines them locally (Kim et al., 2021b; Hou et al., 2022; Ye et al., 2024).
-Additionally, AM-driven constructive policies have been effectively applied in hardware routing (Kim et al., 2020, 2021a) and placement tasks (Park et al., 2022; Kim et al., 2023b).
+Notably, the attention model (AM) [41] has been pivotal in vehicle routing problems (VRPs), establishing a new standard for stabilizing REINFORCE training.
+Subsequent advancements in AM have included enhanced decoder architecture [67, 78, 47], improved REINFORCE [76] baseline assessments [43, 37], leveraging symmetries for sample efficiency [32], the application of population-based reinforcement learning (RL) [21], ensemble-based learning [27, 86] and refined solution-building methods [16].
+Recent advancements in hierarchical algorithms exploit AM-based policies for solution optimization: one policy broadly generates solutions while another refines them locally [36, 24, 80].
+Additionally, AM-driven constructive policies have been effectively applied in hardware routing [34, 35] and placement tasks [60, 31].
 Graph combinatorial optimization, particularly in problems like maximum independent sets, has diversified from contrive VRPs policies due to their local decomposability and the relative ease of training value functions for partial solutions where VRPs are not locally decomposable due to their global contains.
 Starting with the influential works by Khalil et al.
-(2017), which employed deep Q networks (Mnih et al., 2013) for graph policy training, creating solutions constructively, Ahn et al.
-(2020) introduced an innovative Markov Decision Process (MDP) to generate multiple solution components per iteration by exploiting the local decomposability of graph combinatorial problems.
+[30], which employed deep Q networks [55] for graph policy training, creating solutions constructively, Ahn et al.
+[1] introduced an innovative Markov Decision Process (MDP) to generate multiple solution components per iteration by exploiting the local decomposability of graph combinatorial problems.
 Zhang et al.
-(2023a) further developed this concept, training their MDP with forward-looking GFlowNets (Pan et al., 2023a).
+[84] further developed this concept, training their MDP with forward-looking GFlowNets [59].
 Our research aligns with the deep constructive policy framework, where GFlowNets constructively build actions from an initial empty solution ($s_{0}$) to a final complete solution ($s_{n}$).
@@
 F.2 Deep improvement policy for neural combinatorial optimization
-The improvement policy focuses on refining complete solutions, similar to local search methods like 2opt (Croes, 1958).
+The improvement policy focuses on refining complete solutions, similar to local search methods like 2opt [11].
 Its advantage lies in achieving comparable performances to constructive policies, especially when allotted sufficient time for extensive improvement iterations.
-Researchers typically adapt existing heuristic local search techniques, creating learning policies that emulate these methods through maximizing expected returns via reinforcement learning (d O Costa et al., 2020; Hottung & Tierney, 2019; Ma et al., 2021, 2022, 2023; Kim et al., 2022a).
-Also, recent works have explored active search to refine solutions via policy adjustments (Hottung et al., 2021; Son et al., 2023b).
-Moreover, alternative strategies using discrete Langevin sampling instead of neural networks have shown promising outcomes (Sun et al., 2023).
+Researchers typically adapt existing heuristic local search techniques, creating learning policies that emulate these methods through maximizing expected returns via reinforcement learning [12, 22, 48, 49, 50, 33].
+Also, recent works have explored active search to refine solutions via policy adjustments [23, 66] or sampling the conditions from a continuous latent space [7].
+Moreover, alternative strategies using discrete Langevin sampling instead of neural networks have shown promising outcomes [69].
 This approach serves as a complementary extension to deep constructive policies, as it often enhances solutions initially formulated by constructive methods.
@@
 F.3 Heatmap-based neural combinatorial optimization
-The heatmap-based method (Joshi et al., 2021) entails the acquisition of a graph representation, referred to as the “Heatmap,” from input graph instances.
+The heatmap-based method [29] entails the acquisition of a graph representation, referred to as the “Heatmap," from input graph instances.
 This Heatmap serves as a prior distribution, influencing the generation of a solution.
-However, this approach is not self-contained and requires an additional solution construction step, which may involve techniques like beam search (Joshi et al., 2021), dynamic programming (Kool et al., 2022), or Monte Carlo tree search (Fu et al., 2021; Qiu et al., 2022; Sun & Yang, 2023).
+However, this approach is not self-contained and requires an additional solution construction step, which may involve techniques like beam search [29], dynamic programming [42], Monte Carlo tree search [20, 62, 70] or greedy search with a guide using an objective function [44].
 The key advantage of the heatmap-based method lies in its exceptional generalizability when compared to constructive policies.
@@
 A key advantage of our method over many other supervised learning-based heatmap approaches is its independence from optimally labeled data.
+Broader Impact
+We proposed a new probabilistic approach that can be trained using off-policy learning and applied to various combinatorial optimization problems, such as vehicle routing and scheduling.
+This method is relevant to industrial engineering, particularly in optimizing transportation and manufacturing processes, and can potentially extend to scientific discovery through chemical structure and biological sequence optimization.
+It is important to consider the ethical implications of probabilistic optimization methods, as they carry the potential for misuse in fields such as harmful drug discovery.
+While these algorithms offer several benefits, their application should be carefully monitored to ensure responsible and ethical use, minimizing the risk of unintended negative consequences.

### v4

@@
 Abstract
-This paper introduces the Generative Flow Ant Colony Sampler (GFACS), a neural-guided probabilistic search algorithm for solving combinatorial optimization (CO).
-GFACS integrates generative flow networks (GFlowNets), an emerging amortized inference method, with ant colony optimization (ACO), a promising probabilistic search algorithm.
-Specifically, we use GFlowNets to learn a constructive policy in combinatorial spaces for enhancing ACO by providing an informed prior distribution over decision variables conditioned on input graph instances.
-Furthermore, we introduce a novel off-policy training algorithm for scaling conditional GFlowNets into large-scale combinatorial spaces by leveraging local search and shared energy normalization.
-Our experimental results demonstrate that GFACS outperforms baseline ACO algorithms in seven CO tasks and is competitive with problem-specific heuristics for vehicle routing problems.
+We present the Generative Flow Ant Colony Sampler (GFACS), a novel meta-heuristic method that hierarchically combines amortized inference and parallel stochastic search.
+Our method first leverages Generative Flow Networks (GFlowNets) to amortize a multi-modal prior distribution over combinatorial solution space that encompasses both high-reward and diversified solutions.
+This prior is iteratively updated via parallel stochastic search in the spirit of Ant Colony Optimization (ACO), leading to the posterior distribution that generates near-optimal solutions.
+Extensive experiments across seven combinatorial optimization problems demonstrate GFACS’s promising performances.
 1 Introduction
-Combinatorial optimization (CO) is a crucial domain involving high-impact applications, including vehicle routing, hardware design optimization, and biological sequence design.
-Ant colony optimization (ACO) [14, 68] is a general-purpose meta-heuristic algorithm designed to address these tasks.
-ACO employs a constructive solution generation scheme that sequentially adds elements to a partial solution, known for its flexibility across CO tasks.
-This approach involves collecting experiences from multiple agents (ants) to iteratively update a distribution (pheromone update) based on a prior distribution known as the heuristic matrix.
-The major benefits of ACO include: (1) its robustness against local optima, enabled by the algorithm’s stochastic nature, which dynamically balances exploration and exploitation, preventing premature convergence [14]; and (2) its flexible solution generation scheme, which allows it to effectively solve a wide range of combinatorial optimization tasks [79].
-The main challenge with ACO lies in designing effective prior distributions (i.e., the heuristic matrix) for their probability models.
-Inadequate priors can lead to time-consuming pheromone update iterations.
-Recent advancements, known as DeepACO [79], address this by using policy gradient reinforcement learning (RL) [76, 71] to pretrain the priors conditioned on given problem instances.
-RL trains these networks to guide highly rewarded regions, allowing generalization to new problems and enhanced performance.
-Motivated by DeepACO, we propose a deep learning-based prior distribution design.
-However, this paper tackles challenges related to the design of prior distribution through deep amortized inference rather than the policy gradient RL method.
-This is motivated by the following three primary limitations in DeepACO’s use of policy gradient RL.
-First ( ), the on-policy nature of DeepACO restricts the flexible use of a behavior policy, potentially undermining performance by limiting effective exploration of vast combinatorial spaces.
-Second ( ), the reward maximizing form of RL in DeepACO struggles to capture the diverse solutions necessary in CO, even though having a variety of good candidates could significantly enhance the likelihood of finding a global optimum after local searches.
-Last ( ), the inherent symmetry in combinatorial solutions—where different paths can lead to identical outcomes—presents a significant challenge not adequately addressed in DeepACO.
+Combinatorial optimization (CO) aims to optimize an objective function over discrete, finite combinatorial structures under given constraints.
+It has broad applications in operations research (Zhang et al.,, 2019), scientific discovery (Naseri and Koffas,, 2020) and machine learning (Chickering et al.,, 1995, 2004).
+Traditionally, integer programming approaches have been extensively studied to solve CO problems (Miller et al.,, 1960).
+While these methods offer guaranteed optimality, they often become impractical for large-scale instances due to their exponential complexity.
+An alternative is to design problem-specific heuristics, but it requires domain knowledge and labor-intensive design, limiting flexibility to variant problems.
+* Equal contribution.
+$\dagger$ Work done while Hyeonah Kim was at KAIST.
+Code available at: https://github.com/ai4co/gfacs.
+Correspondence to: Minsu Kim (minsu.kim@mila.quebec) and Sanghyeok Choi (sanghyeok.choi@kaist.ac.kr).
+Recent advancements in deep learning (LeCun et al.,, 2015) brought alternative directions to CO.
+For a more detailed review, refer to our related works in Appendix F and the survey by Bengio et al., 2021b .
+Among deep learning for CO methods, supervised learning (SL) approaches distill mathematical programming methods or heuristics into powerful deep neural networks, enabling fast and scalable solution generation (Vinyals et al.,, 2015; Joshi et al.,, 2021; Sun and Yang,, 2023; Luo et al.,, 2023).
+However, SL methods rely on expert-labeled data, limiting their applicability as general-purpose CO solvers.
+Reinforcement learning (RL) is another popular approach to CO (Khalil et al.,, 2017; Kool et al.,, 2019; Kwon et al.,, 2020; Ye et al.,, 2023).
+Unlike supervised learning, RL methods train deep neural networks directly using the objective function of a CO problem as a reward, without requiring expert-labeled data.
+This makes RL methods particularly useful because they can adapt flexibly to a wide range of CO problems, even where specific heuristics are unavailable or unexplored.
+However, RL requires extensive exploration of vast solution space and often involves non-trivial credit assignments across multiple decisions, which introduces significant complexity to the learning process compared to SL (Minsky,, 1961).
+Hierarchical methods that leverage RL in a two-stage problem-solving procedure tackle such exploration and credit assignment problems through hierarchical decomposition.
+Previous works by Bello et al., (2016), Hottung et al., (2022), and Son et al., (2023) pretrain large models on massive problem instances using RL and fine-tune these models on specific problem instances at test time using a method called active search (AS).
+Ye et al., (2023) pretrain relatively small models with significantly reduced training times.
+They pretrain a neural network with RL to produce a prior distribution over solution space and iteratively update the distribution using ant colony optimization (ACO), an effective parallel stochastic search algorithm.
 Contribution.
-This paper introduces a novel ACO algorithm, the Generative Flow Ant Colony Sampler (GFACS), which addresses the limitations of DeepACO by integrating the emerging amortized inference method of GFlowNets [4, 5].
-Our twofold contribution can be summarized as follows.
-(1) We introduce a new viewpoint to ACO using amortized inference instead of policy gradient RL.
-(2) Additionally, we present a novel GFlowNet training algorithm for large scale combinatorial tasks, applicable to a range of further uses over ACO.
-  • ACO as off-policy amortized inference with GFlowNets.
-There are three reasons that GFlowNets can improve over the existing DeepACO algorithm.
-First ( ), the off-policy training property of GFlowNets allows for the utilization of samples drawn from any behavior policy, including powerful heuristic search algorithms commonly used in CO problems.
-Secondly ( ), unlike maximum reward-seeking algorithms (e.g., RL training in DeepACO), GFlowNets are sampling algorithms that are well-suited for modeling the multimodal nature of certain distributions.
-Lastly ( ), GFlowNets model solutions as a generative process on a directed acyclic graph (DAG), which can represent multiple trajectories that share identical states.
-This approach allows for accurate probabilistic modeling by considering the symmetry of combinatorial solutions [5].
-  • Scaling conditional GFlowNets into large-scale combinatorial space.
-GFlowNets often struggle with limited scalability due to challenging exploration problems [58, 39], credit assignment issues [59, 26], and unstable training when inferring conditional distributions [82].
-In combinatorial optimization, both limitations must be addressed because we deal with (1) large-scale combinatorial spaces and (2) conditional inference on problem instances.
-This study introduces a new off-policy training method for GFlowNets in large combinatorial spaces by effectively leveraging local search operators.
-It also proposes a novel training method to address instability in training conditional GFlowNets by properly normalizing conditional reward values with sample mean.
-In comparative experiments, our GFACS exhibits significant superiority over the current DeepACO and classical ACO algorithms for various representative CO tasks – vehicle routing problems, scheduling problems, and a grouping problem.
-Moreover, GFACS outperforms or shows similar performance to competitive unsupervised learning-based vehicle routing solvers, which are designed to solve traveling salesman problems (TSP) and capacitated vehicle routing problems (CVRP) in the context of large-scale instances, such as $N=500,1000$.
-Lastly, we compared our approach with previous GFlowNet methods for solving combinatorial optimization problems [82, 84] in vehicle routing tasks, which are known to pose challenges for credit assignment due to their locally undecomposable properties, as highlighted by Ahn et al.
-[1].
-2 Related works
-2.1 Ant colony optimization
-Ant Colony Optimization (ACO) is a meta-heuristic inspired by ant behavior in finding optimal paths.
-Traditional ACO algorithms, such as Ant System (AS) [14] and MAX-MIN AS [68], rely on specific heuristics for success.
-To reduce the reliance on heuristic design, Ye et al.
-[79] introduce a neural-guided ACO, called DeepACO, utilizing graph representation learning and reinforcement learning to replace manual heuristics with automatically acquired knowledge.
-2.2 Neural combinatorial optimization
-Neural combinatorial optimization [3] is a deep learning-based method to train an instance-conditional policy using supervised learning [75, 20, 70] or unsupervised learning [30, 41, 43, 54].
-Usually, supervised learning methods are beneficial when the target problem has an accurate but time-consuming expert solver.
-The oracle-labeled data is collected from the oracle solvers, and deep neural networks amortize the search process, converting an expensive run-time optimization into an also expensive but one-time cost of training the neural net, allowing for scalability and faster run-time inferences than the expert solver.
-On the other hand, unsupervised learning proceeds through self-guided trial and error, using the COP objective as a training signal.
-This makes it applicable even if encountering less-studied problems where powerful expert solvers are unavailable [3, 1].
-Our method can be categorized as an unsupervised learning method, which does not need expert labels.
-See Appendix F for a more comprehensive survey.
-2.3 GFlowNets and combinatorial optimization
-GFlowNets are designed for sampling combinatorial objects proportionally to a given reward function [4].
-Theoretical studies have connected GFlowNets with variational inference [53] and maximum entropy reinforcement learning [56, 13].
-Recent algorithmic advancements in GFlowNets have focused on developing new training objectives [52, 51, 85], enhancing credit assignment [59, 26], and improving exploration strategies [58, 63, 59, 39, 38].
-GFlowNets have also found successful applications in some combinatorial optimization problems (COPs) [82, 84].
-Zhang et al.
-[84] effectively address the symmetric nature of graph constraint optimization, demonstrating remarkable performance in tasks such as finding maximum independent sets.
-Despite their success, there are limitations in expanding to other combinatorial tasks due to their reliance on local credits for the learning objective proposed by Pan et al.
-[59].
-Conversely, Zhang et al.
-[82] address proxy-based scheduling problems by developing a problem-specific neural architecture and introducing novel training methods for general-purpose conditional GFlowNets.
-One of their significant contributions is the proposal of a loss function based on the sample variance of the log partition function, known as VarGrad, which enhances the training process and performance.
+In this work, we propose a novel probabilistic meta-heuristic called Generative Flow Ant Colony Sampler (GFACS), which leverages the synergy between a multi-modal amortized prior and iterative update of it using parallel stochastic search.
+Unlike previous approaches that use RL pretraining aimed at maximizing expected rewards, we employ Generative Flow Networks (GFlowNets; Bengio et al., 2021a, ; Bengio et al.,, 2023), designed to learn a distribution where the probability of a sample is proportional to its reward, i.e., $p(x)\propto R(x)$.
+This leads to a multi-modal prior distribution of solutions, allowing it to benefit more from subsequent iterative posterior updates compared to the unimodal priors pretrained with reward maximizing RL, as illustrated in Figure 1.
+Moreover, to enhance and stabilize GFlowNet training, we introduce several new techniques that have been verified to be significantly more effective than existing GFlowNet training methods.
+Our posterior update rules follow the most progressive methods of ant colony optimization (ACO) (Ye et al.,, 2023), creating synergies between our multi-modal amortized prior and their iterative refinement via parallel stochastic search.
+Additionally, we also validate our method with another posterior search method, active search (AS; Hottung et al.,, 2022), demonstrating the flexibility of our approach.
+Figure 1: Synergy of multi-modal prior distribution of solutions and iterative posterior update with parallel stochastic search in combinatorial optimization.
+Through extensive experiments, we demonstrate the superiority of our training methods over general-purpose reinforcement learning algorithms for providing effective prior distribution on ACO search across seven combinatorial optimization benchmarks:
+  • Routing problems:
+    • Traveling Salesman Problem (TSP)
+    • Capacitated Vehicle Routing Problem (CVRP)
+    • CVRP with Time Windows (CVRPTW)
+    • Prize-Collecting TSP (PCTSP)
+    • Orienteering Problem (OP)
+  • Scheduling problem: Single Machine Total Weighted Tardiness Problem (SMTWTP)
+  • Grouping problem: Bin Packing Problem (BPP)
+For problems such as the Traveling Salesman Problem (TSP) and the Capacitated Vehicle Routing Problem (CVRP), we compare our GFACS method against problem-specific reinforcement learning algorithms that utilize significantly larger and specialized deep networks.
 2 Preliminary
 2.1 Traveling salesman problem
-The traveling salesman problem (TSP) is a fundamental combinatorial optimization problem, which serves as a building block for more complex CO problems like the capacitated vehicle routing problem (CVRP) and orienteering problem (OP).
+The traveling salesman problem (TSP) is a fundamental combinatorial optimization problem, which serves as a building block for more complex CO problems like the capacitated vehicle routing problem and orienteering problem.
 In light of this and for simplicity, this paper presents formulations using the two-dimensional Euclidean TSP as an illustrative example.
-A TSP instance is defined on a (fully connected) graph $\mathcal{G}=\{\mathcal{V},\mathcal{D}\}$, where $\mathcal{V}$ and $\mathcal{D}$ are a node set and edge set, respectively.
+A TSP instance is defined on a (fully connected) graph $\mathcal{G}=(\mathcal{V},\mathcal{D})$, where $\mathcal{V}$ and $\mathcal{D}$ are a set of nodes and edges, respectively.
 Each node $i\in\mathcal{V}$ has two-dimensional coordinates $v_{i}\in[0,1]^{2}$ as a feature, and each edge $(i,j)\in\mathcal{D}$ has a feature $d_{i,j}$ which is the distance between node $i$ and $j$, i.e., $d_{i,j}=\|v_{i}-v_{j}\|_{2}$.
-A solution to TSP, $x(\mathcal{G})$ corresponds to a Hamiltonian cycle for a given problem instance $\mathcal{G}$, which can be represented as a permutation of nodes $\pi=(\pi_{1},\ldots,\pi_{N})$.
-The tour length is computed as $\mathcal{E}(x(\mathcal{G}))=d_{\pi_{N},\pi_{1}}+\sum_{t=1}^{N-1}d_{\pi_{t},\pi_{t+1}}$.
-The optimization objective for TSP is to find $x(\mathcal{G})$ that minimizes $\mathcal{E}(x(\mathcal{G}))$.
+A solution to TSP, $x$ corresponds to a Hamiltonian cycle for a given problem instance $\mathcal{G}$, which can be represented as a permutation of nodes $\pi=(\pi_{1},\ldots,\pi_{N})$.
+The tour length is computed as $\mathcal{E}(x;\mathcal{G})=d_{\pi_{N},\pi_{1}}+\sum_{t=1}^{N-1}d_{\pi_{t},\pi_{t+1}}$.
+The optimization objective for TSP is to find $x$ that minimizes $\mathcal{E}(x;\mathcal{G})$.
+Figure 2: Solution sampling process of GFACS. The GNN, pretraind with a GFlowNet loss, serves as an expert heuristic to pick the next action to construct a solution, such as a tour in the TSP.
+2.2 Generative flow networks
+Generative Flow Networks, or GFlowNets (Bengio et al., 2021a, ; Bengio et al.,, 2023), provide an effective training framework for learning a policy that samples compositional objects $x\in\mathcal{X}$.
+The objective is to generate $x$ with probability proportional to a given non-negative reward function $R(x)$.
+GFlowNets follow a constructive generative process, using discrete actions to iteratively modify a state which represents a partially constructed object.
+All the ways to construct these objects can be described by a directed acyclic graph (DAG), $\mathcal{H}=(\mathcal{S},\mathcal{A})$, where $\mathcal{S}$ is a finite set of all possible states, and $\mathcal{A}$ is a subset of $\mathcal{S}\times\mathcal{S}$, representing directed edges.
+Within this framework, we define the children of state $s\in\mathcal{S}$ as the set of states connected by edges whose head is $s$, and the parents of state $s$ as the set of states connected by edges whose tail is $s$.
+We define a (complete) trajectory $\tau=(s_{0}\rightarrow\ldots\rightarrow s_{N})\in\mathcal{T}$ from the initial state $s_{0}$ to terminal state $s_{n}$.
+We define the forward policy to model the forward transition probability $P_{F}(s^{\prime}|s)$ from $s$ to its child $s^{\prime}$.
+Similarly, we also consider the backward policy $P_{B}(s|s^{\prime})$ for the backward transition $s^{\prime}\dashrightarrow s$, where $s$ is a parent of $s^{\prime}$.
+We can use the forward policy to compute in a forward way the probability of a trajectory: $P_{F}(\tau;\theta)=\prod_{t=1}^{n}P_{F}\left(s_{t}|s_{t-1};\theta\right)$.
+Similarly, we can use the backward policy to compute in a backward way the probability of trajectory ending at a given object $x=s_{n}$: $P_{B}(\tau|x;\theta)=\prod_{t=1}^{n}P_{B}\left(s_{t-1}|s_{t};\theta\right)$.
+Trajectory balance (Malkin et al.,, 2022).
+Trajectory balance (TB) is the most widely used training objective for GFlowNets.
+It works with three key components: a learnable scalar $Z_{\theta}$ for the partition function, a learnable forward policy $P_{F}(\tau;\theta)$, and a backward policy $P_{B}(\tau|x)$ that can either be learnable or fixed with a predefined configuration (Zhang et al., 2022a, ).
+TB loss to minimize for any trajectory $\tau$ is defined as:
+$\mathcal{L}_{\text{TB}}(\tau;\theta)=\left(\log\frac{Z_{\theta}P_{F}(\tau;\theta)}{R(x)P_{B}(\tau|x)}\right)^{2}.$ (1)
+When the trajectory balance loss is minimized to zero for all complete trajectories, the probability of sampling $x$ with the forward policy is $P_{F}^{\top}(x)\propto R(x)$, as desired, where $P_{F}^{\top}(x)=\sum_{(\tau\rightarrow x)\in\mathcal{T}}P_{F}(\tau)$ with $(\tau\rightarrow x)$ indicating that the terminal state of $\tau$ is $x$.
 2.3 Ant colony optimization
-Ant colony optimization is a meta-heuristic to refine the solution sampling probability by an iterative procedure of solution construction and pheromone update.
+Ant colony optimization (ACO; Dorigo et al.,, 2006) is a meta-heuristic to refine the distribution over solution space by an iterative procedure of solution construction and pheromone update.
 Given a problem instance $\mathcal{G}$, a corresponding heuristic matrix $\eta(\mathcal{G})\in\mathbb{R}^{N\times N}$ and a pheromone trails $\rho\in\mathbb{R}^{N\times N}$, a solution is constructed by auto-regressively sampling edges from the probability distribution:
@@
 where $\rho_{i,j}$ and $\eta_{i,j}(\mathcal{G})$ are the pheromone weight and heuristic weight for edge $d_{i,j}$, respectively.
-At each iteration, $K$ artificial ants independently sample solutions $\{\pi^{k}\}_{k=1}^{K}$ by following Equation 1.
-The pheromone trails $\rho$ are then adjusted using the solution set and corresponding costs, thereby updating the solution construction probability.
-The pheromone update rules vary depending on the ACO algorithm used.
-For example, in Ant System (AS) [14], all $K$ solutions contribute to the pheromone update:
+At each iteration, $K$ artificial ants independently sample solutions $\{\pi^{k}\}_{k=1}^{K}$ by following Eq.(2).
+The pheromone trails $\rho$ are then adjusted using the sampled solutions and their corresponding costs, thereby updating the solution distribution.
+The exact pheromone update rules depend on the specific ACO algorithm being used.
+As a representative example, in Ant System (Dorigo et al.,, 1996), all $K$ solutions contribute to the pheromone update as follows:
 $\rho_{i,j}\leftarrow(1-\gamma)\rho_{i,j}+\sum\limits_{k=1}^{K}\Delta^{k}_{i,j},$ (3)
 where $\gamma$ is the decaying coefficient (or evaporation rate) and $\Delta^{k}_{i,j}$ is the amount of pheromone laid by the $k$-th ant on the edge $(i,j)$.
-Here, $\Delta^{k}_{i,j}$ is proportional to the inverse of tour length if $(i,j)$ is included in $\pi^{k}$, and zero otherwise.
-$\Delta^{k}_{i,j}=\begin{cases}C/f(\pi^{k})&\text{if $(i,j)\in\pi^{k}$}\\ 0&\text{otherwise}\end{cases}$ (3)
-where $C$ is a constant. The smaller the cost $f(\pi^{k})$, the more pheromone is laid on the edges of $\pi^{k}$.
-Note that the heuristic matrix $\eta(\mathcal{G})\in\mathbb{R}^{N\times N}$ is not updated during the search process; it only serves as a prior.
-ACO aligns the search process more closely with the most efficacious paths discovered; see Dorigo et al.
-[15] for details.
-In DeepACO [79], the learned graph neural network (GNN), parameterized by $\theta$, is utilized to construct the heuristic matrix $\eta(\mathcal{G},\theta)$ for a given problem instance $\mathcal{G}$, thereby providing a good prior for guiding probabilistic solution search process of ACO.
-Figure 1: Solution sampling process of GFACS. The GNN, pretraind with a GFlowNet loss, serves as an expert heuristic to pick the next action to construct a solution, such as a tour in the TSP.
-3.3 Generative flow networks
-Generative Flow Networks, or GFlowNets [4, 5] provide an effective training framework for learning a policy that samples compositional objects denoted as $x\in\mathcal{X}$ towards the objective of generating $x$ with probability proportional to a given non-negative reward function $R(x)$.
-GFlowNets follow a constructive generative process, using discrete actions to iteratively modify a state which represents a partially constructed object.
-All the ways to construct these objects can be described by a directed acyclic graph (DAG), $\mathcal{H}=(\mathcal{S},\mathcal{A})$, where $\mathcal{S}$ is a finite set of all possible states, and $\mathcal{A}$ is a subset of $\mathcal{S}\times\mathcal{S}$, representing directed edges.
-Within this framework, we define the children of state $s\in\mathcal{S}$ as the set of states connected by edges whose head is $s$, and the parents of state $s$ as the set of states connected by edges whose tail is $s$.
-We define a complete trajectory $\tau=(s_{0}\rightarrow\ldots\rightarrow s_{N})\in\mathcal{T}$ from the initial state $s_{0}$ to terminal state $s_{n}$.
-We denote $\tau_{{}_{\rightarrow}x}$ when the terminal state $s_{n}$ is finalized to the corresponding object $x\in\mathcal{X}$.
-We define the forward policy to model the forward transition probability $P_{F}(s^{\prime}|s)$ from $s$ to its child $s^{\prime}$.
-Similarly, we also consider the backward policy $P_{B}(s|s^{\prime})$ for the backward transition $s^{\prime}\dashrightarrow s$, where $s$ is a parent of $s^{\prime}$.
-We can use the forward policy to compute in a forward way the probability of a trajectory:$P_{F}(\tau;\theta)=\prod_{t=1}^{n}P_{F}\left(s_{t}|s_{t-1};\theta\right)$.
-Similarly, we can use the backward policy to compute in a backward way the probability of trajectory ending at a given object $x=s_{n}$: $P_{B}(\tau_{\rightarrow x}|x;\theta)=\prod_{t=1}^{n}P_{B}\left(s_{t-1}|s_{t};\theta\right)$.
-Trajectory balance [52].
-Trajectory balance is the most widely used training objective for GFlowNets.
-When the trajectory balance objective is minimized, we achieve that the probability of sampling $x$ with the forward policy is $P_{F}^{\top}(x)\propto R(x)$, as desired, where $P_{F}^{\top}(x)=\sum_{\tau_{\rightarrow x}\in\mathcal{T}}P_{F}(\tau)$.
-The trajectory balance loss $\mathcal{L}_{\text{TB}}$ works by training three models, a learnable scalar $Z_{\theta}$ for the partition function, a forward policy $P_{F}(\tau;\theta)$, and a backward policy $P_{B}(\tau_{\rightarrow x}|x;\theta)$ to minimize the following objective:
-$\mathcal{L}_{\text{TB}}(\tau;\theta)=\left(\log\frac{Z_{\theta}P_{F}(\tau;\theta)}{R(x)P_{B}(\tau_{\rightarrow x}|x;\theta)}\right)^{2}.$ (4)
-4 Generative flow ant colony sampler (GFACS)
+The value of $\Delta^{k}_{i,j}$ is proportional to the inverse of tour length if edge $(i,j)$ is part of $\pi^{k}$, and zero otherwise:
+$\Delta^{k}_{i,j}=\begin{cases}C/f(\pi^{k})&\text{if $(i,j)\in\pi^{k}$}\\ 0&\text{otherwise},\end{cases}$ (4)
+where $C$ is a constant.
+The smaller the cost $f(\pi^{k})$, the more pheromone is laid on the edges of $\pi^{k}$.
+This mechanism encourages future ants to follow promising paths that lead to high-quality solutions.
+Note that the heuristic matrix $\eta(\mathcal{G})$ is not updated during the search process; it only serves as an initial prior.
+This prior is augmented with the pheromone $\rho$ that is typically initialized as a constant matrix and then iteratively updated, resulting in the posterior solution distribution that converges to the Dirac-delta distribution of optimal solution under certain conditions (Dorigo and Gambardella,, 1997; Dorigo and Blum,, 2005).
+In DeepACO (Ye et al.,, 2023), the learned graph neural network (GNN), parameterized by $\theta$, is utilized to construct the heuristic matrix $\eta(\mathcal{G},\theta)$ for a given problem instance $\mathcal{G}$, providing an informative prior for guiding probabilistic solution search process of ACO.
+The limitation of the approach is that they use REINFORCE (Williams,, 1992) for GNN training, which suffers from mode-collapse, restricting further improvement from the subsequent iterative pheromone update.
+Our work seeks to overcome this limitation by incorporating a diversified, multi-modal solution distribution, which shares similar motivation with approximate global optimization methods such as simulated annealing (Kirkpatrick et al.,, 1983).
+3 Generative flow ant colony sampler
 In this section, we provide the inference procedure for the Generative Flow Ant Colony Sampler (GFACS).
-GFACS is a sampling algorithm that enhances Ant Colony Optimization (ACO) by integrating it with GFlowNets.
-The central concept is to use GFlowNets to generate an appropriate instance-conditioned prior (edge) distribution (heuristic matrix), which more effectively guides the probabilistic search procedure of ACO.
+GFACS enhances Ant Colony Optimization (ACO) by integrating it with multi-modal prior from GFlowNet-based training.
 3.1 MDP formulation
-The solution construction procedure by GFlowNets can be modeled as a Markov decision process (MDP).
-We illustrate this procedure using the Traveling Salesman Problem (TSP) as an example.
-This formulation can be easily adapted to general combinatorial optimization problem with minor modifications.
+We illustrate the solution construction procedure as a Markov decision process (MDP), using the Traveling Salesman Problem (TSP) as an example.
+This formulation can be easily adapted to general combinatorial optimization problems with minor modifications.
 State $s$: The initial state $s_{0}$ is an empty set.
-The state $s_{t}$ at $0<t<n$ is a partial combinatorial solution of prensented with previous actions: $\{a_{0},\ldots,a_{t-1}\}$ and $s_{n}$ is completed solution.
-Action $a$: The action $a_{t}$ is a selection of a node to visit at time $t$ among un-visited nodes: $a_{t}\in\{1,\ldots,N\}\setminus\{a_{0},\ldots,a_{t-1}\}$.
-If all nodes are visited, a deterministic terminating action is applied, which maps a final state $s_{N}$ into a solution, $x(\mathcal{G})$.
-Reward $R(x(\mathcal{G});\beta)$: The reward is defined as the unnormalized Boltzmann distribution of the solution $x(\mathcal{G})$ with energy $\mathcal{E}(x(\mathcal{G}))$.
-Specifically, the reward is given by $R(x(\mathcal{G});\beta)=e^{-\beta\mathcal{E}(x(\mathcal{G}))}$.
-Here, $\mathcal{E}(x(\mathcal{G}))$ represents the energy of the solution $x(\mathcal{G})$, which corresponds to metrics such as the tour length in the TSP.
-The coefficient $\beta$, known as the inverse temperature, is a hyperparameter.
-4.2 Policy configuration
-We parameterize the forward policy $P_{F}$ using a graph neural network (GNN) with parameter $\theta$ that maps an input graph instance $\mathcal{G}$ to the output edge matrix $\eta(\mathcal{G};\theta)$, following Ye et al.
-[79].
-Then, the forward policy when $t<N$ is expressed as follows:
-$P_{F}(s_{t+1}|s_{t};\eta(\mathcal{G};\theta),\rho)\propto\begin{cases}\rho_{a_{t-1},a_{t}}\eta_{a_{t-1},a_{t}}(\mathcal{G};\theta)&\text{if $a_{t}\notin s_{t}$ }\\ 0&\text{otherwise}\end{cases}$ (5)
-Note that $\rho$ and $\eta(\mathcal{G};\theta)$ denote the pheromone trails and learned heuristic matrix in ACO.
-Subsequently, we can define a trajectory from the initial state to a terminal solution, with $\tau=(s_{0},\ldots,s_{N},x(\mathcal{G}))$, and the forward probability of that trajectory,
-$P_{F}(\tau;\eta(\mathcal{G};\theta),\rho)=\prod_{t=0}^{N-1}P_{F}(s_{t+1}|s_{t};\eta(\mathcal{G};\theta),\rho).$ (6)
-Note that the forward probability for the transition from $s_{N}$ to $x(\mathcal{G})$ is omitted as it follows a deterministic rule.
-For the backward policy, we set $P_{B}=\frac{1}{2N}$, emulating the maximum entropy GFlowNets [83] to make uniform exploration over symmetric trajectories which terminate into identical solution.
-Here, $P_{B}$ represents a uniform distribution across all $2N$ feasible trajectories leading to a solution $x(\mathcal{G})$.
-This is particularly pertinent in the context of the TSP involving $N$ cities, where $2N$ permutations culminate in the same Hamiltonian cycle $x(\mathcal{G})$.
-The count of symmetric trajectories depends on the task’s characteristics.
-For an exhaustive discussion on solution symmetry across various COPs, please refer to Appendix E.
-4.3 Solution sampling with GFACS
+The state $s_{t}$ at $0<t<n$ is a partial combinatorial solution presented with previous actions: $\{a_{0},\ldots,a_{t-1}\}$ and $s_{n}$ is a completed solution.
+Action $a$: The action $a_{t}$ is a selection of the next node to visit at time $t$ among un-visited nodes: $a_{t}\in\{1,\ldots,N\}\setminus\{a_{0},\ldots,a_{t-1}\}$.
+If all nodes are visited, a deterministic terminating action is applied, which maps a final state $s_{N}$ into a solution, $x$.
+Reward $R$: The reward is given by $R(x;\mathcal{G},\beta)=e^{-\beta\mathcal{E}(x;\mathcal{G})}$.
+Here, $\mathcal{E}(x;\mathcal{G})$ represents the energy of the solution $x$, which corresponds to metrics such as the tour length in the TSP.
+The inverse temperature $\beta$ is a hyperparameter.
+For simplicity, we will omit $\mathcal{G}$ from $R$ and $\mathcal{E}$ except when its inclusion is necessary for clarity.
+3.2 Policy configuration for prior modeling
+We parameterize the forward policy $P_{F}$ using a graph neural network (GNN) with parameter $\theta$ that maps an input graph instance $\mathcal{G}$ to the output edge matrix $\eta(\mathcal{G};\theta)$, following Ye et al., (2023).
+For simplicity, we use notation $\eta(\mathcal{G})$ without $\theta$.
+With a given $\eta(\mathcal{G})$, the forward policy when $t<N$ is expressed as follows:
+$P_{F}(s_{t+1}|s_{t};\eta(\mathcal{G}),\rho)\propto\begin{cases}\rho_{a_{t-1},a_{t}}\eta_{a_{t-1},a_{t}}(\mathcal{G})&\text{if $a_{t}\notin s_{t}$ }\\ 0&\text{otherwise}\end{cases}$ (5)
+Note that $\rho$ and $\eta(\mathcal{G})$ denote the pheromone trails and learned heuristic matrix in ACO.
+Subsequently, we can define a trajectory from the initial state to a terminal solution, with $\tau=(s_{0},\ldots,s_{N},x)$, and the forward probability of that trajectory as
+$P_{F}(\tau;\eta(\mathcal{G}),\rho)=\prod_{t=0}^{N-1}P_{F}(s_{t+1}|s_{t};\eta(\mathcal{G}),\rho).$ (6)
+The forward probability for the transition from $s_{N}$ to $x$ is omitted since it is deterministic.
+During training, we fix $\rho$ to $\mathbf{1}$, and thus we omit it from subsequent sections unless its presence is required for clarity.
+For the backward policy, we first define $h(x)$ as the number of possible trajectories that yield the solution $x$, and we set $P_{B}(\tau|x)=\frac{1}{h(x)}$.
+This setting emulates the maximum entropy GFlowNets (Zhang et al., 2022a, ), treating all symmetric trajectories that terminate in the same solution equally.
+In the case of the TSP, where $h(x)=2N$, we have $P_{B}=\frac{1}{2N}$, thereby imposing a uniform backward distribution across all $2N$ feasible trajectories leading to the same Hamiltonian cycle $x$.
+It is important to note that the number of symmetric trajectories varies depending on the specific CO problem; for a more detailed discussion on solution symmetry across various CO problems, please refer to Appendix E
+3.3 ACO update on learned prior
 In the sampling process, we assume the existence of the learned GNN parameterized with $\theta^{*}$; see Section 4 for details of the training procedure.
-The learned GNN provides the heuristic matrix $\eta(\mathcal{G};\theta^{*})$ for ACO algorithm.
-Thus, the remaining process is to update pheromone parameter $\rho$ with solutions sampled by $P_{F}(\tau;\eta(\mathcal{G};\theta^{*}),\rho)$ during the solution searching process.
-As the $\eta(\mathcal{G};\theta^{*})$ is trained with GFlowNets to sample diverse high-rewarded (i.e., low energy) samples, $P_{F}$ generates good candidates for updating $\rho$.
+This learned GNN provides the prior parameter $\eta^{*}(\mathcal{G})$ for the ACO algorithm.
+The next step is to update the pheromone parameter $\rho$, using $K$ agents (ants) to search over samples drawn from the amortized prior $P_{F}(\tau;\eta^{*}(\mathcal{G}))$ as described in Section 2.3.
+Since the trained $\eta^{*}(\mathcal{G})$ samples diverse, high-reward (i.e., low-energy) candidates, the distribution $P_{F}$ provides strong prior candidates for the ants in the ACO algorithm.
+This learned prior is expected to highly improve the likelihood of discovering the global optimum after the iterations of the pheromone update.
 The overall process is illustrated in Figure 2.
@@
 This section introduces how to train GFACS, especially the GNN component parameterized with $\theta$, in the GFlowNet framework.
-GFACS is trained via iterations of the off-policy experience collection (Step A) and training (Step B).
-5.1 Step A: Off-policy experience collection with local search and energy reshaping
-Figure 2: Hindsight observation with local search and energy reshaping to control the trade-off between exploration and exploitation.
+GFACS is trained via iterations of the off-policy experience collection (Section 4.1) and training (Section 4.2).
+4.1 Off-policy experience collection
+Figure 3: Experience collection procedure. Energy reshaping compensates for the energy of underrated samples that have potential to become low-energy samples after the local search.
 In this step, we collect experiences on given graph instances.
-The experiences in step A will be used for training GFACS in step B.
-For every iteration, we first sample $M$ problem instances from a random sampler $P(\mathcal{G})$, i.e., $\mathcal{G}_{1},\ldots,\mathcal{G}_{M}\sim P(\mathcal{G})$.
-Details of instance sampling are described in Appendix B.
-For each problem instance $\mathcal{G}_{m}$, we collect $K$ solutions by following Eq.
-6 as $\mathcal{B}_{\text{explore}}=\{(\mathcal{G}_{m},\tau^{m,k}_{\rightarrow x_{m,k}},\mathcal{E}(x_{m,k})\}_{m=1,\ldots,M,k=1,\ldots,K}$.
-To exploit high-reward regions in the large-scale combinatorial space, we also collect $K$ exploitation samples for each $\mathcal{G}_{m}$ using a local search operator ($LS$), such as 2-opt for TSP, as a powerful behavioral policy.
-The local search operator directly refines a given solution into a locally optimized one, i.e., $x^{\prime}\leftarrow LS(x)$.
-We then use the backward policy $P_{B}$ to sample a trajectory $\tau_{\rightarrow x^{\prime}}$ for each $x^{\prime}$, obtaining a batch of “exploitation experiences,” $\mathcal{B}_{\text{exploit}}=\{(\mathcal{G}_{m},\tau^{m,k}_{\rightarrow x^{\prime}_{m,k}},\mathcal{E}(x^{\prime}_{m,k})\}_{m=1,\ldots,M,k=1,\ldots,K}$.
-Please note that we also suggest a general-purpose local search operator that partially destroys a solution and then repairs for problems that lack domain-specific local search methods.
-Please refer to Appendix C for details about the destroy-and-repair local saerch.
-While the concept of using local search for exploitation has been explored in recent GFlowNet works [39], our distinct approach involves utilizing the local search (LS) operation to facilitate hindsight observation—a forward-looking insight derived from extending the search beyond the current state.
-We focus on evaluating samples that were initially overlooked due to their high energy (i.e., low reward) but can be transformed to exhibit low energy (i.e., high reward) following the local search.
-To achieve this, we modify the energy (in a way we call energy reshaping) for the exploration around $x$ in $\mathcal{B}_{\text{explore}}$ by using the energy of the refined solution $x^{\prime}=LS(x)$ as follows:
-$\mathcal{E}_{\text{reshaped}}(x)=\alpha\mathcal{E}(x^{\prime})+(1-\alpha)\mathcal{E}(x).$ (7)
-5.2 Step B: Training with collected experiences with shared energy normalization
-Given the collected experience batches $\mathcal{B}_{\text{explore}}$ and $\mathcal{B}_{\text{exploit}}$, we aim to minimize the following loss function:
-$\mathcal{L}(\theta)=\mathcal{L}(\theta;\mathcal{B}_{\text{explore}})+\mathcal{L}(\theta;\mathcal{B}_{\text{exploit}}),$ (8)
-where ($\mbox{where }\mathcal{L}(\theta;\mathcal{B})=\sum_{m=1}^{M}\sum_{k=1}^{K}\left(\log\frac{2NP_{F}(\tau^{m,k}_{\rightarrow x_{m,k}};\eta(\mathcal{G}_{m};\theta),\rho)Z_{\theta}(\mathcal{G}_{m})}{R(x_{m,k};\beta)}\right)^{2}.$ (9)
-The $2N$ comes from the fact that TSP has $2N$ permutations, as discussed in Section 4.2.
-In addition, our target sampling distribution is proportional to reward $p(x;\mathcal{G})\propto R(x(\mathcal{G}))$.
-However, in most COPs, including TSP, the scale of $R(x(\mathcal{G}))$ can vary significantly across instances, possibly leading to unstable training due to the high variance of logarithms of rewards.
-Similarly to previous works that stabilize REINFORCE training [43, 37], we normalize the energy $\mathcal{E}=-\log R$ using the sample mean over $K$ samples for the same $\mathcal{G}_{m}$, i.e.,
-$-\log\tilde{R}(x_{m,k})=\tilde{\mathcal{E}}(x_{m,k})=\mathcal{E}(x_{m,k})-\frac{1}{K}\sum_{k=1}^{K}\mathcal{E}(x_{m,k}).$ (10)
-We call this technique a shared energy normalization.
-Note that the shared energy normalization is applied independently to the two batches, $\mathcal{B}_{\text{explore}}$ and $\mathcal{B}_{\text{exploit}}$, as the energy distribution between the two batches is often significantly different from each other, especially when a heuristic local search is applied.
-To this end, the gradient loss is computed as follows:
-$\nabla_{\theta}\mathcal{L(\theta)}\approx\nabla_{\theta}\sum_{m=1}^{2M}\sum_{i=1}^{I}\left(\log\frac{2NP_{F}(\tau^{i,m}_{\rightarrow x_{i,m}};\eta(\mathcal{G}_{m};\theta),\rho)Z_{\theta}(\mathcal{G}_{m};\theta)}{\tilde{R}(x_{i,m})}\right)^{2}.$ (11)
-6 Experiments
-This section presents experimental results to validate our algorithm [^1] Source code is available: https://github.com/ai4co/gfacs.
-.
-We compare GFACS with other ACO algorithms, and with competitive reinforcement learning algorithms for vehicle routing problems.
-Moreover, we conduct an ablation study, and perform an empirical comparison with previous GFlowNets training methods, which serve a similar purpose to our method.
-Below, we provide a brief overview of the experimental setup.
-Implementation and hyperparameters.
-GFACS is built upon the implementation established by DeepACO [79].
-As our methodology mainly focuses on improving training algorithms rather than network architecture, we simply adopt the GNN architecture proposed in DeepACO, except the additional parameters for partition function $Z_{\theta}$ in Equation 9.
-For detailed network architecture, hyperparameters and hardware settings, please refer to Appendix A.
-It is worth noting that one important hyperparameter that GFACS newly introduces is the inverse energy temperature $\beta$, which enables us to control the diversity-optimality trade-offs (see Section D.2 for detailed analysis).
-In addition, we employ Ant System [14] for the pheromone updates throughout the experiments, while we also provide experimental results with other ACO variants in Section D.3.
-Test settings.
-Each algorithm is tested on the held-out datasets, each with 100 instances generated in advance except TSP, where we obtain the TSP test dataset with 128 instances from Ye et al.
-[79].
-Unless otherwise stated, the results presented throughout this section are the averaged values from three independent models, each trained with a distinct random seed.
-6.1 Comparison with ACO algorithms
-Here, we compare GFACS against the vanilla ACO with and the DeepACO [79] algorithm across various CO tasks.
-Problems.
-Following the approach of Ye et al.
-[79], we evaluate our algorithm in a range of COPs.
-This includes four routing problems – TSP, Capacitated Vehicle Routing Problem (CVRP), CVRP with Time Window (CVRPTW), Prize-Collecting Vehicle Routing Problem (PCTSP), and Orienteering Problem (OP) – along with a scheduling problem – Single Machine Total Weighted Tardiness Problem (SMTWTP) – and a grouping problem, the Bin Packing Problem (BPP).
-For detailed problem configurations, we direct readers to Appendix B.
-For TSP, CVRP and CVRPTW, we use local search algorithms with perturbation similar to Ye et al.
-[79].
-For others, we use the destroy-and-repair method (see Appendix C for more details).
-Figure 3: Results of ACO algorithms on various CO tasks.
+The collected experiences will be used to train GFACS, as described in Section 4.2.
+At each training round, we first sample $M$ problem instances from a random instance generator $P(\mathcal{G})$, i.e., $\mathcal{G}_{1},\ldots,\mathcal{G}_{M}\sim P(\mathcal{G})$.
+Details of instance generation are described in Appendix B.
+For each problem instance $\mathcal{G}_{m}$, we collect $K$ solutions by following Eq.(6), resulting in a batch of “exploration experiences,”
+$\mathcal{B}_{\text{explore}}=\left\{\left(\mathcal{G}_{m},\tau_{m,k},\mathcal{E}(x_{m,k})\right)\mid m\in[M],k\in[K]\right\}.$
+To exploit high-reward regions in the large-scale combinatorial space, we collect additional $K$ experiences for each $\mathcal{G}_{m}$ using a local search (LS) operator, such as 2-opt for TSP, as a powerful behavior policy.
+The LS operator directly refines a given solution into a locally optimized one, i.e., $\hat{x}\leftarrow\text{LS}(x)$.
+We then use the backward policy $P_{B}$ to sample a trajectory $\hat{\tau}$ for each $\hat{x}$, obtaining a batch of “exploitation experiences,”
+$\mathcal{B}_{\text{exploit}}=\left\{\left(\mathcal{G}_{m},\hat{\tau}_{m,k},\mathcal{E}(\hat{x}_{m,k})\right)\mid m\in[M],k\in[K]\right\}.$
+For problems that lack domain-specific LS operators, we also suggest a general-purpose destroy-and-repair method that partially destroys a solution and then repairs; see Appendix C for more details.
+While the concept of using local search for exploitation has been explored in recent GFlowNet works (Kim et al., 2024b, ), our distinct approach involves utilizing the local search operation to facilitate hindsight observation—a forward-looking insight derived from extending the search beyond the current state.
+Methodologically, we compensate for the solutions that were initially overlooked due to high energy (i.e., low reward) but that can be readily transformed into low energy (i.e., high reward) solutions through local search.
+To achieve this, we modify the energy (in a way we call energy reshaping) of $x$ in the exploration experiences, $\mathcal{B}_{\text{explore}}$.
+Specifically, for a given solution $x\in\mathcal{B}_{\text{explore}}$ and its corresponding refined solution $\hat{x}\leftarrow\text{LS}(x)$, we adjust the energy of $x$ as follows:
+$\mathcal{E}^{\prime}(x)=\alpha\mathcal{E}(\hat{x})+(1-\alpha)\mathcal{E}(x).$ (7)
+This reshaping captures the potential for samples to become low-energy solutions after local search.
+The overall procedure of experience collection via sampling and local search, along with the energy reshaping, is illustrated in Figure 3.
+4.2 Training with collected experiences
+Given a batch of experiences $\mathcal{B}$, the trajectory balance (TB) loss in Eq.(1) can be rewritten as
+$L(\mathcal{B})=\frac{1}{MK}\sum_{m=1}^{M}\sum_{k=1}^{K}\left(\log\frac{P_{F}(\tau_{m,k};\eta(\mathcal{G}_{m}))Z(\mathcal{G}_{m})}{(1/h(x_{m,k}))R(x_{m,k};\mathcal{G}_{m},\beta)}\right)^{2},$ (8)
+where $1/h(x_{m,k})$ is to the uniform backward policy as described in Section 3.2.
+Since the loss is computed over multiple instances $\mathcal{G}_{m}$, the objective corresponds to the conditional trajectory balance.
+Figure 4: Results of ACO algorithms with different priors on various CO tasks.
 Our GFACS outperforms every ACO baseline.
 The results are averaged over 3 independent models evaluated on the held-out test sets, and the shade indicates the min-max range of the 3 models.
-Baselines.
-We use vanilla ACO with a human-designed heuristic matrix, DeepACO as baseline algorithms.
-We also test GFACS without the proposed off-policy training scheme (called On-policy TB).
-For inference, we use $K=100$ ants and $T=100$ ACO iterations for all the algorithms.
-Results.
-Figure 3 show GFACS consistently outperforms both the classic ACO algorithm and DeepACO across all benchmark problems.
-Notably, for routing problems, GFACS surpasses DeepACO by a significant margin.
-This is particularly remarkable given that both algorithms use identical neural networks and local search operators, underscoring the importance of our training strategy.
-Additionally, the proposed off-policy training method enhances GFACS performance, demonstrating its effectiveness for GFlowNets.
-6.2 Comparison with other reinforcement learning methods for combinatorial optimization
-In this section, we present experimental results comparing our method with reinforcement learning (RL)-based approaches, as both aim to develop combinatorial solvers without the need for labeled optimal data points, unlike supervised learning methods (e.g., [16, 70]).
-Baselines.
-For TSP baselines, we select the AM [41], POMO [43], Sym-NCO [37] and Pointerformer [28] as RL-based constructive methods, DIMES [62] as RL-based heatmap-based methods, the Select and Optimize (SO) [9] as RL-based improvement methods.
-We also applied two search algorithms, the Efficient Active Search (EAS) [23] and Simulation Guided Beam Search (SGBS) [10], to POMO for a more thorough comparison.
-For CVRP, [^2] We also adopt TAM as a CVRP baseline; however, due to the inaccessibility of source code, we compare it using the reported result in the paper [24].
-See Section D.5.
-we employ POMO and Sym-NCO as baselines with EAS and SGBS.
-Note that we restrict the duration of these searches to no more than 5 and 10 minutes for problems with 500 and 1000 nodes, respectively.
-We use Concorde [2] for TSP and PyVRP [77] for CVRP as oracle heuristic solvers to obtain near-optimal solutions that serve as baselines with $0.00\%$ gap.
-For ACO algorithms, we use 100 ants with $T=10$ with local search algorithms similar to Ye et al.
-[79].
-Results.
-Among the baseline methods evaluated, our GFACS algorithm consistently delivers either superior or highly competitive performances.
-When compared to faster methods like AM or POMO, GFACS demonstrates substantially better results.
-Even after an additional search with EAS or SGBS, they fail to match GFACS’s performance.
-Compared to algorithms like DIMES and Pointformer, which operate at similar speeds, GFACS achieves improved performance.
-GFACS performs slightly worse than the oracle heuristic solver LKH-3, but offers an advantage of faster inference, especially in CVRP.
-Therefore, we can conclude that GFACS, as a general-purpose algorithm, holds significant appeal in the highly competitive benchmarks of both TSP and CVRP.
+As discussed in Section 2.2, our target sampling distribution is proportional to reward, i.e., $p(x;\mathcal{G})\propto R(x;\mathcal{G},\beta)$.
+However, in most CO problems, including TSP, the scale of $R(x;\mathcal{G},\beta)$ can vary significantly across instances, possibly leading to unstable training due to the high variance of logarithms of rewards.
+Similar to previous works that stabilize REINFORCE training (Kwon et al.,, 2020; Kim et al.,, 2022), we normalize the energy $\mathcal{E}$ using the sample mean over $K$ samples for the same $\mathcal{G}_{m}$ as follows:
+~$\tilde{\mathcal{E}}(x_{m,k};\mathcal{G}_{m})=\mathcal{E}(x_{m,k};\mathcal{G}_{m})-\frac{1}{K}\sum_{k=1}^{K}\mathcal{E}(x_{m,k};\mathcal{G}_{m}).$ (9)
+Accordingly, we have $\tilde{R}(x_{m,k};\mathcal{G}_{m},\beta)=e^{-\beta\tilde{\mathcal{E}}(x_{m,k};\mathcal{G}_{m})}$.
+We call this technique a shared energy normalization.
+Note that this technique is applied independently to the two batches, $\mathcal{B}_{\text{explore}}$ and $\mathcal{B}_{\text{exploit}}$, as the energy distribution between the two batches is often significantly different from each other, especially when a heuristic local search is applied.
+By applying shared energy normalization to both batches, we obtain ~$\tilde{\mathcal{B}}_{\text{explore}}$ and ~$\tilde{\mathcal{B}}_{\text{exploit}}$, in which every $\mathcal{E}$ is replaced by $\tilde{\mathcal{E}}$.
+The final loss to minimize is then defined as
+$\mathcal{L}(\theta):=\frac{1}{2}L(\mathcal{B}_{\text{explore}})+\frac{1}{2}L(\mathcal{B}_{\text{exploit}}),$ (10)
+where $L$ is as defined in Eq.(8).
+Note the parameters $\theta$ govern the functions $Z(\mathcal{G}_{m})$ and $P_{F}(\cdot;\eta(\mathcal{G}_{m}))$, although this dependence has been omitted for simplicity.
+The gradient of the loss with respect to $\theta$ is given by $\nabla_{\theta}\mathcal{L}(\theta)$, and the parameters are optimized using stochastic gradient descent methods.
+5 Experiments
+This section presents experimental results to validate our algorithm.
+We compare GFACS against other ACO algorithms with different priors and also against competitive reinforcement learning algorithms for vehicle routing problems.
+Moreover, we conducted an ablation study and an empirical comparison with previous GFlowNets training methods.
+Last but not least, we validate the efficacy of multi-modal prior with another posterior update method beyond ACO, the active search (AS).
+Below, we provide a brief overview of the experimental setup.
+Implementation.
+GFACS is built upon the implementation established by DeepACO (Ye et al.,, 2023).
+As our methodology mainly focuses on improving training algorithms rather than network architecture, we simply adopt the GNN architecture proposed in DeepACO, except the additional parameters for partition function $Z(\cdot;\theta)$ in Eq.(8).
+For detailed network architecture, hyperparameters, and hardware settings, please refer to Appendix A.
+It is worth noting that one important hyperparameter that GFACS newly introduces is the inverse energy temperature $\beta$, which enables us to control the diversity-optimality trade-offs (see Section D.2 for detailed analysis).
+In addition, we employ Ant System (Section 2.3) for the pheromone updates throughout the experiments.
+We provide experimental results with other ACO variants in Section D.3.
+Test settings.
+Each algorithm is tested on the held-out datasets, each with 100 instances generated in advance except TSP, where we obtain the TSP test dataset with 128 instances from Ye et al., (2023).
+Unless otherwise stated, the results presented throughout this section are the averaged values from three independent models, each trained with a distinct seed.
+Table 1: Experimental results on real-world datasets, TSPLib and CVRPLib.
+‘$N$’ denotes the range of instance size and ‘#’ denotes the number of instances in each set.
+We report the average optimality gap using the best-known costs.
+$N$ ACO DeepACO GFACS (ours)
+TSPLib 100-299 30 1.71% 1.25% 1.21%
+300-699 10 4.26% 2.70% 2.60%
+700-1499 12 7.01% 4.08% 3.88%
+CVRPLib 100-299 43 2.50% 3.27% 2.69%
+300-699 40 3.75% 4.22% 3.71%
+700-1001 17 4.64% 5.07% 4.32%
 Table 2: Experimental results on TSP and CVRP.
-Results for methods with * are drawn from Ye et al.
-[79], Cheng et al.
-[8], and Jin et al.
-[28].
+Results for methods with * are drawn from Ye et al., (2023), Cheng et al., (2023), and Jin et al., (2023).
 The reported objective value is averaged over 128 instances of TSP and 100 instances of CVRP.
@@
     +SGBS 18.86 13.98 318s 28.58 23.61 605s     +SGBS 65.44 3.93 5.2m 127.90 7.21 10.1m
-DIMES* 17.01 2.78 11s 24.45 5.75 32s SymNCO 68.81 9.28 2s 141.82 18.86 3s
+DIMES* 17.01 2.78 11s 24.45 5.75 32s Sym-NCO 68.81 9.28 2s 141.82 18.86 3s
 SO* 16.94 2.40 15s 23.77 2.80 26s     +EAS 64.63 2.65 5.0m 125.58 5.27 10.0m
@@
 GFACS (ours) 16.80 1.56 15s 23.72 2.63 66s GFACS (ours) 64.19 1.95 36s 121.80 2.11 1.3m
-Table 2: Experimental results on real-world datasets, TSPLib and CVRPLib-X. ‘#’ denotes the number of instances in each set. We report the average optimality gap using the best-known costs.
-TSPLib # ACO DeepACO Ours CVRPLib-X # ACO DeepACO Ours
-100-299 30 1.71% 1.25% 1.21% 100-299 43 2.50% 3.27% 2.69%
-300-699 10 4.26% 2.70% 2.60% 300-699 40 3.75% 4.22% 3.71%
-700-1499 12 7.01% 4.08% 3.88% 700-1001 17 4.64% 5.07% 4.32%
+5.1 Comprison with different priors
+Baselines.
+We compare GFACS against the vanilla ACO (heuristic prior) and the DeepACO (Ye et al.,, 2023) (on-policy RL prior) and ACO with TB prior as baseline methods.
+We aim to show our multi-modal prior from our off-policy GFlowNets training methods is beneficial when this is integrated with the iterative pheromone update of ACO.
+We use $K=100$ ants and $T=100$ ACO iterations for all algorithms.
+Problems.
+We evaluate our algorithm in a range of CO problems, including five routing problems – TSP, Capacitated Vehicle Routing Problem (CVRP), CVRP with Time Window (CVRPTW), Prize-Collecting TSP (PCTSP), and Orienteering Problem (OP) –, a scheduling problem, the Single Machine Total Weighted Tardiness Problem (SMTWTP), and a grouping problem, the Bin Packing Problem (BPP).
+Detailed descriptions and configurations for each problem can be found in Appendix B.
+For TSP, CVRP and CVRPTW, we use local search heuristics (2-opt, Swap* and LocalSearch operator in PyVRP (Wouda et al.,, 2024), respectively) with perturbation scheme similar to Ye et al., (2023) for training GFACS and testing all the ACO algorithms.
+For others, we use the destroy-and-repair method.
+In Appendix C, further details are provided.
+Results.
+Figure 4 shows GFACS consistently outperforms both the classic ACO algorithm and DeepACO across all benchmark problems.
+Notably, for routing problems, GFACS surpasses DeepACO by a significant margin.
+This is particularly remarkable given that both algorithms use identical neural networks and local search operators, underscoring the importance of the multi-modal prior.
+Additionally, the proposed off-policy training method enhances GFACS performance, demonstrating its effectiveness for GFlowNets.
 Evaluation on real-world datasets.
-We compare the performance of our model against baseline models on real-world TSP and CVRP instances from the TSPLib [64] and CVRPLib [72] X datasets.
+We compare the performance of our model against baseline models on real-world TSP and CVRP instances from the TSPLib (Reinelt,, 1991) and CVRPLib X-instances (Uchoa et al.,, 2017).
 The models were trained on random uniform instances of TSP/CVRP with sizes 200 for dataset sizes 100-299, 500 for sizes 300-699, and 1000 for sizes 700-1499, with all other settings remaining the same.
@@
 As the problem scale increases, ACO’s performance declines, underscoring the importance of effective priors using GFlowNets for efficient pheromone posterior updates in large combinatorial spaces.
+5.2 Comparison with RL solvers
+In this section, we present experimental results comparing our method with problem-specific RL approaches that leveraged specialized neural architectures or learning heuristics for the TSP and CVRP.
+Baselines.
+For TSP baselines, we select the AM (Kool et al.,, 2019), POMO (Kwon et al.,, 2020), Sym-NCO (Kim et al.,, 2022) and Pointerformer (Jin et al.,, 2023) as RL-based constructive methods, DIMES (Qiu et al.,, 2022) as RL-based heatmap-based methods, the Select and Optimize (SO; Cheng et al.,, 2023) as RL-based improvement methods.
+We also applied two search algorithms, the Efficient Active Search (EAS; Hottung et al.,, 2022) and Simulation Guided Beam Search (SGBS; Choo et al.,, 2022), to POMO for a more thorough comparison.
+For CVRP, [^1] Due to the inaccessibility of source codes, we compare ours with TAM using the reported result in the paper (Hou et al.,, 2023).
+See Section D.4.
+we employ POMO and Sym-NCO as baselines with EAS and SGBS.
+Note that we restrict the duration of these searches to no more than 5 and 10 minutes for problems with 500 and 1000 nodes, respectively.
+We use Concorde (Applegate et al.,, 2006) for TSP and PyVRP (Wouda et al.,, 2024) for CVRP as oracle heuristic solvers to obtain near-optimal solutions that serve as baselines with $0.00\%$ gap.
+For ACO algorithms, we use $K=100$ ants with $T=10$ with local search algorithms similar to Ye et al., (2023).
+Results.
+GFACS consistently shows superior or highly competitive performance compared to evaluated baseline methods (Table 2).
+Ours outperforms fast zero-shot methods like AM and POMO—even when they employ search techniques such as EAS or SGBS—and achieves better results than similarly paced algorithms like DIMES and Pointformer.
+Although it performs slightly worse than the heuristic solver LKH-3 (Helsgaun,, 2017), GFACS offers faster inference times, especially in CVRP.
+As a general-purpose algorithm, it holds significant appeal in the competitive benchmarks of both TSP and CVRP.
+Notably, despite its small neural architecture and requiring less than three hours of training, GFACS outperforms larger models like POMO and Sym-NCO, which need over a week of training (Kim et al.,, 2022).
 5.3 Ablation study
-Table 3: Results of the ablation study for the proposed training techniques.
+Table 3: Results of the ablation study.
 Method TSP500 CVRP500 BPP120
 GFACS 16.80 64.19 51.84
-   – Off-policy Training 17.09 64.42 52.32
+   – Off-policy 17.09 64.42 52.32
    – Energy Reshaping 16.87 64.41 52.07
-   – Shared $\varepsilon$-norm 17.52 65.51 90.73
-    + VarGrad [82] 16.86 64.28 51.95
+   – Shared Energy Norm. 17.52 65.51 90.73
+    + VarGrad 16.86 64.28 51.95
 This section reports the ablation study results, validating our proposed training techniques for stabilizing and enhancing GFlowNets in large-scale combinatorial optimization tasks.
 As detailed in Table 3, our off-policy training method, incorporating local search, energy reshaping to value potential solutions, and shared $\epsilon$-norm to stabilize conditional GFlowNets training, significantly improves performance across three tasks: TSP, CVRP, and BPP.
-For comprehensive results, please refer to Table 10.
-6.4 Comparative analysis with previous GFlowNets training methods
-While this paper proposes a novel training method for GFlowNets designed for large-scale combinatorial optimization problems that are not locally decomposable due to their cyclic nature (e.g., TSP, CVRP) [1], we also compare existing GFlowNets training methods within this domain.
-Figure 4: Validation costs during training with different GFlowNet losses.
-Forward-looking detailed balance (FL-DB) [59, 84].
+For comprehensive results, including standard deviations, refer to Section D.5.
+5.4 Comparison with other GFlowNet-based training methods
+While this paper proposes a novel training method for GFlowNets designed for large-scale combinatorial optimization problems that are not locally decomposable due to their cyclic nature, such as in TSP and CVRP, we also compare existing GFlowNets training methods within this domain.
+Figure 5: Validation cost on TSP with 200 nodes during training, compared to forward-looking detailed balance (FL-DB).
+Forward-looking detailed balance (FL-DB; Pan et al.,, 2023; Zhang et al.,, 2023).
 FL-DB leverages local credits when the application benefits from useful local rewards for partial solutions.
-While graph combinatorial optimization tasks are generally locally decomposable [1], yielding useful local rewards, tasks like TSP do not offer substantial local rewards due to their cyclic nature, which imposes a global constraint, thus complicating local credit assignment.
-As shown in Figure 4, various GFlowNet methods were used to train the ACO’s heuristic matrix $\eta(\mathcal{G};\theta)$.
+While graph combinatorial optimization tasks are generally locally decomposable (Ahn et al.,, 2020), yielding useful local rewards, tasks like TSP do not offer substantial local rewards due to their cyclic nature, which imposes a global constraint, thus complicating local credit assignment.
+As shown in Figure 5, various GFlowNet methods were used to train the ACO’s heuristic matrix $\eta(\mathcal{G})$.
 While FL-DB did not surpass on-policy TB for TSP, our off-policy method demonstrated significant improvement.
-This trend holds across different deep learning models and inference tasks beyond ACO (see Section D.1).
-Gradient of log-parition variance (VarGrad) [65, 82].
-The VarGrad loss, initially proposed by Richter et al.
-[65] for variational inference, was applied to conditional GFlowNets for solving scheduling problems by Zhang et al.
-[82].
-This approach is crucial for combinatorial optimization tasks that require conditional inference on problem instances ($\mathcal{G}$).
-The key idea involves using a generative distribution and energy to compute the log-partition function through the consistency property of log-partitions over trajectories conditioned on identical instances ($\mathcal{G}$), thereby bypassing the explicit training of the (conditional) log-partition ($\log Z(\mathcal{G})$).
-Explicit training of the log-partition is challenging in conditional GFlowNets due to the varying energy landscapes across different instance conditions ($\mathcal{G}$).
-On the other hand, our method, which utilizes explicit training of the log-partition (see Equation 9), outperforms VarGrad in training conditional GFlowNets (see Table 3).
-This improvement is achieved because our Share $\epsilon$-norms enable stable and explicit training of the log-partition by normalizing energy values into sample means, resulting in an energy distribution across different instances that maintains a mean of zero.
-Note that accurately estimated the explicit log-partition value is known to stabilize the amortized inference process in GFlowNets [53].
-7 Discussion
-This paper proposes a novel probabilistic search method leveraging the off-policy training capabilities of GFlowNets.
-This strategy can be applied not only to ant-colony optimization (ACO) but also to other probabilistic methods requiring an effective learned prior (see Section D.4).
-Our off-policy framework enables the integration of learning algorithms with offline datasets containing optimal combinatorial solutions for specific problem instances, combined with online exploration to fine-tune the policy through off-to-on training.
-This approach holds promise for future work.
-Limitation.
-A limitation of our method is that its performance depends on the effectiveness of the local search.
-A promising future direction could involve leveraging advanced neural-guided improvement heuristics, including neural local search approaches [48, 50], to create a hybrid algorithm that combines GFlowNets with neural improvement-based local search.
+This trend holds across different deep learning models and inference tasks beyond ACO; see Section D.1.
+Gradient of log-partition variance (VarGrad).
+Originally proposed by Richter et al., (2020) for variational inference, the VarGrad loss was applied to conditional GFlowNets for scheduling problems by Zhang et al., 2022b .
+This approach is vital for combinatorial optimization tasks requiring conditional inference on the problem instance $\mathcal{G}$.
+It computes the log-partition function via the consistency property over trajectories conditioned on identical instances, thus bypassing explicit training of $\log Z(\mathcal{G})$.
+Explicit training is challenging due to varying energy landscapes across different instances.
+Our method, utilizing explicit training of the log-partition $\log Z(\mathcal{G})$ as in Eq.(8), outperforms VarGrad in training conditional GFlowNets (Table 3).
+This improvement stems from our Shared energy normalization, which stabilizes training by normalizing energy values to sample means, resulting in an energy distribution with a zero mean across instances.
+Note that accurately estimating the explicit log-partition value is known to stabilize training of GFlowNets (Malkin et al.,, 2023).
+5.5 Integrating GFACS with active search
+Background.
+In this section, we explore the integration of active search (AS) as a posterior update on a parameterized prior distribution $p_{\theta}(\cdot|\mathcal{G})$.
+In AS, the prior parameter $\theta$, pretrained on a large set of random problem instances $\mathcal{G}$, is adaptively updated for specific test-time instances $\mathcal{G}_{\text{test}}$ using the same objective function as in the prior training.
+We trained this prior $p_{\theta}$ using two methods: on-policy RL method (Kwon et al.,, 2020) that employ REINFORCE with shared baselines, and our off-policy GFlowNet training method as explained in Section 4.
+Setting. For the active search procedure, we sample 20 solutions per test-time instance and perform 5 gradient updates at each iteration, updating $\theta$ over a total of 50 iterations.
+Table 4: Evaluation results combine with active search (AS) as a posterior search method.
+TSP200 CVRP200
+Method Obj. Gap (%) Obj. Gap (%)
+Concorde / PyVRP 10.72 0.00 27.93 0.00
+RL prior + AS 10.79 $\pm$ 0.00 0.65 28.43 $\pm$ 0.02 1.79
+GFlowNet prior + AS 10.75 $\pm$ 0.00 0.29 28.34 $\pm$ 0.01 1.47
+Results.
+As shown in Table 4, the off-policy GFlowNet prior combined with AS posterior inference outperforms the RL counterpart.
+This result provides further evidence supporting our hypothesis that the multi-modal prior from GFlowNet training synergizes effectively with the subsequent iterative posterior update algorithm.
+Additionally, the results demonstrate the flexibility of our method beyond the ACO, suggesting a promising direction for future research into the benefits of using a multi-modal informative prior.
+6 Conclusion
+In this paper, we introduced a novel probabilistic meta-heuristic that hierarchically combines GFlowNets-based amortized multi-modal priors with posterior search using ant colony optimization (ACO) for combinatorial optimization problems.
+Our strategy achieved remarkable success across seven benchmark combinatorial optimization tasks.
+Furthermore, by applying our meta-heuristic to other posterior search methods like active search (AS), we demonstrated the expandability and modularity of our approach.
+Limitation of our work is that, while our method empirically improves solution quality, we do not provide formal theoretical guarantees in how our method gives better optimality than the reinforcement learning counterpart.
+Acknowledgment
+The authors thank Haoran Ye and Federico Berto for valuable discussions for this project.
+This research is based on funding from Samsung, Intel, CIFAR and the CIFAR AI Chair program.
+The research was enabled in part by computational resources provided by the Digital Research Alliance of Canada (https://alliancecan.ca), Mila (https://mila.quebec), and NVIDIA.
+This work was also partially supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No.
+RS-2024-00410082).
 Appendix A Implementation
 A.1 Learning architecture and optimizer
-We utilize the anisotropic graph neural network with edge gating mechanisms, which is proposed by [29] and adopted by [79], as the foundational GNN architecture for the construction of a heuristic matrix.
-This network offers node and edge embeddings that are more beneficial for edge prediction than other popular GNN variants, such as Graph Attention Networks [73] or Graph Convolutional Networks [40].
+We utilize the anisotropic graph neural network with edge gating mechanisms, which is proposed by Joshi et al., (2021) and adopted by Ye et al., (2023), as the foundational GNN architecture for the construction of a heuristic matrix.
+This network offers node and edge embeddings that are more beneficial for edge prediction than other popular GNN variants, such as Graph Attention Networks (Veličković et al.,, 2018) or Graph Convolutional Networks (Kipf and Welling,, 2017).
 We represent the node feature of node $v_{i}$ at the $\ell$-th layer as $h_{i}^{\ell}$, and the feature of the edge between nodes $i$ and $j$ at the same layer as $e_{ij}^{\ell}$.
@@
 $e_{ij}^{\ell+1}=e_{ij}^{\ell}+\text{ACT}(\text{BN}(m^{\ell}))$ (13)
-where $U^{\ell},V^{\ell},P^{\ell},Q^{\ell},R^{\ell}\in R^{d\times d}$ are the learnable parameters of layer $\ell$, ACT denotes the activation function, BN[25] denotes batch normalization, $\mathcal{A}$ denote the aggregation function, $\sigma$ is the sigmoid function, $\odot$ is Hadamard product, $\mathcal{N}_{i}$ denotes the neighborhoods of node $v_{i}$.
-In this paper we use $\alpha$ is SiLU[17] and $\mathcal{A}$ is mean pooling.
+where $U^{\ell},V^{\ell},P^{\ell},Q^{\ell},R^{\ell}\in R^{d\times d}$ are the learnable parameters of layer $\ell$, ACT denotes the activation function, BN (Ioffe and Szegedy,, 2015) denotes batch normalization, $\mathcal{A}$ denote the aggregation function, $\sigma$ is the sigmoid function, $\odot$ is Hadamard product, $\mathcal{N}_{i}$ denotes the neighborhoods of node $v_{i}$.
+In this paper we use ACT is SiLU (Elfwing et al.,, 2018) and $\mathcal{A}$ is mean pooling.
 We transform the extracted edge features into real-valued heuristic metrics by employing a 3-layer Multilayer Perceptron (MLP) that includes skipping connections to the embedded node feature $h$.
 The SiLU is the activation function for all layers except the output layer, in which the sigmoid function is applied to generate normalized outputs.
-The training process is optimized using the AdamW optimizer[46] in conjunction with a Cosine Annealing Scheduler[45]
+The training process is optimized using the AdamW optimizer (Loshchilov and Hutter, 2017a, ) in conjunction with a Cosine Annealing Scheduler (Loshchilov and Hutter, 2017b, )
 A.2 Hyperparameters
@@
 We report the $\beta$ value used for each task in Table 6. Note that $\beta$ remains consistent across different scales of the same problem class.
-For TSP and CVRP, where we have local search operators, we also need $\alpha$ for energy reshaping (Eq.
-7).
-We empirically found that linearly increasing $\alpha$ from $0.5$ to $1.0$ along the epochs works well, and we use that setting for all TSP and CVRP experiments.
+In Equation 7, we also need a hyperparameter $\alpha$ that controls how much we augment the energy of a solution with the energy of the refined solution.
+If $\alpha=0$, we do not augment at all, and if $\alpha=1$, the energy of a solution is defined to be equal to the energy of the refined solution.
+We empirically found that linearly increasing $\alpha$ from $0.5$ to $1.0$ along the epochs works well, and we use that setting for all experiments.
 A.3 Computing resource
@@
 These encompass vehicle routing problems, bin packing problems, and scheduling problems.
-For a comprehensive understanding of each problem’s specific settings, we direct readers to the work by Ye et al.
-[79].
-Furthermore, for access to the exact source code pertaining to each problem, please refer to the repository available at  [^3] https://github.com/henry-yeh/DeepACO from Ye et al.
-[79].
+For a comprehensive understanding of each problem’s specific settings, we direct readers to the work by Ye et al., (2023).
+Furthermore, for access to the exact source code pertaining to each problem, please refer to the repository available at  [^2] https://github.com/henry-yeh/DeepACO from Ye et al., (2023).
 Traveling salesman problem
@@
 A CVRPTW instance consists of a set of customer nodes, a single depot node, the demands of each customer node, the vehicle capacity (as in CVRP), and additionally, the earliest and latest arrival times for each node, as well as the service time at each node.
-We generate the CVRPTW instance following the method provided by Berto et al.
-[6], which can be accessed here [^4] https://github.com/ai4co/rl4co/blob/main/rl4co/envs/routing/cvrptw/generator.py.
+We generate the CVRPTW instance following the method provided by Berto et al., (2023), which can be accessed here.
+[^3] https://github.com/ai4co/rl4co/blob/main/rl4co/envs/routing/cvrptw/generator.py
 Orienteering problem
@@
 We sample a random OP instance by first sampling the position of nodes, including the starting node, from a unit square $[0,1]^{2}$.
-And we set the prize for each node, following [19, 79], to be $p_{i}=(1+\lfloor 99\cdot\frac{d_{0,i}}{\max_{j=1}^{n}{d_{0,j}}}\rfloor)/100$, where $d_{0,i}$ is the Euclidean distance from the depot to node $i$.
-We set the maximum length to 4 in the OP100 task, following Kool et al.
-[41].
+And we set the prize for each node, following Fischetti et al., (1998); Ye et al., (2023), to be $p_{i}=(1+\lfloor 99\cdot\frac{d_{0,i}}{\max_{j=1}^{n}{d_{0,j}}}\rfloor)/100$, where $d_{0,i}$ is the Euclidean distance from the depot to node $i$.
+We set the maximum length to 4 in the OP100 task, following Kool et al., (2019).
 Prize-collecting traveling salesman problem
@@
 The bin packing problem (BPP) aims to minimize the number of bins when objects with different volumes need to be packed into a finite number of bins.
-The cost functions are often modified to maximize the utility of used bins [18, 79].
+The cost functions are often modified to maximize the utility of used bins (Falkenauer and Delchambre,, 1992; Ye et al.,, 2023).
 However, the direct use of the objective function employed in DeepACO can lead to an excessively large $\beta$ since the objective values tend to be saturated near 1.0; note that the objective values can not exceed 1.0 in that setting.
@@
 As demonstrated in Section 4, our algorithm utilizes the local search for off-policy training.
-The refined solutions from local search serve as exploitation batches, and they are also used for energy-reshaping (Eq.
-7).
-For some problems like traveling salesman problem (TSP), capacitated vehicle routing problem (CVRP), and CVRP with time window (CVRPTW), we used the well-established local search operators such as 2-opt, Swap* [74] and granular neighborhood search from PyVRP [77].
+The refined solutions from local search serve as exploitation batches, and they are also used for energy-reshaping (Equation 7).
+For some problems like traveling salesman problem (TSP), capacitated vehicle routing problem (CVRP), and CVRP with time window (CVRPTW), we used the well-established local search operators such as 2-opt, Swap* (Vidal,, 2022) and granular neighborhood search from PyVRP (Wouda et al.,, 2024).
 While domain-specific local search techniques are well developed for many combinatorial optimization problems, there are still tasks that lack effective local search solvers like Swap*.
-A potential candidate for a general-purpose local search solver in combinatorial optimization is the destroy-and-reconstruct local search operator.
-This method is inspired by the large neighborhood search approach described by Pisinger and Ropke [61].
+A potential candidate for a general-purpose local search solver in combinatorial optimization is the destroy-and-repair local search operator.
+This method is inspired by the large neighborhood search approach described by Pisinger and Ropke, (2019).
 Our implementation utilizes an evolutionary-style algorithm that repeatedly performs destruction, reconstruction, and top-K selection over a specified number of rounds.
@@
 Note that the number of symmetric trajectories for a given solution differs from problem types, as discussed in Appendix E.
-Reconstruction with temperature annealing.
+Repair with temperature annealing.
 Given a partial solution, we reconstruct a new solution $x^{\prime}$ using the policy similar to the decoding steps explained in Section 3.2.
@@
 Top-K selection.
-For a given solution $x$, we perform the aforementioned destruction and reconstruction multiple times (in batch), obtaining a set of refined solutions.
+For a given solution $x$, we perform the aforementioned destruction and repair multiple times (in batch), obtaining a set of refined solutions.
 Then we evaluate them and leave only top-K solutions among them in terms of solution quality.
@@
 D.1 Comparison between trajectory balance and forward-looking detailed balance
-We trained Attention Model (AM) [41] and heatmap-based policy from Joshi et al.
-[29] in an unsupervised way using FL-DB and TB loss.
-Table 6 shows the test results for 128 random TSP50 instances.
-The result is aligned with the previous observation from GFACS, showing the limitation of FL-DB loss for solving complex CO tasks.
-Table 6: Comparison between forward-looking detailed balance (FL-DB) and trajectory balance (TB) loss in TSP50, using Attention Model (AM) and heatmap-based method from Joshi et al.
-[29] as backbone policies.
+Table 7: Comparison between forward-looking detailed balance (FL-DB) and trajectory balance (TB) loss in TSP50, using Attention Model (AM) and heatmap-based method from Joshi et al., (2021) as backbone policies.
 TSP50
@@
 […]
+We trained Attention Model (AM; Kool et al.,, 2019) and heatmap-based policy from Joshi et al., (2021) in an unsupervised way using FL-DB and TB loss.
+Table 7 shows the test results for 128 random TSP50 instances.
+The result is aligned with the previous observation from GFACS, showing the limitation of FL-DB loss for solving complex CO tasks.
 D.2 Analysis of $\beta$ and diversity-optimality trade-off
@@
 D.3 Evaluation with other ACO variants
-Table 7: Evaluation results with different pheromone update rules, Elitist AS and MAX-MIN AS. We run $T=10$ ACO rounds with 100 ants in all settings.
+Table 8: Evaluation results with different pheromone update rules, Elitist Ant System (Elitist) and MAX-MIN Ant System (MAX-MIN). We run $T=10$ ACO rounds with 100 ants in all settings.
 TSP200 CVRP200
@@
 Concorde / PyVRP 10.72 0.00 27.93 0.00
-AS 10.91 1.77 28.45 1.86
-DeepACO + AS 10.77 0.45 28.44 1.80
-GFACS + AS 10.75 0.26 28.32 1.40
-Elitist AS 10.97 2.31 28.46 1.90
-DeepACO + Elitist AS 10.77 0.49 28.45 1.84
-GFACS + Elitist AS 10.75 0.28 28.34 1.47
-MAX-MIN AS 10.97 2.34 28.46 1.88
-DeepACO + MAX-MIN AS 10.78 0.57 28.44 1.82
-GFACS + MAX-MIN AS 10.75 0.32 28.33 1.43
+Ant System 10.91 1.77 28.45 1.86
+DeepACO + Ant System 10.77 0.45 28.44 1.80
+GFACS + Ant System 10.75 0.26 28.32 1.40
+Elitist 10.97 2.31 28.46 1.90
+DeepACO + Elitist 10.77 0.49 28.45 1.84
+GFACS + Elitist 10.75 0.28 28.34 1.47
+MAX-MIN 10.97 2.34 28.46 1.88
+DeepACO + MAX-MIN 10.78 0.57 28.44 1.82
+GFACS + MAX-MIN 10.75 0.32 28.33 1.43
 Our method is focused on enhancing the design of the heuristic matrix within the ACO framework.
 Thus, the advancement of the ACO algorithm itself (usually in pheromone update and action selection rule) is orthogonal to our work.
-To validate this claim, we evaluate our method with different ACO variants other than AS, namely Elitist Ant System [14], MAX-MIN Ant System [68].
+To validate this claim, we evaluate our method with different ACO variants other than Ant System, namely Elitist Ant System (Dorigo et al.,, 1996), MAX-MIN Ant System (Stützle and Hoos,, 2000).
 The results in Table 8 demonstrate that GFACS consistently outperforms both ACO with human-designed heuristic and DeepACO under all pheromone update rules tested.
 This underscores the superiority of GFACS in generating a high-quality heuristic matrix.
-D.4 Applying our method to other search algorithms
-Table 8: Evaluation results combine with active search as posterior inference method.
-TSP200 CVRP200
-Method Obj. Gap (%) Obj. Gap (%)
-Concorde / PyVRP 10.72 0.00 27.93 0.00
-Pretrained GFACS
-   – Off-policy TB Active Search 10.75 0.29 28.34 1.47
-   – On-policy RL Active Search 10.76 0.40 28.38 1.62
-Pretrained DeepACO
-   – On-policy RL Active Search 10.79 0.65 28.43 1.79
-In this subsection, we study active search [3] rather than the pheromone update of ACO.
-In ACO, pheromone update is a posterior inference given a prior distribution (heuristic) for the probability model that generates the solutions.
-Similarly, we can think of active search as a posterior update that implicitly updates a given probability model by adjusting the parameters that generate it.
-We sample 20 solutions and take 5 gradient steps to update the parameter at each iteration.
-After 50 iterations, we report the average performance of the active search using our method with 3 seeds of 128 instances for each task.
-Table 8 shows that GAFCS integrates better with active search than DeepACO.
-This indicates that GAFCS is effective not only with heuristic-based posterior inference but also with learning-based posterior inference.
-D.5 Comparison against TAM
-Table 9: Results for CVRP instances used in TAM. The results for AM and TAM is drawn from Hou et al. [24], and the result for DeepACO is drawn from Ye et al. [79].
+D.4 Comparison against TAM
+Table 9: Results for CVRP instances used in TAM. The results for AM and TAM is drawn from Hou et al., (2023), and the result for DeepACO is drawn from Ye et al., (2023).
 CVRP100 CVRP400
@@
 AM* 16.42 0.06 29.33 0.20
-TAM* 16.08 0.086 25.93 1.35
+TAM* 16.08 0.09 25.93 1.35
 DeepACO ($T=4$) 16.08 2.97 25.31 3.65
 DeepACO ($T=10$) 15.77 3.87 25.27 5.89
-GFACS ($T=4$) 15.77 1.07 24.95 3.23
-GFACS ($T=10$) 15.72 1.40 24.87 5.21
-We compare GFACS and DeepACO against CVRP instances used in TAM [24].
+GFACS ($T=4$) 15.77 2.91 24.95 3.23
+GFACS ($T=10$) 15.72 3.88 24.87 5.21
+We compare GFACS and DeepACO against CVRP instances used in Two-stage Divide Method (TAM; Hou et al.,, 2023).
 These instances have a different capacity constraint from those we used for the main experiment (Table 2).
-Please refer to [24] for more details regarding the instance generation.
-Moreover, to match the processing time to be comparative to DeepACO, we set $K$ (equal to the number of ants for ACO) to 16.
-D.6 Ablation study
+Please refer to Hou et al., (2023) for more details regarding the instance generation.
+Moreover, to match the processing time to be comparative to DeepACO, we set $K$ (the number of ants for ACO) to 16.
+D.5 Full results of the ablation study
 Table 10: Results of the ablation study.
 The mean and standard deviation are reported based on three independent runs.
-The reported gap is calculated against Concorde [2] for TSP and PyVRP [77] for CVRP.
+The reported gap is calculated against Concorde (Applegate et al.,, 2006) for TSP and PyVRP (Wouda et al.,, 2024) for CVRP.
 TSP500 CVRP500 BPP120
 Method Obj. Gap (%) Obj. Gap (%) Obj.
-GFACS 16.0842 $\pm$ 0.0043 1.56 64.1858 $\pm$ 0.0071 1.95 51.8428 $\pm$ 0.0706
+GFACS 16.8042 $\pm$ 0.0043 1.56 64.1858 $\pm$ 0.0071 1.95 51.8428 $\pm$ 0.0706
    – Off-policy Training 17.0861 $\pm$ 0.0028 3.26 64.4248 $\pm$ 0.0155 2.33 52.3158 $\pm$ 0.1171
@@
    – Shared $\varepsilon$-norm 17.5209 $\pm$ 0.0062 5.87 65.5069 $\pm$ 0.0171 4.05 90.7332 $\pm$ 9.1649
-    + VarGrad [82] 16.8570 $\pm$ 0.0086 1.88 64.2771 $\pm$ 0.0293 2.09 51.9504 $\pm$ 0.1222
-In this section, we present the comprehensive results of our ablation study, incorporating standard deviations from multiple independent runs (three runs) to ensure statistical significance.
-The findings demonstrate that each component of our technical proposals contributes to improvement, supported by empirical and statistically rigorous evidence.
+    + VarGrad Zhang et al., 2022b 16.8570 $\pm$ 0.0086 1.88 64.2771 $\pm$ 0.0293 2.09 51.9504 $\pm$ 0.1222
+In this section, we present the comprehensive results of our ablation study, including standard deviations from multiple independent runs (three runs) to ensure statistical significance.
+Table 10 demonstrates that each component of our technical proposals contributes to improvement, supported by empirical and statistically rigorous evidence.
+D.6 Results without local search on TSP and CVRP
+Table 11: Results obtained with and without local search (LS) at test time on TSP and CVRP are reported.
+The values represent the objective value averaged over 128 test instances and further averaged across three independent runs.
+Method TSP200 TSP500 TSP500 CVRP500
+ACO w/o LS 14.19 24.05 42.95 95.24
+DeepACO w/o LS 11.59 18.83 32.12 73.20
+GFACS w/o LS 12.63 21.72 34.04 81.03
+ACO w/ LS 10.91 17.38 28.45 64.62
+DeepACO w/ LS 10.77 16.84 28.44 64.49
+GFACS w/ LS 10.75 16.80 28.32 64.19
+We conduct an additional ablation study to assess the impact of local search at test time.
+Note that we use the same policies as in Section 5.1 or Section 5.2, which were trained with local search.
+The results are presented in Table 11.
+As expected, DeepACO (RL-based) performs better than ours when local search is not applied.
+This outcome can be attributed to two factors: (1) The RL-based method tends to generate non-diverse low-energy samples, whereas the GFlowNet-based method (GFACS) produces a more diverse set of candidates; and (2) Due to the energy reshaping (Eq.
+(7) in Section 4.1), the policy is optimized with respect to a modified energy map that assumes the application of local search, resulting in suboptimal performance when local search is omitted.
+This second factor also explains why GFACS outperforms other methods when local search is applied.
+D.7 Runtime analysis
+Table 12: Average running time (in seconds) to solve a single instance for Table 1.
+$N$ ACO DeepACO GFACS
+TSPLib 100-299 1.3 1.3 1.3
+300-699 11.2 10.5 10.8
+700-1499 89.3 72.5 77.0
+CVRPLib 100-299 15.5 15.5 15.5
+300-699 34.2 35.1 35.5
+700-1001 60.0 59.8 61.8
+For completeness, in Table 12, we report the runtime for each algorithm on the real-world instances (Table 1 in Section 5.1).
+Although all algorithms have similar inference times, DeepACO and GFACS occasionally run faster than ACO.
+This is attributed to the increased complexity of real-world instances: both DeepACO and GFACS benefit from an NN-based amortized prior that enables them to start closer to a local optimum, resulting in earlier termination of the local search.
+This also explains why DeepACO is slightly faster than GFACS, as it produces a more locally optimized initial solution.
 Appendix E Symmetric solutions of combinatorial optimization
@@
 In contrast, for a single-vertex solution ($N=1$), the trajectory is identical, reflecting the singularity of the route.
-Appendix F Extended related works
+Appendix F Related works
 F.1 Deep constructive policy for neural combinatorial optimization
 Constructive policy methods systematically build solutions by incrementally adding elements, beginning with an empty set and culminating in a complete solution.
-This approach ensures compliance with constraints by progressively narrowing the scope of actions, making it versatile for application in domains like vehicle routing [57, 41], graph combinatorial optimization [30, 1, 84], and scheduling [81].
+This approach ensures compliance with constraints by progressively narrowing the scope of actions, making it versatile for application in domains like vehicle routing (Nazari et al.,, 2018; Kool et al.,, 2019), graph combinatorial optimization (Khalil et al.,, 2017; Ahn et al.,, 2020; Zhang et al.,, 2023), and scheduling (Zhang et al.,, 2020).
 In deep learning, various techniques adopt this constructive policy framework.
-Notably, the attention model (AM) [41] has been pivotal in vehicle routing problems (VRPs), establishing a new standard for stabilizing REINFORCE training.
-Subsequent advancements in AM have included enhanced decoder architecture [67, 78, 47], improved REINFORCE [76] baseline assessments [43, 37], leveraging symmetries for sample efficiency [32], the application of population-based reinforcement learning (RL) [21], ensemble-based learning [27, 86] and refined solution-building methods [16].
-Recent advancements in hierarchical algorithms exploit AM-based policies for solution optimization: one policy broadly generates solutions while another refines them locally [36, 24, 80].
-Additionally, AM-driven constructive policies have been effectively applied in hardware routing [34, 35] and placement tasks [60, 31].
+Notably, the attention model (AM; Kool et al.,, 2019) has been pivotal in vehicle routing problems (VRPs), establishing a new standard for stabilizing REINFORCE training.
+Subsequent advancements in AM have included enhanced decoder architecture and decoding procedure (Son et al.,, 2024; Xin et al.,, 2021; Luo et al.,, 2023; Drakulic et al.,, 2023), improved REINFORCE (Williams,, 1992) baseline assessments (Kwon et al.,, 2020; Kim et al.,, 2022), leveraging symmetries for sample efficiency (Kim et al., 2024a, ), the application of population-based reinforcement learning (RL) (Grinsztajn et al.,, 2023), ensemble-based learning (JIANG et al.,, 2023; Zhou et al.,, 2024).
+Recent advancements in hierarchical algorithms exploit AM-based policies for solution optimization: one policy broadly generates solutions while another refines them locally (Kim et al.,, 2021; Hou et al.,, 2023; Ye et al.,, 2024; Zheng et al.,, 2024).
 Graph combinatorial optimization, particularly in problems like maximum independent sets, has diversified from contrive VRPs policies due to their local decomposability and the relative ease of training value functions for partial solutions where VRPs are not locally decomposable due to their global contains.
-Starting with the influential works by Khalil et al.
-[30], which employed deep Q networks [55] for graph policy training, creating solutions constructively, Ahn et al.
-[1] introduced an innovative Markov Decision Process (MDP) to generate multiple solution components per iteration by exploiting the local decomposability of graph combinatorial problems.
-Zhang et al.
-[84] further developed this concept, training their MDP with forward-looking GFlowNets [59].
-Our research aligns with the deep constructive policy framework, where GFlowNets constructively build actions from an initial empty solution ($s_{0}$) to a final complete solution ($s_{n}$).
-Additionally, our methods include a heatmap-based approach, detailed in Section F.3.
+Starting with the influential works by Khalil et al., (2017), which employed deep Q networks (Mnih et al.,, 2013) for graph policy training, creating solutions constructively, Ahn et al., (2020) introduced an innovative Markov Decision Process (MDP) to generate multiple solution components per iteration by exploiting the local decomposability of graph combinatorial problems.
+Zhang et al., (2023) further developed this concept, training their MDP with forward-looking GFlowNets (Pan et al.,, 2023).
+Our research aligns with the deep constructive policy framework, where we build a solution from an initial empty solution ($s_{0}$) based on the policy trained with GFlowNets.
+Additionally, our method is included in the heatmap-based approaches, detailed in Section F.3.
 F.2 Deep improvement policy for neural combinatorial optimization
-The improvement policy focuses on refining complete solutions, similar to local search methods like 2opt [11].
+The improvement policy focuses on refining complete solutions, similar to local search methods like 2opt (Croes,, 1958).
 Its advantage lies in achieving comparable performances to constructive policies, especially when allotted sufficient time for extensive improvement iterations.
-Researchers typically adapt existing heuristic local search techniques, creating learning policies that emulate these methods through maximizing expected returns via reinforcement learning [12, 22, 48, 49, 50, 33].
-Also, recent works have explored active search to refine solutions via policy adjustments [23, 66] or sampling the conditions from a continuous latent space [7].
-Moreover, alternative strategies using discrete Langevin sampling instead of neural networks have shown promising outcomes [69].
+Researchers typically adapt existing heuristic local search techniques, creating learning policies that emulate these methods through maximizing expected returns via reinforcement learning (d O Costa et al.,, 2020; Hottung and Tierney,, 2020; Ma et al.,, 2021, 2022, 2023; Kim et al.,, 2023).
+Also, recent works have explored active search to refine solutions via policy adjustments (Hottung et al.,, 2022; Son et al.,, 2023) or sampling the conditions from a continuous latent space (Chalumeau et al.,, 2023).
+Moreover, alternative strategies using discrete Langevin sampling instead of neural networks have shown promising outcomes (Sun et al.,, 2023).
 This approach serves as a complementary extension to deep constructive policies, as it often enhances solutions initially formulated by constructive methods.
@@
 F.3 Heatmap-based neural combinatorial optimization
-The heatmap-based method [29] entails the acquisition of a graph representation, referred to as the “Heatmap," from input graph instances.
+The heatmap-based method (Joshi et al.,, 2021) entails the acquisition of a graph representation, referred to as the “Heatmap," from input graph instances.
 This Heatmap serves as a prior distribution, influencing the generation of a solution.
-However, this approach is not self-contained and requires an additional solution construction step, which may involve techniques like beam search [29], dynamic programming [42], Monte Carlo tree search [20, 62, 70] or greedy search with a guide using an objective function [44].
+However, this approach is not self-contained and requires an additional solution construction step, which may involve techniques like beam search (Joshi et al.,, 2021), dynamic programming (Kool et al.,, 2022), Monte Carlo tree search (Fu et al.,, 2021; Qiu et al.,, 2022; Sun and Yang,, 2023; Xia et al.,, 2024) or greedy search with a guide using an objective function (Li et al.,, 2023).
 The key advantage of the heatmap-based method lies in its exceptional generalizability when compared to constructive policies.
@@
 A key advantage of our method over many other supervised learning-based heatmap approaches is its independence from optimally labeled data.
-Broader Impact
-We proposed a new probabilistic approach that can be trained using off-policy learning and applied to various combinatorial optimization problems, such as vehicle routing and scheduling.
-This method is relevant to industrial engineering, particularly in optimizing transportation and manufacturing processes, and can potentially extend to scientific discovery through chemical structure and biological sequence optimization.
-It is important to consider the ethical implications of probabilistic optimization methods, as they carry the potential for misuse in fields such as harmful drug discovery.
-While these algorithms offer several benefits, their application should be carefully monitored to ensure responsible and ethical use, minimizing the risk of unintended negative consequences.
+F.4 GFlowNets for neural combinatorial optimization
+Originally proposed as a framework for learning policies over discrete compositional variables (Bengio et al., 2021a, ), GFlowNets have since been applied to combinatorial optimization problems (Zhang et al.,, 2023; Zhang et al., 2022b, ).
+They offer unique advantages over reinforcement learning methods, particularly in generating diverse candidate solutions and enabling off-policy training.
+Although prior studies have explored the diversity aspect (Zhang et al.,, 2023; Zhang et al., 2022b, ), our work is, to our knowledge, the first to empirically validate the benefits of off-policy training in combinatorial optimization settings.
+To underscore our contribution, we conceptually compare our method against two recent approaches that incorporate off-policy training for GFlowNets: local search GFlowNets (LS-GFN) (LS-GFN; Kim et al., 2024b, ) and the hindsight-like off-policy strategy proposed by Zhu et al., (2023).
+Note that, in Section 5.4, we provide an empirical comparison of our approach with FL-DB (Zhang et al.,, 2023) and VarGrad (Zhang et al., 2022b, ).
+The key difference between LS-GFN and our approach is that our method addresses conditional inference problems.
+This requires training conditional GFlowNets that operate under distinct reward landscapes for different conditions, whereas LS-GFN is designed for unconditional problems.
+Moreover, the local search operators themselves differ.
+LS-GFN uses a back-and-forth mechanism that partially backtracks trajectories and reconstructs them using its own policy, thereby limiting exploration to sequences sharing identical root segments.
+In contrast, our local search methods, such as the 2-opt operator for the TSP, explore the solution space within a Hamming ball, allowing for more diverse exploration without the constraint of shared initial sequences.
+Furthermore, our destroy-and-repair local search method (Appendix C) can be viewed as an extension of LS-GFN, incorporating symmetricity-aware destruction to mitigate the limitations of the back-and-forth operator.
+In Zhu et al., (2023), they tries to model a family of policies using hypernetworks to tackle multi-objective optimization problems.
+They also propose a hindsight-like off-policy strategy, which enables sharing of off-policy samples among policies configured with varying objective mixing coefficients, thereby enhancing sample efficiency during training.
+In contrast, our method leverages hindsight observations from a (deterministic) local search (Section 4.1).
+Specifically, we adjust the energy of an original sample by incorporating the energy after local search, thereby quantifying how readily the sample can be converted into a low-energy solution.